A Complete Bibliography of Publications in Computational Statistics & Data Analysis (2020–2029)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

10 April 2024 Version 1.31

Title word cross-reference

2 [196, 480]. ^{2} [40]. ^{nd} [670]. C_{p} [486]. D [424]. ℓ_{0} [466]. ℓ_{1} [664, 19]. ℓ_{2} [664]. ℓ_{q} [185]. I [387]. K [76, 9, 136, 432, 314]. L [622]. L_{0} [782, 805]. L_{q} [567]. λ [217]. M [116, 185, 508, 118, 351]. p [712]. R [234]. R^{2} [695]. S [693]. t [221, 128]. U [547]. V [547]. ε [653].

-balls [185]. -D [480]. -estimation [118, 351]. -estimators [693, 185]. -explorations [653]. -level [196]. -loss [567]. -means [136, 314]. -moments [622]. -optimal [424]. -partite [9]. -penalized [19]. -process [221]. -quantile [508]. -Regularized [466, 664]. -statistics [547].

-values [712]. -vine [234].

19 [761, 654, 314].

2020 [48, 122, 171, 11, 107, 114, 37, 90, 157, 139, 129]. 2021 [242, 322, 372, 212, 193, 289, 307, 224, 259, 363, 354, 338]. 2022 [421, 489, 560, 392, 474, 383, 464, 399, 446, 548, 521, 501]. 2023 [635, 696, 753, 588, 574, 671, 689, 657, 605, 729, 724, 707]. 2024 [801, 776, 763, 828, 842, 820, 789]. 2nd [706].

abdominal [534]. absolute [7]. abundance [75]. Accelerate [700, 117]. accelerated [466, 701, 344, 850, 523]. Accelerating [709]. acceleration [228]. accelerator [577]. accommodation [629]. accuracy [200]. achieve [513]. achievement [286]. acquired [240]. activation [152]. Active [365, 95]. actuarial [835]. acute [587]. adapted [177]. Adaptive [609, 405, 167, 210, 106, 811, 300, 693, 458, 590, 632, 583]. addition [276, 385]. Additive [741, 784, 617, 598, 817, 680, 214, 358, 347, 780, 240, 554, 401, 584, 76, 414, 100]. addresses [668]. Adjacency [94]. Adjacency-based [94]. Adjusting [642]. adjustment [162]. ADMM [14]. adulteration [293]. advanced [192]. Aerobics [437]. affine [493, 513]. affine-equivariant [493]. against [164, 271]. Agglomerative [569]. aggregation [99, 841]. Airflow [534]. AJIVE [643]. aka [691]. algorithm [425, 416, 208, 761, 804, 499, 121, 337, 660, 196, 125, 10, 614, 467, 734, 403, 563, 844, 493, 286, 14, 806, 731, 228]. algorithms [309, 365, 830, 50, 779, 424, 262, 226, 109]. allelic [492]. alliance [305]. allocation [565]. allowing [484]. altered [9]. Alternating [147, 738, 810]. alternatives [164, 513]. Alzheimer [282, 153]. analyses [271]. Analysis [254, 147, 425, 543, 587, 296, 706, 537, 562, 282, 819, 405, 453, 536, 264, 592, 593, 360, 200, 22, 125, 371, 340, 387, 177, 352, 524, 794, 812, 163, 70, 25, 459, 629, 233, 765, 176, 737, 54, 705, 747, 615, 241, 397, 284, 283, 62, 166, 616, 630, 430, 59, 21, 32, 482, 718, 618, 681, 669, 13]. analyte [47]. analyzers [325]. analyzing [831, 849]. Angle [244, 505]. Angle-based [244, 505]. anomalous [443]. anomaly [443]. ANOVA [830]. Application [36, 87, 414, 609, 580, 295, 799, 296, 216, 282, 42, 33, 371, 396, 177, 305, 663, 717, 187, 631, 437, 557, 718, 15, 681]. Applications [604, 602, 761, 545, 835, 453, 126, 99, 827, 655, 654, 767, 54, 386, 67, 278, 484, 314, 16, 482]. applied [556]. approach [362, 728, 435, 532, 20, 598, 27, 296, 217, 571, 407, 106, 778, 211, 93, 252, 849, 178, 780, 291, 98, 53, 241, 235, 250, 832, 266, 100, 168, 198, 700, 111]. approaches [611, 634]. Approximate [267, 462, 106, 41, 87, 29, 710, 50, 778, 467, 734]. approximated [783]. approximation [514, 155, 608, 10]. approximations [261, 214, 347, 329, 197]. April [48, 242, 421, 635, 801]. Archimedean [218, 232]. area [409, 92, 703, 702, 96]. area-interaction [409]. area-wise [96]. areal [816]. arm [544, 632]. ARMA [254, 822]. armed [653]. arrays [691]. arrival [230]. article [13]. aspect [649]. assess [499, 240]. Assessing [362, 371]. Assessment [585, 591]. assisted [182]. association [682, 587, 165, 326]. assuming [271]. assumptions [390]. asymmetric [218, 237, 272, 5]. Asymptotic [113, 550, 487, 513, 547]. asynchronous [284]. AUC [220]. augmentation [426]. Augmented [141, 798, 402]. August [122, 322, 489, 696]. auto [520, 233]. auto- [233]. auto-covariances [520]. autocorrelated [519]. autocorrelation [233]. autocovariances [205]. automated [639]. Automatic [126, 647]. autonomous [571]. autoregression [570, 448, 445]. autoregressive [680, 357, 550, 536, 633, 595, 390, 41, 480, 229, 666]. available [453]. avenue [659]. average [623, 480, 62, 772, 303]. averaging [701, 182, 331, 402, 713, 84, 273, 79, 380, 509]. awareness [825].

B [318]. B-spline [318]. backfitting [97]. bagging [582]. balanced [796]. Balancing [632, 370]. balls [185]. band [238, 411]. banded [766]. bandit [653]. bands [735]. bandwidth [201, 318]. BART [638, 780]. BART-based [638]. based [728, 622, 20, 425, 710, 492, 416, 799, 511, 543, 817, 704, 293, 80, 78, 553, 628, 451, 694, 434, 481, 92, 505, 209, 558, 367, 272, 816, 214, 264, 774, 319, 495, 371, 328, 300, 703, 211, 387, 336, 638, 800, 148, 579, 186, 432, 285, 240, 645, 94, 510, 552, 152, 818, 739, 554, 586, 513, 634, 61, 721, 433, 443, 187, 815, 348, 332, 226, 276, 410, 149, 398, 380, 24, 244, 478, 266, 170, 168, 198, 711, 32, 234, 766]. bases [318]. Bayes [156, 340, 387, 440, 250]. Bayesian [539, 265, 239, 104, 568, 569, 258, 587, 409, 760, 151, 562, 222, 350, 92, 546, 394, 536, 462, 214, 47, 633, 792, 593, 319, 360, 106, 371, 387, 726, 755, 36, 448, 347, 352, 794, 849, 767, 134, 780, 374, 733, 765, 659, 797, 497, 737, 516, 386, 615, 739, 53, 732, 181, 45, 676, 473, 39, 666, 65, 184, 368, 578, 62, 423, 535, 82, 74, 170, 281, 711, 408, 749]. behavior [815]. Benchmark [44]. benthic [609]. Berk [714]. Bernoulli [609, 504]. Bernstein [597]. best [324]. beta [408, 644]. Beta-GOS [644]. better [719]. between [180, 34, 587, 218, 819, 257, 134]. between-block [218]. Beyond [441, 610, 235, 398]. bi [337, 304]. bi-clustering [337]. bi-level [304]. Bias [89, 102, 332, 5, 423]. Bias-corrected [332]. biased [28]. big [78, 131, 198, 197]. binary [425, 587, 431, 231, 750, 300, 95, 252, 663, 6, 326, 720]. Biodosimetry [67]. biological [827]. biomedical [36]. biomonitoring [784]. BIOSTATISTICS [670]. bipartite [767]. biplots [369]. Birnbaum [345]. bitcoin [668]. bivariate [771, 450, 203, 271, 554]. Block [830, 148, 621, 362, 296, 218, 770, 571, 781, 767, 189, 479, 215, 731, 810, 814, 410]. Block-diagonal [621]. block-missing [410]. Block-wise [830]. blocking [196]. blocks [218, 189]. bMOM [416]. Board [172, 1, 12, 38, 49, 91, 108, 115, 123, 130, 140, 158, 194, 213, 225, 243, 260, 290, 308, 323, 339, 355, 364, 373, 384, 393, 400, 422, 447, 465, 475, 490, 502, 522, 549, 561, 575, 589, 606, 636, 658, 672, 690, 697, 708, 725, 730, 754, 764, 777, 790, 802, 821, 829, 843]. body [771]. boosting [425, 206]. Bootstrap [553, 416, 602, 93, 148, 4]. bootstrap-based [148]. Bootstrapping [63, 390, 699]. Borrowing [82]. boundary [797]. bounded [408]. Brain [536, 9, 717, 74]. breast [681]. bridge [405, 65]. bridge-randomized [65]. Brownian [405]. building [356]. Burn [808]. Burn-in [808]. bursty [230].

calculation [712, 56]. Calibrated [787, 241]. calibration [286, 7]. cancer [681]. canonical [105]. capability [389]. cardinal [497]. Carlo [556, 562, 160, 79, 39, 52, 368, 640]. case [498, 64, 76]. case-cohort [498, 64]. case-control [64]. Categorical [369, 582, 673, 180, 376, 132, 736, 285, 26, 531]. Cauchy [387]. Causal [256, 543, 642, 587, 499, 702, 437, 311]. cause [847, 681]. cause-of-failure [681]. cause-specific [847]. Cellwise [116]. Censored [310, 201, 378, 282, 498, 346, 347, 344, 849, 204, 291, 503, 76, 247, 826, 24, 153, 59]. censoring [69, 145, 391, 240, 554, 76]. Center [437, 698]. Centered [622]. central [366]. chain [562, 444]. Change [677, 824, 553, 405, 31, 788]. Change-point [677, 405]. changepoint [580, 472]. characteristics [180, 251]. checking [3]. checks [80]. childhood [587]. choice [217]. Cholesky [288, 756, 254, 471]. chordal [491]. chromosome [165]. circle [660]. circular [442, 603, 45, 666]. class [309, 376, 264, 345, 352, 95]. classical [88]. Classification [630, 349, 44, 216, 668, 637, 750, 389, 36, 177, 95, 244]. classifier [58, 663]. classifiers [682]. clear [196]. clinical [541, 457, 565, 590, 681]. cloning [40]. Cluster [619, 481, 271, 796]. Cluster-robust [619]. cluster-weighted [481]. clustered [592, 428, 583]. Clustering [303, 431, 246, 800, 517, 620, 582, 416, 569, 761, 770, 313, 337, 571, 781, 827, 136, 147, 848, 181, 45, 61, 720, 187, 184, 478]. Clustering-based [800]. Clusterwise [302]. Co [147, 275, 61]. Co-clustering [147, 61]. co-sparse [275]. coalescents [257]. coding [398]. coefficient [180, 190, 97, 188, 269, 103, 506, 178, 127, 273, 500, 306, 3, 519, 100, 454, 583, 749]. coefficients [317, 16]. cohort [498, 64]. Collaboration [432]. collaborative [767]. combination [274, 276]. Combining [348]. Common [15]. Communication [351, 745, 263, 353]. Communication-efficient [351, 745, 263, 353]. Community [266, 432, 184, 731]. comparative [34, 6]. Comparing [634, 270]. Comparison [164, 143, 50, 328, 565, 497, 420, 472]. comparisons [281]. Competing [165, 304, 245, 146, 120]. competitions [305]. complementary [476]. complete [530]. Completely [155]. complex [151, 694, 47, 98]. Complexity [514, 844]. component [163, 629, 176, 705, 397, 678, 283, 32, 482]. components [298, 283]. Composite [331, 681, 799, 222, 353]. compositional [845, 125, 388]. compound [88, 723]. Compromise [274]. Computation [450, 106, 297, 2, 267, 405, 47, 54, 415, 700]. Computational [13]. computationally [577, 666]. computations [357]. computer [711]. computerized [286]. Computing [477, 547, 71]. concomitant [128]. Conditional [207, 428, 439, 57, 795, 7, 451, 179, 692, 270, 357, 462, 407, 774, 336, 655, 87, 306, 772, 491, 651]. Conditional-mean [795]. Confidence [301, 445, 769, 553, 238, 200, 411, 809, 71, 460]. Conformal [735]. confounding [642, 618]. conjugate [608, 189, 228]. connected [664]. connection [359]. connectivity [9, 717, 152]. conquer [78]. considerations [827]. consistency [4, 315]. constrained [365, 597]. constraints [394, 585, 19, 460]. constructed [151]. Construction [232, 476]. contaminated [542]. contamination [600]. contemporaneous [68]. context [300]. context-dependent [300]. contingency [460]. continuous [313, 444]. contour [441]. contrast [650]. control [564, 677, 64, 812, 714, 503, 667]. controls [736]. conversion [153]. convex [752, 117, 848, 433]. convexity [441]. convolution [694]. convolutional [216]. Copula [554, 341, 564, 755, 146, 625, 552, 661]. copulas [504, 218, 232, 462, 159]. corporate [414]. correct [131]. Corrected [84, 332]. Correcting [423, 741]. correction [5]. correlated [486, 262, 419]. Correlation [374, 180, 20, 33, 800, 417, 428, 127]. cost [271, 244]. cost-effectiveness [271]. cost-sensitive [244]. count [757, 265, 576, 452, 470, 246, 87, 395, 795, 620]. counts [579]. covariables [433, 332]. covariance [568, 288, 89, 144, 694, 774, 320, 25, 717, 226, 380, 604, 723, 487]. covariances [520]. Covariate [644, 648, 370, 557]. Covariate-modulated [648]. covariates [63, 282, 596, 101, 41, 388, 26, 538, 128, 632, 59, 100, 263]. COVID [761, 654, 314]. COVID-19 [761, 654, 314]. Cox [650, 106, 748]. Cramér [710]. cranial [587]. credit [775]. crime [87]. criterion [84, 486, 332]. cross [594, 728, 209, 227, 592, 563, 79, 370]. cross-distance [563]. cross-sectional [209, 370]. cross-validation [594, 227, 592, 79]. cumulative [359, 380]. cure [692, 64, 344, 775, 247, 414]. curve [126, 23, 809]. curves [270, 656, 397, 143]. CUSUM [148]. CV [652]. CV-entropy [652]. CVA [369].

D [480]. Data

[40, 613, 13, 3, 582, 441, 728, 435, 542, 771, 514, 201, 265, 425, 831, 576, 452, 413, 44, 223, 799, 817, 813, 642, 602, 684, 78, 138, 611, 706, 378, 102, 564, 628, 537, 470, 426, 677, 761, 69, 282, 190, 209, 217, 337, 453, 637, 656, 816, 431, 246, 450, 83, 304, 792, 827, 833, 231, 22, 125, 371, 345, 320, 778, 703, 599, 726, 347, 177, 131, 254, 524, 787, 688, 803, 103, 344, 202, 751, 849, 85, 120, 285, 428, 28, 204, 271, 639, 508, 291, 784, 467, 127, 94, 507, 26, 98, 497, 737, 516, 503, 152, 386]. data [615, 160, 81, 278, 500, 181, 326, 471, 45, 420, 87, 17, 61, 823, 57, 351, 472, 406, 720, 461, 518, 96, 284, 631, 76, 166, 361, 353, 327, 395, 411, 519, 630, 410, 8, 247, 685, 826, 153, 414, 621, 643, 742, 59, 783, 509, 74, 71, 170, 198, 557, 419, 661, 454, 620, 718, 310, 197, 681]. Data-cloning [40]. Data-driven [613, 3, 217, 524]. datasets [362, 112, 374, 477, 348, 745]. Dealing [470]. debiased [433, 791]. December [171, 372, 560, 753]. decision [590, 535]. decision-theoretic [535]. decomposition [756, 747]. decompositions [254, 471]. deconvolution [676]. decorrelation [210]. decreasing [119, 612]. Decrement [245]. Deep [637, 391, 321, 528, 476, 469]. DeepQuantreg [391]. default [775, 414]. defining [555]. definite [526]. degree [42, 79]. Degrees [136]. Delaunay [173]. delayed [535]. delta [329]. delta-variance [329]. denoising [452]. dense [513]. Densely [664]. densities [532, 762, 451, 272]. Density [249, 46, 685, 195, 365, 550, 268, 5, 77, 345, 318, 359, 581, 676, 170]. dependence [174, 209, 203, 336, 648, 430]. dependencies [68]. dependent [582, 644, 69, 145, 22, 300, 186, 486, 508, 59, 485]. depression [352]. depth [555, 746, 297]. depths [267]. derivative [486]. Design [110, 201, 274, 778, 476, 654, 818]. designs [609, 641, 50, 196, 300, 457, 476, 565, 98, 424, 377, 141, 616, 299, 600, 385]. detail [223]. Detecting [9, 655, 788]. detection [542, 580, 817, 135, 236, 190, 280, 379, 210, 389, 432, 629, 184, 647, 731, 266, 824]. determination [306, 683, 111]. Determining [46]. deterministic [493]. developmental [352]. devices [580]. diagnosis [233]. diagnostic [769, 200]. diagonal [621]. difference [827, 655]. differencing [581]. different [69]. differential [319, 340, 654, 507, 54, 262]. diffusion [31, 443, 29]. diffusions [444]. Dimension [252, 255, 410, 651, 576, 434, 182, 750, 438, 177, 563, 513, 646, 343, 226, 149, 24]. Dimension-reduced [255]. dimension-reduction [750]. dimension-wise [576]. dimensional [728, 735, 532, 205, 758, 617, 156, 44, 295, 543, 602, 144, 466, 706, 237, 207, 72, 367, 656, 331, 162, 312, 30, 524, 812, 803, 70, 417, 459, 751, 120, 178, 837, 559, 127, 507, 159, 26, 625, 510, 503, 219, 386, 6, 81, 500, 133, 721, 433, 14, 567, 436, 806, 277, 810, 809, 621, 382, 429, 604, 785, 32, 723, 718, 791, 310, 445]. Dimensionality [111]. dimensions [241, 330]. direct [241]. directed [203]. direction [810]. directional [74]. directions [541]. Dirichlet [608, 620]. Dirichlet-multinomial [620]. disclosure [564]. discovery [334, 812, 714, 667]. Discrepancy [819, 261]. discretely [596, 31]. discriminant [592, 177, 54, 241, 21]. disease [320, 282]. dispersion [31]. Dissimilarity [313, 559]. distance [710, 774, 563, 332, 226]. distortion [7]. Distributed [458, 361, 198, 791, 758, 351, 353, 327, 731, 263]. Distribution [268, 42, 238, 359, 321, 624, 496, 67, 547, 255, 411, 478, 56]. distributional [743, 423]. distributions [442, 155, 511, 2, 113, 692, 264, 203, 291, 645, 181, 850, 79, 96, 187]. divergence [655]. divergent [683, 604]. diverging [263]. diversity [257]. divide [78]. divisive [569]. dose [299, 600]. dose-response [600]. double [347, 111]. Doubly [437, 378, 830, 304, 57]. doubly-censored [378]. driven [613, 217, 524, 3]. driving [571]. drug [276]. drugs [274]. dual [830]. duration [146, 294]. during [413]. dwell [496]. dwell-time [496]. Dynamic [717, 768, 576, 223, 741, 371, 744, 188, 523, 519, 687, 220]. Dynamical [507, 248]. dynamics [138, 633, 381].

earthquake [216]. ecological [827]. edge [189]. Editorial [706, 669, 172, 1, 12, 38, 49, 91, 108, 115, 123, 130, 140, 158, 194, 213, 225, 243, 260, 290, 308, 323, 339, 355, 364, 373, 384, 393, 400, 422, 447, 465, 475, 490, 502, 522, 549, 561, 575, 589, 606, 636, 658, 672, 690, 697, 708, 725, 730, 754, 764, 777, 790, 802, 821, 829, 843]. educational [386]. EEG [413, 74]. effect [642, 540, 145, 558, 585, 506, 404, 850, 523, 311]. Effective [46, 362, 509]. effectiveness [271]. effects [728, 190, 357, 488, 371, 10, 75, 702, 619, 262, 128, 661, 329, 772]. efficacy [781]. Efficient [412, 236, 662, 533, 625, 54, 818, 262, 56, 499, 351, 377, 666, 14, 353, 411, 806, 745, 266, 199, 822, 263]. eigensystems [813]. elastic [693]. ellipsoids [477]. elliptical [556, 705]. EM-type [109]. embedded [809]. Embedding [287]. embeddings [285]. Empirical [653, 66, 826, 557, 104, 80, 2, 113, 324, 142, 359, 787, 186, 98, 737, 443, 250, 73, 175]. enclosing [477]. endogenous [253]. endpoint [300]. endpoints [33, 499]. Ensemble [756, 320, 58, 343]. ensembles [173, 634]. Entropy [511, 586, 652]. Entropy-based [511]. enumeration [641]. epidemic [654, 381]. equality [34, 510, 382, 785, 460]. equation [543, 664, 121, 654, 219, 262]. equations [231, 340, 428, 507, 530]. Equivalence [376]. equivariant [493]. ERGMs [432]. Erlang [794]. Error [350, 60, 494, 258, 346, 238, 773, 211, 387, 131, 219, 279, 411, 454, 822]. errors [89, 191, 237, 97, 390, 486, 3, 519, 7, 557]. errors-in-variables [89, 97, 3]. Escaping [309]. Essential [439]. estimands [460]. estimate [532, 89]. estimates [773]. Estimating [203, 305, 702, 151, 365, 231, 428, 530, 366, 168]. Estimation [223, 145, 656, 23, 716, 46, 79, 96, 738, 584, 538, 76, 277, 805, 454, 766, 220, 594, 441, 673, 494, 174, 195, 155, 539, 265, 758, 710, 576, 452, 183, 740, 63, 799, 248, 208, 526, 191, 216, 611, 680, 102, 650, 572, 236, 40, 161, 692, 92, 190, 597, 232, 167, 64, 268, 695, 101, 320, 703, 599, 693, 318, 444, 520, 359, 800, 715, 533, 787, 836, 727, 186, 249, 202, 486, 463, 294, 41, 204, 374, 837, 623, 717, 734, 415, 775, 160, 496, 241, 118, 586, 57, 351, 279, 397, 483, 518, 666, 52, 607, 14, 361, 255, 523]. estimation [395, 411, 647, 437, 685, 745, 601, 88, 478, 840, 29, 315, 311, 791, 822, 487]. estimator [566, 550, 77, 433, 263]. estimators [156, 312, 5, 345, 693, 185, 343, 619, 427]. Evaluating [211]. evaluation [33]. even [641]. even-odd [641]. event [552, 554, 160, 661]. events [230]. evolution [106]. evolutionary [39]. evolving [607]. exact [778, 377]. exact-approximate [778]. example [662]. Exceedance [714]. exchange [286]. exciting [456]. expected [43]. expectile [163, 646]. experiment [413, 540, 274]. experimental [50, 377]. experiments [525, 841, 276, 711, 385]. experts [291]. explained [551]. Explaining [682]. Explicit [294]. Explicit-duration [294]. exploiting [488]. explorations [653]. Exploring [587, 352, 320]. explosive [550]. exponential [27, 229]. exposure [617]. extended [105]. Extending [631, 128]. extension [516, 678]. Extrapolation [494]. Extremal [570]. extreme [230, 733].

f [10]. f-SAEM [10]. Factor

[841, 425, 196, 264, 162, 387, 325, 751, 275, 159, 440, 674, 277]. factored [629]. factorial [196]. factorization [288]. factors [768]. failure [466, 701, 439, 850, 675, 76, 247, 59, 681]. false [334, 812, 714, 667]. families [556]. family [442, 272, 23]. Fast [673, 265, 248, 357, 593, 347, 359, 639, 712, 734, 404, 273, 786, 804, 214, 577, 10, 467]. Faster [160]. FDR [677, 503]. feasible [625]. Feature [26, 503, 366, 137, 44, 30, 120, 285, 469]. features [503, 667]. February [11, 212, 392, 588, 776]. feedforward [824]. Fellegi [631]. FGM [504]. field [141]. fields [799, 62, 292]. file [435]. file-matching [435]. filter [44, 366, 652]. filtering [767, 381, 797, 87]. Filters [341, 666]. Finding [121, 50]. finite [412, 703, 325, 298, 128]. finite-population [703]. first [834, 342, 305]. first-order [834, 342]. Fisher [415]. fit [762, 239, 834, 556, 628, 613, 75, 67, 55, 699, 215, 380, 419]. fitted [678]. Fitting [358, 728, 227, 660, 530]. Fixed [201, 190, 33, 654]. Flexible [441, 837, 496, 264, 792, 815]. FMR [70]. fMRI [371, 152, 420]. follow [692]. follow-up [692]. footrule [721]. forecast [786]. forecasting [20]. forest [468]. forests [662, 488, 469]. form [774]. forward [234]. Fourier [427]. fractional [196]. fragmentary [509]. framework [831, 611, 499, 101, 733, 176, 348, 555]. free [528, 179, 312, 85, 120, 652]. freedom [136]. French [631]. frequency [413, 283]. Frequentist [329]. frontier [743]. Full [765, 426]. fully [639]. fully-automated [639]. FunCC [337]. Function [566, 18, 804, 238, 77, 359, 395, 555, 687]. Function-on-Function [566, 18, 687]. Functional [135, 668, 571, 233, 705, 720, 153, 441, 735, 542, 183, 740, 813, 80, 138, 706, 150, 337, 597, 431, 596, 811, 778, 302, 727, 463, 176, 747, 665, 278, 110, 397, 406, 584, 256, 411, 630, 370]. functions [169, 568, 151, 222, 313, 319, 838, 415, 483, 368, 255, 567, 772]. Fused [8, 810]. Fusing [438]. fusion [539, 787, 348, 620].

Galerkin [318]. game [435]. GARCH [699]. Gaussian [426, 594, 435, 622, 528, 258, 799, 511, 664, 537, 650, 222, 357, 217, 292, 227, 319, 534, 520, 780, 839, 507, 219, 443, 398, 281, 661, 711, 482, 487, 197]. Gaussianity [610]. gene [396]. General [118, 556, 50, 703, 554, 221, 419]. general-purpose [50]. Generalisations [535]. generalised [511, 189, 734]. generalization [624]. Generalized [825, 147, 655, 19, 344, 275, 484, 149, 314, 673, 704, 694, 835, 394, 214, 231, 312, 773, 486, 463, 751, 428, 306, 62, 630, 247, 16, 263, 197]. Generating [33]. generation [836]. genetic [257, 165, 320]. geometric [181]. geometry [752]. georeferenced [516]. Gittins [653]. GLMs [314]. global [40, 512, 616, 832]. GMM [680, 253]. Goodness [762, 834, 215, 239, 556, 628, 613, 75, 67, 55, 699, 380, 419]. Goodness-of-fit [762, 834, 215, 239, 556, 628, 613, 75, 67, 55, 699, 380, 419]. GOS [644]. GP [780]. GP-BART [780]. GPU [523]. gradient [228]. gradients [51]. Graph [816, 375, 174, 27, 326, 281]. Graph-based [816]. Graphical [251, 752, 258, 376, 537, 572, 207, 396, 68, 601]. grids [288]. Group [616, 807, 482, 673, 420, 766]. Grouped [595, 81, 703, 417, 96, 485]. groups [121, 83, 403]. guarantees [824]. guided [381].

habitat [609]. Hamiltonian [368]. handling [309, 611]. Harmless [349]. hazard [847]. hazards [679, 344, 849, 76, 414]. health [631]. heavy [692, 835, 570, 699]. heavy-tailed [692, 835, 570, 699]. Heckman [109]. Hermitian [526]. heterogeneity [346, 565, 720, 647]. Heterogeneous [833, 558, 219, 702, 348, 840]. heteroscedastic [783, 429, 822]. heteroscedasticity [839, 651]. heteroskedastic [86]. heteroskedasticity [674]. hh} [799]. Hidden [316, 684, 451, 506, 496, 39, 192, 294]. Hierarchical [218, 409, 592, 812, 569, 357, 313, 544, 232, 633, 773, 124, 82]. High [295, 543, 162, 6, 567, 728, 205, 758, 617, 156, 44, 602, 144, 466, 706, 237, 207, 72, 367, 656, 331, 312, 177, 148, 524, 812, 803, 70, 417, 459, 751, 120, 178, 837, 714, 559, 127, 507, 159, 625, 510, 219, 386, 241, 81, 330, 513, 721, 433, 14, 436, 806, 277, 621, 382, 429, 604, 785, 32, 723, 718, 791, 445]. high-dimension [177]. High-dimensional [543, 162, 6, 567, 728, 205, 758, 617, 44, 602, 144, 466, 706, 237, 207, 72, 656, 312, 524, 812, 803, 70, 417, 459, 751, 837, 127, 507, 625, 510, 219, 386, 721, 433, 436, 806, 382, 429, 604, 785, 32, 723, 718, 791, 445]. higher [781]. higher-order [781]. Hilbert [150, 178, 436]. HiQR [806]. histograms [639]. HMM [844]. homogeneous [154, 39]. homoskedasticity [785]. homotopic [700]. honest [418]. Horseshoe [539]. hospital [240]. hospital-acquired [240]. hourly [192]. Huber [163, 458]. Huber-type [163]. Hybrid [778, 479, 199, 563]. hydrology [484]. hypercube [616]. hypersphere [254, 471, 478]. Hypothesis [317, 723, 342, 615, 429].

i.i.d [164]. Identification [581, 667, 126, 233, 823, 311]. identify [582]. Identifying [488]. ignorable [94, 137]. II [300, 387, 457]. image [292, 36, 480, 793]. imaging [718]. imbalanced [8]. Impact [546, 257, 535]. impartial [660]. implement [698]. implementation [691]. implication [156]. implications [119]. importance [562, 662, 832]. improper [440]. Improved [200, 529, 733]. Improving [426]. imputation [751, 137]. Imputed [688]. incident [220]. income [96]. incomplete [611, 333, 803]. Independence [531, 655, 134, 439, 721, 310, 491]. independent [398]. index [558, 167, 407, 653, 82, 380, 66, 826, 51, 531, 4, 253, 687].

Indicator [711, 137]. Indicator-based [711]. indices [132]. individual [551]. individualized [817, 95, 168]. induced [297]. inequality [394]. infections [240]. Inference [407, 189, 839, 381, 528, 813, 558, 499, 394, 214, 626, 230, 827, 593, 319, 448, 347, 638, 579, 459, 767, 709, 479, 124, 159, 659, 625, 273, 844, 739, 665, 326, 401, 96, 437, 262, 643, 74, 170, 73, 557, 175, 485, 822]. inferential [119]. inferiority [544]. Inferring [154]. inflated [75]. information [300, 415, 681]. informative [349, 284, 76]. informatively [498]. informed [375]. INGARCH [579]. Ingersoll [106]. inhomogeneous [63]. INLA [659]. insufficient [692]. insulin [587]. integer [760, 456]. integer-valued [760, 456]. Integrated [823, 609, 568, 261]. Integrating [792]. integration [643]. integrative [333]. intensity [63, 834, 87]. inter [230, 800]. inter-arrival [230]. inter-regional [800]. Interaction [488, 409, 389, 727, 652]. interactions [617, 196, 812, 687]. interdirections [495]. interpretable [488]. interpretations [845]. Interval [88, 594, 180, 498, 347, 344, 849, 204, 76, 247, 59]. interval-censored [498, 347, 344, 849, 204, 76, 247, 59]. intervals [553, 301, 71, 445, 460]. intracranial [74]. intractable [309, 715]. intrinsic [526, 357]. Introducing [13]. invariance [420]. invariant [313, 203]. inverse [839, 375, 343, 427, 410, 149]. inversion [714, 460]. invertible [256]. irregular [454]. irregularly [170]. Ising [386]. isometric [746]. issue [706, 669, 670]. items [719, 286]. Iterative [562, 253].

Jackknife [175]. Jackstraw [643]. January [193, 383, 574, 763]. Joint [771, 231, 520, 120, 397, 166, 587, 101, 252, 467, 734, 551, 160, 523, 601, 315, 661, 56]. joint-modeling [101]. Jones [714]. July [114, 307, 474, 689, 842]. jump [358]. jumps [167, 791]. June [107, 289, 464, 671, 828].

K-bMOM [416]. Kalman [381]. Kendall [269]. Kernel [451, 46, 63, 150, 5, 345, 336, 359, 178, 436, 149]. Kernel-based [451, 336, 149]. kernels [536]. kinetic [709]. knockoff [503]. Kolmogorov [846, 712]. Kriging [276]. Kronecker [288]. Kullback [452, 332].

label [349]. labelling [299]. labels [349]. lacking [223]. Lagrangian [402]. Laplace [261, 214, 328, 347, 815, 62, 783, 197]. Laplace-based [815]. Laplacian [747, 601]. Large [174, 525, 362, 288, 102, 770, 333, 830, 112, 639, 477, 518, 250, 648, 731]. Large-scale [174, 525, 770, 333, 830, 639, 518, 250, 648]. lasso [104, 572, 804, 199, 16, 766, 566, 132, 513, 700]. lasso-type [199]. Latent [326, 582, 673, 576, 40, 825, 571, 132, 352, 614, 461, 607, 166, 215, 414]. Latin [616]. lattice [666]. ‘LatticeKrig’ [516]. LDA [756]. LEAN [557]. learner [173]. Learning [219, 413, 293, 664, 333, 287, 656, 304, 792, 577, 608, 391, 95, 529, 848, 68, 6, 256]. Least [147, 89, 378, 701, 121, 77, 750, 41]. least-squares [378]. Lehmann [23]. Leibler [452, 332]. leptokurtosis [674]. leukemia [587]. level [641, 196, 324, 304, 513]. Lévy [836]. life [28]. lift [495]. lift-interdirections [495]. Likelihood [769, 528, 362, 425, 104, 710, 831, 799, 40, 161, 142, 444, 787, 204, 479, 686, 98, 404, 737, 732, 57, 527, 395, 66, 826, 29, 783, 73, 557, 175]. likelihood-based [425]. Likelihood-free [528]. Likelihood-type [769]. likelihoods [151]. line [524]. linear [673, 758, 183, 598, 543, 817, 704, 248, 566, 191, 80, 680, 664, 237, 150, 481, 346, 835, 394, 312, 371, 773, 340, 693, 402, 177, 302, 19, 463, 291, 784, 529, 497, 219, 818, 665, 53, 241, 81, 401, 584, 62, 66, 826, 128, 153, 7, 253, 429, 604, 482, 791, 263]. linearization [738]. Linearly [228]. Link [240, 453, 554, 395]. Link-based [240, 554]. linkage [492, 631]. links [379]. Lloyd [416]. Lloyd-type [416]. Local [665, 512, 395, 309, 174, 201, 525, 273, 79, 832]. Locally [650, 727, 534]. location [72, 387, 347, 493]. location-scale [347]. location-shifts [387]. Log [515, 650, 746]. log-Gaussian [650]. log-ratio [746]. Log-regularly [515]. logarithm [395]. Logistic [101, 609, 34, 811, 573, 455, 578]. logspline [195]. long [587]. long-term [587]. Longitudinal [437, 413, 833, 231, 599, 254, 103, 467, 737, 615, 160, 471, 720, 284, 100, 557, 661, 768, 454]. longitudinally [568]. Lorenz [407]. loss [846, 229, 14, 567]. Low [452, 261]. Low-rank [452]. lower [844]. lymphoblastic [587].

machine [792, 750, 747]. machines [505]. Mallows [84, 509]. manifolds [532]. Mann [406, 145]. MANOVA [602, 429]. map [686]. mapping [609, 276]. mapping-based [276]. March [37, 224, 399, 605, 789]. Marginal [334, 508, 262]. marginalisation [261]. Marginally [161]. marked [251]. Markov [292, 294, 684, 562, 451, 545, 637, 444, 506, 316, 590, 6, 496, 39, 192]. Markov-switching [545]. martingale [655]. massive [799, 353, 327, 8, 745, 71]. matching [435, 207, 585, 118, 311]. matrices [288, 526, 144, 819, 33, 315]. matrix [452, 526, 371, 629, 415, 241, 133, 17, 187, 14, 621, 745, 742, 604, 723, 766]. matrix-valued [629]. matrix-variate [371, 187]. MAVE [4, 335]. Max [721]. Max-sum [721]. maxima [309]. maximization [846]. Maximum [444, 710, 40, 161, 204, 479, 686, 57, 29]. May [90, 259, 446, 657, 820]. MCMC [426, 456]. Mean [624, 179, 367, 264, 162, 520, 336, 655, 428, 28, 645, 510, 306, 366, 795, 7, 785, 310]. mean-mixture [264]. mean-variance [645]. mean-variance-correlation [428]. means [416, 407, 136, 233, 839, 314, 382]. Measure [546, 336, 559, 443]. measurement [60, 494, 258, 131, 279, 7, 557]. measurements [3]. measures [682, 113, 662, 300]. mechanism [684]. mechanisms [432]. media [630]. median [416, 297]. median-of-means [416]. Mediation [718, 543, 587, 32, 618]. mediator [587, 618]. mediator-outcome [618]. mediators [718]. medical [177]. members [305]. membership [767, 837]. meta [200, 619]. meta-analysis [200]. meta-regression [619]. method [594, 261, 216, 556, 227, 585, 389, 318, 131, 245, 30, 713, 714, 439, 39, 810, 783, 640]. methods [582, 539, 34, 44, 720, 149]. microbial [633, 667]. microsimulations [223]. MINAR [757]. minimizing [609]. minimum [650, 477]. minimum-volume [477]. misalignment [337]. Mises [710]. Missing [453, 831, 817, 684, 282, 379, 101, 30, 202, 751, 94, 26, 351, 433, 518, 235, 332, 437, 410, 66, 137, 73, 681]. misspecification [148]. MIXANDMIX [2]. mixed [564, 585, 10, 402, 75, 751, 767, 837, 240, 404, 665, 401, 61, 631, 62, 619, 262, 128, 329]. mixed-effects [619, 262, 128, 329]. mixed-membership [767]. mixed-type [564, 751, 631]. Mixing [155]. Mixture [612, 291, 586, 414, 542, 762, 710, 119, 536, 515, 64, 264, 360, 608, 325, 298, 794, 344, 765, 645, 775, 45, 530, 247, 840, 424]. Mixture-based [586]. mixtures [2, 47, 268, 685, 128]. MLE [156]. MM [226]. modal [279, 327, 51]. mode [825, 186]. Model [80, 179, 182, 190, 402, 85, 61, 299, 398, 542, 60, 758, 617, 34, 817, 293, 566, 296, 684, 741, 466, 680, 664, 237, 69, 701, 825, 558, 257, 394, 571, 324, 331, 64, 633, 136, 595, 86, 608, 101, 371, 744, 533, 624, 713, 334, 727, 84, 103, 70, 25, 849, 767, 146, 847, 120, 189, 178, 28, 233, 291, 784, 480, 386, 591, 81, 45, 420, 229, 678, 461, 192, 584, 187, 815, 306, 631, 3, 76, 221, 523, 395, 674, 332, 519, 277, 276, 652, 826, 153, 414, 786, 88, 478, 509, 74, 661, 109, 583, 749, 15, 687, 350]. Model-based [61, 398, 293, 187, 478]. Model-free [179, 85, 120, 652]. Modeling [106, 579, 757, 20, 295, 587, 409, 760, 537, 525, 42, 121, 830, 47, 165, 792, 132, 101, 726, 252, 794, 645, 507, 516, 480, 152, 471, 17, 512, 768]. Modelling [98, 771, 568, 27, 251]. models [169, 673, 609, 494, 174, 265, 752, 425, 156, 576, 183, 740, 598, 258, 543, 704, 376, 208, 80, 741, 426, 845, 119, 694, 545, 770, 481, 701, 40, 161, 692, 207, 190, 209, 357, 43, 835, 637, 550, 99, 227, 214, 781, 167, 97, 612, 312, 360, 22, 773, 358, 10, 211, 396, 93, 755, 402, 448, 347, 188, 298, 715, 95, 302, 390, 75, 344, 202, 463, 506, 751, 654, 709, 189, 41, 204, 837, 479, 240, 124, 765, 467, 734, 625, 404, 497, 68, 552, 219, 386, 775, 273, 739, 554, 160, 665, 496, 53, 500, 110, 401, 473, 17]. models [743, 850, 57, 841, 39, 530, 666, 18, 607, 62, 3, 348, 699, 215, 795, 423, 832, 262, 731, 814, 82, 380, 247, 66, 128, 456, 805, 840, 600, 683, 4, 7, 100, 253, 711, 454, 791, 329, 263, 294]. modified [756, 471]. modularity [266]. modulated [648]. moment [343]. moments [622]. monitoring [580]. monotone [155, 407, 397, 168]. Monte [556, 562, 160, 79, 39, 52, 368, 640]. motif [356]. Motor [52]. movements [534]. moving [553, 480, 62]. Multi [112, 848, 810, 333, 726, 653, 204, 634, 823, 649, 809, 632]. multi-arm [632]. multi-armed [653]. multi-aspect [649]. Multi-block [810]. multi-dimensional [809]. multi-output [634]. multi-response [333]. Multi-scale [112]. multi-source [726, 823]. multi-state [204]. Multi-task [848]. multi-way [726]. multicategory [505, 244]. Multiclass [455, 747, 592]. Multiclass-penalized [455]. multicollinearity [517]. multidomain [514]. multilayer [27, 461]. Multilevel [618, 296, 98]. Multimodal [368, 181, 718]. multinomial [156, 578, 620]. multiple [673, 553, 540, 95, 459, 146, 127, 403, 759, 472, 235, 648, 277, 768]. multiple-class [95]. multiple-trial [540]. Multiplicative [5, 211, 795, 7, 350]. multiplier [602, 93]. multipliers [810]. multiply [611]. multiresolution [516]. multiscale [623]. multitype [251]. Multivariate [481, 495, 743, 601, 441, 757, 542, 20, 504, 556, 470, 69, 251, 33, 499, 571, 431, 585, 593, 755, 359, 254, 352, 579, 390, 716, 508, 467, 734, 160, 493, 278, 471, 676, 634, 667, 666, 283, 250, 619, 620]. multivariate-marked [251].

national [631]. near [688]. near-sited [688]. Nearest [799, 426]. Neighbor [426]. Neighborhood [728]. Neighborhood-based [728]. neighbors [799]. nested [261, 64, 146, 512]. net [693]. Network [485, 580, 9, 216, 396, 476, 448, 249, 85, 54, 461, 607, 256, 74, 281, 807, 620]. network-valued [85]. networks [27, 296, 9, 770, 42, 825, 379, 319, 438, 744, 36, 188, 715, 469, 6, 356, 473, 731, 824]. Neural [715, 216, 438, 36, 469, 807, 824]. neuroimaging [545, 482]. NN [432]. NN-walktrap [432]. noise [349]. noisy [596, 368]. nominal [513]. nomination [105]. Non [430, 435, 154, 580, 799, 208, 222, 270, 544, 571, 536, 64, 164, 520, 117, 189, 581, 124, 94, 507, 513, 39, 256, 361, 137, 482, 822, 197]. non- [799]. non-asymptotic [513]. non-conjugate [189]. non-convex [117]. non-Gaussian [435, 507, 482, 197]. non-homogeneous [154, 39]. non-i.i.d [164]. non-ignorable [94, 137]. non-inferiority [544]. non-invertible [256]. non-mixture [64]. Non-parametric [430, 580, 208, 270, 571, 536, 520]. non-randomly [361]. non-stationary [222, 581, 124, 822]. non-uniformly [361]. nonconvex [185, 266, 335]. nonignorable [434, 282, 533, 202, 255, 73]. Nonlinear [541, 10, 755, 404, 678, 832, 143, 228]. nonnegative [345]. Nonparametric [582, 568, 692, 626, 83, 798, 318, 469, 793, 170, 156, 217, 238, 387, 679, 347, 30, 794, 202, 134, 765, 615, 184, 250, 436, 685, 21, 788]. nonparametrics [792]. nonresponse [434, 255]. norm [805]. Normal [268, 60, 762, 69, 43, 264, 291, 645, 127, 625, 181, 478, 429]. normality [550, 271]. normals [515]. Novel [602, 261, 761, 318, 780, 291, 53]. November [157, 363, 548, 729]. Number [46, 155, 298, 836, 189, 52, 683, 263]. numbering [13]. numerical [741, 2, 245, 6].

obesity [587]. objective [440]. observation [284]. observational [403, 702, 370]. observations [66, 485]. observed [673, 596, 381, 705, 31]. October [139, 354, 521, 724]. odd [641]. odds [247]. One [796, 825, 712, 361, 809, 429]. one-dimensional [809]. one-mode [825]. one-sided [712]. one-step [361]. one-way [429]. ones [634]. Online [580, 608, 131, 751, 722, 742, 750]. OofA [424]. operational [580]. operator [581, 795]. Optimal [99, 304, 600, 385, 769, 50, 663, 818, 110, 424, 518, 286]. optimising [565]. optimization [50, 117, 266, 199]. optimizer [40]. Oracle [411, 822]. Oracle-efficient [411, 822]. Order [683, 834, 342, 779, 781, 839, 547, 276, 385, 56]. order-of-addition [276, 385]. ordered [514, 540, 450, 377]. ordering [55, 281]. orders [15]. ordinal [180, 132, 484, 430]. ordinary [340, 654, 507]. organizations [296]. organizing [686]. Ornstein [154]. orthogonal [691, 476]. other [772]. outbreak [314]. outcome [587, 304, 618]. outcomes [333, 498, 554, 437, 718]. Outer [232, 51]. Outlier [542, 379, 135, 236, 629]. output [634]. outputs [741]. overdispersion [470].

P [214]. P-splines [214]. pages [512]. paired [83, 497]. pairs [799, 218]. Pairwise [845, 292, 637, 6]. pandemic [453]. panel [190, 209, 395, 519]. Parallel [333, 227, 577, 677]. Parallel-and-stream [577]. Parameter [836, 415, 478, 248, 268, 695, 101, 715, 31]. parameterizations [637]. parameterized [161]. Parameters [46, 34, 236, 703, 445]. parametric [580, 494, 208, 216, 270, 571, 536, 774, 520, 732, 279, 483, 430, 600]. parsimonious [17]. Partial [736, 20, 543, 251, 207, 121, 233, 584, 395, 7]. Partially [150, 514, 758, 598, 817, 680, 83, 744, 533, 727, 178, 784, 94, 381, 705, 584, 66, 826, 253, 791]. Particle [341, 262]. partite [9]. Partition [285]. Partition-based [285]. partitions [492]. partly [253]. path [125]. paths [16]. pathway [459]. patients [713]. pattern [280, 792, 491]. patterns [342, 484, 430]. Pearson [180]. Penalised [204]. Penalized [195, 804, 830, 695, 334, 846, 19, 455, 404, 65, 527, 14, 805, 71, 335]. penalties [806]. penalty [532, 195, 325, 782]. performance [580, 113, 656, 719]. performances [426]. periodograms [328]. permutation [662]. persistence [148]. perturbation [487]. phase [457, 415, 675, 300]. piece [442, 60]. Pietra [175]. plans [110]. plant [633]. point [63, 834, 677, 251, 405, 342, 792, 796, 555, 746, 824]. points [553]. Poisson [42, 22, 638, 624, 75, 67, 87, 88]. Poisson-stopped-sum [42]. political [22]. Polya [640]. Polymatching [403]. polynomial [201, 77]. polynomials [597]. polytomous [573]. pooled [784]. population [102, 2, 703, 840]. populations [34, 412, 839]. portmanteau [390]. positive [526]. possibly [237]. Post [827]. Post-clustering [827]. Posterior [124, 43]. Potts [748]. Power [387, 819, 43, 232, 513]. power-expected-posterior [43]. powerful [55, 382]. practical [546, 827]. practice [220]. precision [714, 241, 14, 621, 315, 766]. preconditioned [228]. Predicting [552, 153]. Prediction [222, 596, 594, 735, 101, 529, 523, 509]. predictions [713]. Predictive [719, 541, 656, 726, 717]. predictors [169, 324, 132, 702]. prepivots [367]. presence [293, 538, 76, 649, 651]. preserving [325]. Primal [125, 830]. primal-dual [830]. Principal [750, 283, 396, 163, 629, 176, 705, 397, 678, 32]. prior [492, 546]. priors [43, 387, 440]. probabilities [223]. Probability [775, 412, 798, 713, 88]. problem [89, 145, 184, 235]. problems [72, 206, 95, 653, 518, 199]. procedure [544, 210, 535]. procedures [553, 83, 698]. process [757, 154, 644, 222, 227, 626, 534, 608, 106, 838, 717, 590, 705, 356, 87, 348, 221, 555, 29, 711, 746, 426]. process-based [348]. processes [594, 528, 63, 834, 80, 409, 650, 251, 520, 638, 836, 780, 31, 284, 443, 197]. product [51]. products [288]. profile [352]. profiled [758]. programming [818, 433]. progression [352]. projected [80]. Projection [417, 646, 380, 267, 127, 738, 297]. Projection-averaging-based [380]. Promote [315]. Propagation [217]. propensity [533, 370]. properties [786]. proportion [714]. proportional [679, 849]. proportionality [144, 847]. pseudo [262]. pseudo-marginal [262]. psychological [719]. purpose [50]. pursuit [325, 738]. PX [228]. PX-EM [228].

Q [476]. Q-network [476]. quadratic [740, 319, 54, 18, 14, 806]. quadrature [576]. qualitative [488]. quantification [47, 36, 418]. quantifying [546]. Quantile [78, 674, 24, 441, 603, 216, 92, 405, 272, 331, 450, 626, 570, 833, 391, 599, 58, 688, 727, 103, 19, 417, 316, 508, 759, 818, 739, 483, 433, 722, 65, 353, 436, 793, 810, 745, 71, 73, 749, 687, 681]. quantile-based [272, 739]. quantiles [317, 118, 255]. quantitative [488]. quantities [716]. quantum [294]. quasi [151, 141, 783]. quasi-likelihood [783]. quasi-likelihoods [151]. quasi-sudoku [141].

radiation [587]. rainfall [192]. Raman [47]. random [174, 504, 799, 27, 357, 33, 662, 468, 836, 134, 469, 127, 26, 415, 356, 330, 850, 841, 62, 523, 332, 600, 661, 487]. randomisation [535]. randomization [590, 632]. randomized [495, 271, 65]. randomly [361, 153]. Rank [138, 209, 452, 72, 313, 22, 325, 202, 510, 698, 168]. Rank-based [209, 510, 168]. rank-invariant [313]. rank-preserving [325]. ranked [94]. ranks [495]. rate [106, 334, 812, 667, 247]. rates [387, 245]. ratio [845, 175, 746, 111]. ratios [365]. Rayleigh [480]. record [492, 631]. recovery [162, 534, 185, 524]. Recursive [779]. reduced [22, 255]. reduction [514, 102, 434, 182, 750, 438, 252, 563, 646, 343, 226, 410, 149, 24, 651]. reference [429]. regenerative [295]. regime [277]. regimes [304, 168]. region [809]. Regional [759, 223, 800, 317]. regions [769, 200, 506]. registration [368]. Regression [169, 282, 284, 59, 609, 494, 239, 34, 183, 740, 295, 566, 526, 191, 80, 78, 378, 150, 845, 69, 236, 481, 179, 333, 92, 346, 217, 116, 43, 597, 272, 779, 238, 331, 515, 97, 626, 407, 596, 833, 534, 132, 360, 391, 101, 206, 811, 599, 693, 298, 302, 321, 252, 624, 787, 688, 269, 103, 19, 25, 463, 458, 780, 748, 316, 508, 573, 275, 388, 645, 455, 176, 625, 529, 848, 759, 818, 375, 53, 110, 634, 279, 646, 675, 483, 667, 18, 722, 738, 65, 578, 306, 527, 221, 250, 353, 327, 567, 436, 806, 793, 619, 427, 423, 832]. regression [418, 810, 410, 149, 24, 153, 51, 745, 742, 840, 73, 335, 16, 408, 234, 687, 681, 297]. regressions [104, 664, 774, 788]. regressors [253]. Regular [491, 196, 15]. Regularization [151, 258, 627, 94, 621, 583]. Regularized [466, 147, 675, 831, 664, 185, 117, 837, 742, 749]. regularly [515]. relational [246]. relationships [256]. renewable [722]. Reparameterization [733]. replicate [540, 3]. replicate-effect [540]. replicates [676]. reported [757]. representation [504, 288, 838]. reproducing [150, 178, 436]. resampling [412]. residual [603, 550, 211, 28]. residual-based [211]. resistance [587]. Response [300, 817, 222, 333, 30, 252, 590, 433, 235, 299, 826, 600, 137, 73]. response-adaptive [590]. responses [459, 759, 535, 531, 408]. restricted [839]. reversed [679]. revisited [35, 5]. ridge [104, 779, 312, 111]. right [201, 391, 347, 503]. right- [347]. right-censored [503]. RIHT [785]. risk [452, 587, 165, 304, 847, 775, 88, 768]. risks [245, 146, 120, 240]. RKHS [285]. road [295]. robot [305]. robotics [305]. Robust [594, 34, 183, 740, 293, 92, 835, 328, 206, 811, 693, 610, 629, 623, 573, 388, 645, 471, 31, 229, 483, 530, 449, 353, 327, 299, 335, 598, 416, 611, 562, 236, 116, 660, 515, 304, 148, 818, 493, 235, 221, 567, 437, 619]. Robustness [271, 299]. ROC [270, 23, 538]. role [587]. Ross [106]. rotation [676]. Rothman [35]. roughness [532]. ruin [88]. rule [524]. rules [505, 199].

SAEM [10]. Safe [505, 199]. same [786]. same-step [786]. sample [362, 205, 145, 505, 72, 367, 162, 468, 142, 387, 679, 85, 278, 330, 423, 814, 382, 604]. samples [186]. Sampling [133, 309, 562, 232, 614, 28, 98, 110, 796, 456, 105, 198, 15]. sampling-based [198]. SAR [480]. Saunders [345]. SCAD [766]. scalable [357, 227]. scalar [793]. scalar-on-image [793]. scale [174, 525, 770, 236, 333, 830, 515, 203, 112, 347, 639, 291, 518, 250, 648, 731]. scale-invariant [203]. scale-mixture [291]. scan [280, 301, 406]. scatter [610, 493]. scheme [579, 554]. schemes [105]. science [835]. score [207, 804, 533, 370, 768]. screening [622, 505, 736, 30, 655, 417, 120, 285, 26, 503, 81, 500, 841, 8, 652, 531, 137, 310]. search [112, 524]. second [779, 547]. second-order [779, 547]. sectional [209, 370]. seemingly [481]. segmentation [292, 816, 36, 623]. selection [201, 831, 44, 704, 293, 376, 237, 179, 498, 257, 304, 136, 695, 101, 693, 318, 177, 298, 808, 803, 305, 846, 469, 439, 404, 386, 563, 53, 591, 330, 841, 229, 39, 449, 353, 332, 427, 423, 226, 247, 73, 198, 711, 807, 782, 109, 15]. Selective [401]. self [686, 241, 456]. self-calibrated [241]. self-exciting [456]. self-organizing [686]. Semi [758, 183, 216, 120, 316, 273, 496, 732, 436, 519, 600, 454]. semi-competing [120]. semi-functional [183]. semi-Markov [496]. semi-Markov-switching [316]. semi-nonparametric [436]. semi-parametric [216, 732, 600]. Semi-profiled [758]. semi-varying [273, 519, 454]. semicontinuous [166]. Semiparametric [191, 378, 272, 64, 713, 25, 28, 152, 850, 687, 350, 331, 167, 86, 93, 847, 676, 420, 57, 255, 772]. sensitive [244]. sensitivity [769, 616]. separability [819, 342, 610]. separable [744]. separate [634]. Separating [117]. Separation [217]. September [129, 338, 501, 707]. sequence [582, 42, 306]. sequences [261]. Sequential [607, 840, 135, 637, 52]. serial [430]. Series [669, 757, 735, 20, 295, 760, 572, 571, 570, 230, 593, 328, 579, 808, 316, 581, 623, 233, 472, 192, 283, 430, 795, 456, 822, 651]. set [365]. sets [102, 677, 450, 639]. setting [164]. settings [656]. several [785]. Shape [597, 365, 420]. Shape-constrained [597, 365]. shaped [374]. shapes [169, 126]. share [845]. Shared [578, 523]. shifts [387, 277]. shotgun [112]. shrinkage [539, 578, 601]. shrinking [686]. sided [712]. sign [315]. signal [210]. signatures [287]. signed [698]. signed-rank [698]. Significance [772]. Silhouette [303]. SIMEX [279]. simple [127, 832]. Simplified [234]. simulating [808, 356]. simulation [525]. Simultaneous [813, 809, 238, 411]. Single [51, 558, 167, 407, 472, 380, 66, 826, 4, 253, 687]. Single-index [51, 558, 407, 380, 66, 826, 253, 687]. singular [517]. sited [688]. sites [512]. size [362, 102, 509]. sJIVE [551]. skew [164]. skew-symmetric [164]. skewed [271, 181]. skewness [113, 674]. slice [614]. sliced [628, 375, 410]. Small [703, 92, 702]. SMC [40]. Smirnov [846, 712]. SMLSOM [686]. Smooth [566, 238, 97, 722]. Smoothed [599, 73, 626]. smoothing [201, 555, 71]. Smoothly [627]. social [42, 630]. sociology [296]. soft [349]. soft-labels [349]. Solution [16, 433, 366]. Some [67, 164]. source [726, 823]. space [208, 545, 434, 825, 99, 755, 178, 461, 607, 436, 809]. space-time [99]. spaced [170]. spaces [150]. span [844]. Sparse [185, 176, 436, 32, 583, 617, 572, 633, 320, 778, 727, 275, 124, 507, 513, 366, 427, 410, 398, 601, 807, 445]. sparseness [325]. sparsity [798]. Spatial [457, 647, 362, 609, 265, 63, 223, 799, 288, 760, 680, 628, 72, 357, 280, 99, 816, 595, 112, 688, 41, 301, 406, 229, 96, 62, 348, 197]. spatial-temporal [688]. spatially [796, 16, 583]. Spatio [694, 409, 650, 761, 251, 161, 342, 473]. Spatio-temporal [694, 409, 650, 761, 251, 161, 342, 473]. Spearman [721]. special [706, 669, 670]. species [155]. specific [847]. Specification [93, 654]. specificity [769]. specified [533]. spectral [526, 2, 770, 781, 593, 581, 105, 170, 604]. spectrometric [537]. spectroscopy [47]. spectrum [216, 604]. spells [146]. spheres [568]. spherical [685, 478]. spiked [683]. spiked-type [683]. spikes [683]. spline [598, 318]. splines [214, 71]. Split [142]. square [750]. squared [229]. Squares [147, 89, 378, 701, 121, 77, 41]. squares-type [77]. stabilizations [613]. stable [248, 324]. stage [457, 518, 700]. Standard [773]. state [208, 545, 755, 294, 204, 496]. state-space [545]. States [761, 314]. stationary [222, 550, 593, 534, 808, 581, 124, 822]. statistic [35, 301, 406, 785]. Statistical [435, 558, 230, 459, 443, 746, 682, 564, 585, 805]. statisticians [432]. Statistics [13, 257, 546, 714, 712, 547, 604, 56]. step [361, 786]. stepwise [161]. Stochastic [504, 154, 309, 296, 525, 770, 779, 781, 10, 112, 767, 709, 189, 838, 479, 676, 743, 55, 262, 731, 814]. stopped [42]. strategies [653, 565]. strategy [78]. stratum [691]. stream [577, 473]. streaming [742]. streams [131]. strength [641, 82]. strengths [305]. strong [691, 199]. strongly [491]. Structural [68, 543, 664, 121, 219, 772]. structure [320, 6, 454, 723]. Structured [70, 89, 819, 159, 618]. structures [582, 119, 833, 601, 788]. student [836]. studies [498, 64, 75, 403, 702, 675, 299, 276, 370]. study [34, 413, 6, 557, 437]. sub [664]. sub-Gaussian [664]. subdata [198]. Subgroup [817, 311, 833, 360, 823]. subject [69, 460]. Subsampling [704, 770, 518]. subscales [719]. subset [280, 449]. subspace [366]. substance [352]. sudoku [141]. sufficient [182, 438, 736, 563, 646, 343, 427, 226, 149, 531]. Sum [288, 553, 42, 721]. Sunter [631]. Superiority [156]. Supervised [551, 349, 577]. Support [280, 527, 505, 162, 750, 524, 747]. supports [408]. sure [310]. surface [538]. surfaces [526, 143]. Surrogate [434, 525, 33, 499, 846]. survey [223]. survival [642, 69, 453, 713, 794, 334, 812, 240, 748, 467, 500, 783, 768, 310, 681]. SUTSE [786]. switching [545, 316]. symmetric [191, 164, 476, 616, 723]. symmetrical [616]. symmetry [35, 164, 698]. synchrosqueezing [534]. synthetic [732]. system [631]. Systematic [641]. systems [514, 248].

tables [460]. tail [771, 716, 815]. tailed [692, 835, 570, 699]. task [371, 848, 152]. task-based [371, 152]. taxonomy [135]. techniques [2, 381, 472]. temporal [409, 650, 694, 761, 251, 161, 342, 688, 473, 720]. temporally [607, 485]. tenets [739]. tensor [599, 688]. term [587]. Test [460, 769, 239, 834, 511, 813, 684, 270, 72, 819, 35, 367, 162, 200, 679, 75, 85, 134, 127, 439, 510, 330, 721, 31, 443, 699, 215, 814, 382, 429, 772]. Test-inversion [460]. Testing [144, 346, 342, 774, 306, 762, 34, 9, 556, 677, 540, 544, 165, 662, 468, 827, 654, 581, 317, 615, 648, 380, 429, 785, 723]. Tests [319, 86, 205, 602, 628, 209, 613, 164, 328, 387, 610, 93, 148, 390, 719, 67, 278, 513, 721, 286, 55, 649, 419]. text [22, 512]. their [484]. theoretic [435, 535]. theoretical [786, 824]. theory [156]. therapy [587]. thoracic [534]. three [544, 674]. three-factor [674]. threshold [277, 456, 29]. threshold-type [277]. Thresholding [513, 111]. ties [484, 649]. Time [526, 572, 324, 22, 757, 735, 154, 20, 413, 295, 760, 466, 701, 99, 571, 246, 570, 230, 593, 328, 444, 579, 808, 838, 316, 581, 623, 233, 152, 554, 160, 496, 850, 472, 675, 192, 666, 283, 76, 430, 795, 247, 153, 456, 59, 661, 822, 766, 651, 669]. time-average [623]. Time-dependent [22, 59]. time-frequency [413]. time-series [20, 295]. time-to-event [554, 160, 661]. Time-varying [526, 757, 246, 152, 666, 766]. times [230, 439, 552]. TMB [329]. tool [546, 750, 439]. topics [512]. Topology [628, 9]. Topology-based [628]. total [195]. traffic [295]. trait [106]. trans [804]. trans-lasso [804]. transaction [326]. transfer [529]. transform [534, 427]. transformation [86, 93, 334, 202, 57, 746]. Transformations [732, 413, 135, 232]. transformed [699]. transition [223]. treatment [728, 587, 558, 633, 304, 403, 805, 168, 772]. treatments [377, 370]. Tree [634, 389, 75, 374, 311, 640]. Tree-based [634]. tree-shaped [374]. trees [431, 780, 418]. trend [797, 822]. trends [413]. trial [540, 565, 681]. trials [541, 544, 300, 457, 271, 590, 141, 632]. triangulation [173]. trimming [660]. triplet [637]. Truncated [463, 624, 57]. truncation [439]. Tukey [799]. Tukey- [799]. Tuning [695, 312]. Tuning-free [312]. Two [205, 367, 679, 278, 330, 187, 814, 735, 442, 60, 532, 34, 144, 641, 145, 825, 72, 196, 83, 162, 468, 274, 387, 457, 85, 133, 675, 518, 601, 168, 382]. two-dimensional [735, 532]. two-factor [196]. two-level [641]. two-mode [825]. two-piece [442, 60]. Two-sample [367, 679, 278, 330, 814, 145, 72, 162, 468, 387, 85, 382]. two-stage [457, 518]. type [769, 154, 34, 416, 564, 77, 345, 387, 163, 751, 712, 415, 61, 698, 631, 277, 199, 683, 109]. types [69].

Uhlenbeck [154]. ultra [331, 120, 81, 621]. ultra-high [331, 120, 81, 621]. ultrahigh [30, 26, 503, 500, 810, 310]. ultrahigh-dimensional [30, 503, 500]. ultrasonic [126]. unbiased [452]. Uncertainty [36, 418, 765, 591]. Under-reported [757]. unequal [412]. Unified [652, 831, 611, 555]. uniform [813, 55]. uniformly [361]. unifying [691, 348]. unit [324, 52, 478]. unit-level [324]. United [761, 314]. univariate [634]. universal [276]. unknown [298, 189, 698, 395, 281, 408]. unmeasured [642, 618]. unrelated [481]. unspecified [850]. Unsupervised [710, 292, 637, 431]. updating [131]. upgraded [361]. urban [87]. use [468, 352, 417, 719, 380]. users [630]. using [169, 195, 831, 452, 504, 261, 642, 216, 222, 272, 779, 656, 431, 47, 304, 534, 77, 360, 106, 36, 347, 254, 787, 780, 734, 127, 404, 775, 702, 181, 471, 420, 530, 283, 368, 306, 430, 262, 555, 685, 782].

Valid [827]. validation

[594, 642, 227, 571, 592, 79]. validity [86]. value [751, 733]. valued [760, 744, 85, 629, 456]. values [33, 712]. VAR [572, 68]. Variable [622, 237, 498, 298, 803, 846, 500, 247, 673, 617, 831, 576, 704, 293, 179, 40, 662, 304, 736, 693, 177, 655, 417, 53, 81, 229, 39, 578, 353, 427, 832, 226, 8, 531, 73, 711, 807]. Variables [782, 673, 180, 504, 89, 313, 97, 117, 415, 3, 166, 414]. variance [609, 268, 774, 428, 565, 623, 645, 747, 329, 310]. variances [219]. variate [371, 187]. variation [195, 551]. Variational [448, 849, 767, 159, 21]. Variations [43]. Varying [188, 506, 757, 526, 190, 246, 515, 97, 627, 103, 178, 152, 317, 273, 500, 666, 3, 519, 100, 16, 454, 749, 766]. Varying-coefficient [188, 506, 103, 3, 100]. Vecchia [197]. vector [350, 505, 280, 633, 750, 390, 747, 527, 445]. vectors [162, 134, 510, 517, 785]. versions [835]. Vertex [105]. via [710, 526, 664, 217, 756, 536, 592, 695, 608, 693, 402, 95, 147, 185, 751, 285, 581, 717, 714, 848, 737, 273, 676, 79, 397, 738, 52, 14, 652, 51, 632, 266, 71, 198, 311, 807, 700, 620, 583, 746, 824]. Vine [564, 20, 552, 234]. vines [15, 491]. Visualization [591]. vocalization [126]. volatility [20]. volume [477, 538].

Wald [34]. Wald-type [34]. walktrap [432]. warm [700]. warm-up [700]. warping [838]. Wasserstein [546]. wastewater [633]. Wavelet [540, 526]. waves [536]. way [726, 429]. weak [390]. weakly [608, 186]. web [512]. weight [119]. Weighted [701, 103, 202, 799, 27, 650, 481, 804, 781, 304, 750, 79, 530]. weighting [798]. weights [612]. Whitney [145, 406]. Width [303]. wilcox.test [698]. Wilcoxon [406]. wild [148]. WIM [546]. wise [576, 830, 96]. Wishart [717]. within [540, 101]. Woodroofe [35]. workflow [733].

X [165]. X-chromosome [165].

Yanai [306]. youth [352].

zero [624, 75, 506, 127]. zero-effect [506]. zero-inflated [75]. zero-truncated [624]. zeroes [581]. Zipf [42].

References

Anonymous:2020:EBa

[1] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 141(??):ii–iv, January 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301793.

Cordero-Grande:2020:MNT

[2] Lucilio Cordero-Grande. MIXANDMIX: numerical techniques for the computation of empirical spectral distributions of population mixtures. Computational Statistics & Data Analysis, 141(??):1–11, January 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301458.

Wang:2020:DDM

[3] Miaomiao Wang, Chunling Liu, Tianfa Xie, and Zhihua Sun. Data-driven model checking for errors-in-variables varying-coefficient models with replicate measurements. Computational Statistics & Data Analysis, 141(??): 12–27, January 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301379.

Zhang:2020:BCM

[4] Hong-Fan Zhang, Lei Huang, and Lian-Lian Liu. On bootstrap consistency of MAVE for single index models. Computational Statistics & Data Analysis, 141(??):28–39, January 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301367.

Igarashi:2020:MBC

[5] Gaku Igarashi and Yoshihide Kakizawa. Multiplicative bias correction for asymmetric kernel density estimators revisited. Computational Statistics & Data Analysis, 141(??): 40–61, January 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301446.

Pensar:2020:HDS

[6] Johan Pensar, Yingying Xu, Santeri Puranen, Maiju Pesonen, Yoshiyuki Kabashima, and Jukka Corander. Highdimensional structure learning of binary pairwise Markov networks: a comparative numerical study. Computational Statistics & Data Analysis, 141 (??):62–76, January 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930146X.

Zhang:2020:CAM

[7] Jun Zhang, Bingqing Lin, and Zhenghui Feng. Conditional absolute mean calibration for partial linear multiplicative distortion measurement errors models. Computational Statistics & Data Analysis, 141(??):77– 93, January 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301434.

Xie:2020:FVS

[8] Jinhan Xie, Meiling Hao, Wenxin Liu, and Yuanyuan Lin. Fused variable screening for massive imbalanced data. Computational Statistics & Data Analysis, 141(??):94– 108, January 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301471.

Chen:2020:DTA

[9] Shuo Chen, F. DuBois Bowman, and Yishi Xing. Detecting and testing altered brain connectivity networks with k-partite network topology. Computational Statistics & Data Analysis, 141 (??):109–122, January 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301410.

Karimi:2020:FSF

[10] Belhal Karimi, Marc Lavielle, and Eric Moulines. f-SAEM: a fast stochastic approximation of the EM algorithm for nonlinear mixed effects models. Computational Statistics & Data Analysis, 141 (??):123–138, January 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301483.

Anonymous:2020:F

[11] Anonymous. February 2020. Computational Statistics & Data Analysis, 142 (??):??, February 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2020:EBb

[12] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 142(??):Article 106854, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947319302099.

Sugrue:2020:IAN

[13] Darren Sugrue. Introducing article numbering to Computational Statistics and Data Analysis. Computational Statistics & Data Analysis, 142(??): Article 106857, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302129.

Wang:2020:EAA

[14] Cheng Wang and Binyan Jiang. An efficient ADMM algorithm for high dimensional precision matrix estimation via penalized quadratic loss. Computational Statistics & Data Analysis, 142(??): Article 106812, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301574.

Zhu:2020:CSO

[15] Kailun Zhu, Dorota Kurowicka, and Gabriela F. Nane. Common sampling orders of regular vines with application to model selection. Computational Statistics & Data Analysis, 142(??): Article 106811, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301562.

Zhao:2020:SPG

[16] Yaqing Zhao and Howard Bondell. Solution paths for the generalized lasso with applications to spatially varying coefficients regression. Computational Statistics & Data Analysis, 142(??): Article 106821, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301689.

Sarkar:2020:PMM

[17] Shuchismita Sarkar, Xuwen Zhu, Volodymyr Melnykov, and Salvatore Ingrassia. On parsimonious models for modeling matrix data. Computational Statistics & Data Analysis, 142(??):

Article 106822, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301690.

Sun:2020:FFQ

[18] Yifan Sun and Qihua Wang. Functionon-function quadratic regression models. Computational Statistics & Data Analysis, 142(??):Article 106814, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947319301616.

Liu:2020:GPQ

[19] Yongxin Liu, Peng Zeng, and Lu Lin. Generalized ℓ_{1}-penalized quantile regression with linear constraints. Computational Statistics & Data Analysis, 142(??):Article 106819, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947319301665.

Barthel:2020:PCV

[20] Nicole Barthel, Claudia Czado, and Yarema Okhrin. A partial correlation vine based approach for modeling and forecasting multivariate volatility time-series. Computational Statistics & Data Analysis, 142(??):Article 106810, February 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301550.

Yu:2020:VND

[21] Weichang Yu, Lamiae Azizi, and John T. Ormerod. Variational nonparametric discriminant analysis. Computational Statistics & Data Analysis, 142(??):Article 106817, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947319301641.

Jentsch:2020:TDP

[22] Carsten Jentsch, Eun Ryung Lee, and Enno Mammen. Time-dependent Poisson reduced rank models for political text data analysis. Computational Statistics & Data Analysis, 142(??): Article 106813, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301586.

Jokiel-Rokita:2020:ERC

[23] Alicja Jokiel-Rokita and Rafał Topolnicki. Estimation of the ROC curve from the Lehmann family. Computational Statistics & Data Analysis, 142(??): Article 106820, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301677.

Yan:2020:QBD

[24] Mei Yan, Efang Kong, and Yingcun Xia. Quantile based dimension reduction in censored regression. Computational Statistics & Data Analysis, 142(??): Article 106818, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301653.

Liu:2020:SMC

[25] Jin Liu, Yingying Ma, and Hansheng Wang. Semiparametric model for covariance regression analysis. Computational Statistics & Data Analysis, 142(??): Article 106815, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301628.

Ni:2020:FSU

[26] Lyu Ni, Fang Fang, and Jun Shao. Feature screening for ultrahigh dimensional categorical data with covariates missing at random. Computational Statistics & Data Analysis, 142(??): Article 106824, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301719.

Caimo:2020:MER

[27] Alberto Caimo and Isabella Gollini. A multilayer exponential random graph modelling approach for weighted networks. Computational Statistics & Data Analysis, 142(??):Article 106825, February 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301720.

Ma:2020:SMM

[28] Huijuan Ma, Wei Zhao, and Yong Zhou. Semiparametric model of mean residual life with biased sampling data. Computational Statistics & Data Analysis, 142(??):Article 106826, February 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301732.

Yu:2020:AML

[29] Ting-Hung Yu, Henghsiu Tsai, and Heiko Rachinger. Approximate maximum likelihood estimation of a threshold diffusion process. Computational Statistics & Data Analysis, 142(??): Article 106823, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301707.

Li:2020:NFS

[30] Xiaoxia Li, Niansheng Tang, Jinhan Xie, and Xiaodong Yan. A nonparametric feature screening method for ultrahigh-dimensional missing response. Computational Statistics & Data Analysis, 142(??):Article 106828, February 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301756.

Song:2020:RTD

[31] Junmo Song. Robust test for dispersion parameter change in discretely observed diffusion processes. Computational Statistics & Data Analysis, 142(??): Article 106832, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301872.

Zhao:2020:SPC

[32] Yi Zhao, Martin A. Lindquist, and Brian S. Caffo. Sparse principal component based high-dimensional mediation analysis. Computational Statistics & Data Analysis, 142(??):Article 106835, February 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301902.

Florez:2020:GRC

[33] Alvaro Jóse Flórez, Ariel Alonso Abad, Geert Molenberghs, and Wim Van Der Elst. Generating random correlation matrices with fixed values: an application to the evaluation of multivariate surrogate endpoints. Computational Statistics & Data Analysis, 142(??): Article 106834, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301896.

Bianco:2020:RWT

[34] Ana M. Bianco, Graciela Boente, and Isabel M. Rodrigues. Robust Waldtype methods for testing equality between two populations regression parameters: a comparative study under the logistic model. Computational Statistics & Data Analysis, 142(??): Article 106827, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301744.

Gaigall:2020:RWS

[35] Daniel Gaigall. Rothman–Woodroofe symmetry test statistic revisited. Computational Statistics & Data Analysis, 142(??):Article 106837, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947319301926.

Kwon:2020:UQU

[36] Yongchan Kwon, Joong-Ho Won, Beom Joon Kim, and Myunghee Cho Paik. Uncertainty quantification using Bayesian neural networks in classification: Application to biomedical image segmentation. Computational Statistics & Data Analysis, 142(??): Article 106816, February 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930163X.

Anonymous:2020:Ma

[37] Anonymous. March 2020. Computational Statistics & Data Analysis, 143 (??):??, March 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2020:EBc

[38] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 143(??):Article 106871, March 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0167947319302269.

Spezia:2020:BVS

[39] Luigi Spezia. Bayesian variable selection in non-homogeneous hidden Markov models through an evolutionary Monte Carlo method. Computational Statistics & Data Analysis, 143(??): Article 106840, March 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301951.

Duan:2020:DCS

[40] Jin-Chuan Duan, Andras Fulop, and Yu-Wei Hsieh. Data-cloning SMC^{2}: a global optimizer for maximum likelihood estimation of latent variable models. Computational Statistics & Data Analysis, 143(??):Article 106841, March 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301963.

Ma:2020:ALS

[41] Yingying Ma, Wei Lan, Fanying Zhou, and Hansheng Wang. Approximate least squares estimation for spatial autoregressive models with covariates. Computational Statistics & Data Analysis, 143(??):Article 106833, March 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301884.

Duarte-Lopez:2020:ZPS

[42] Ariel Duarte-López, Marta Pérez-Casany, and Jordi Valero. The Zipf–Poisson-stopped-sum distribution with an application for modeling the degree sequence of social networks. Computational Statistics & Data Analysis, 143(??):Article 106838, March 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301938.

Fouskakis:2020:VPE

[43] Dimitris Fouskakis, Ioannis Ntzoufras, and Konstantinos Perrakis. Variations of power-expected-posterior priors in normal regression models. Computational Statistics & Data Analysis, 143(??): Article 106836, March 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301914.

Bommert:2020:BFM

[44] Andrea Bommert, Xudong Sun, Bernd Bischl, Jörg Rahnenführer, and Michel Lang. Benchmark for filter methods for feature selection in high-dimensional classification data. Computational Statistics & Data Analysis, 143(??): Article 106839, March 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930194X.

Rodriguez:2020:BMM

[45] Carlos E. Rodríguez, Gabriel Núñez-Antonio, and Gabriel Escarela. A Bayesian mixture model for clustering circular data. Computational Statistics & Data Analysis, 143(??):Article 106842, March 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301975.

McCloud:2020:DNE

[46] Nadine McCloud and Christopher F. Parmeter. Determining the number of effective parameters in kernel density estimation. Computational Statistics & Data Analysis, 143(??):Article 106843, March 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301987.

Han:2020:BMC

[47] Ningren Han and Rajeev J. Ram. Bayesian modeling and computation for analyte quantification in complex mixtures using Raman spectroscopy. Computational Statistics & Data Analysis, 143(??):Article 106846, March 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302014.

Anonymous:2020:Aa

[48] Anonymous. April 2020. Computational Statistics & Data Analysis, 144 (??):??, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2020:EBd

[49] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 144 (??):Article 106921, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300128.

Garcia-Rodenas:2020:CGP

[50] Ricardo García-Ródenas, José Carlos García-García, Jesús López-Fidalgo, José Ángel Martín-Baos, and Weng Kee Wong. A comparison of generalpurpose optimization algorithms for finding optimal approximate experimental designs. Computational Statistics & Data Analysis, 144(??):Article 106844, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301999.

Yang:2020:SIM

[51] Jing Yang, Guoliang Tian, Fang Lu, and Xuewen Lu. Single-index modal regression via outer product gradients. Computational Statistics & Data Analysis, 144(??):Article 106867, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302221.

Taylor:2020:MUN

[52] Simon A. C. Taylor, Chris Sherlock, Gareth Ridall, and Paul Fearnhead. Motor unit number estimation via sequential Monte Carlo. Computational Statistics & Data Analysis, 144(??): Article 106845, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302002.

Posch:2020:NBA

[53] Konstantin Posch, Maximilian Arbeiter, and Juergen Pilz. A novel Bayesian approach for variable selection in linear regression models. Computational Statistics & Data Analysis, 144(??): Article 106881, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302361.

Pan:2020:ECD

[54] Yuqing Pan and Qing Mai. Efficient computation for differential network analysis with applications to quadratic discriminant analysis. Computational Statistics & Data Analysis, 144(??): Article 106884, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302397.

Wang:2020:MPG

[55] Dewei Wang, Chuan-Fa Tang, and Joshua M. Tebbs. More powerful goodness-of-fit tests for uniform stochastic ordering. Computational Statistics & Data Analysis, 144(??):Article 106898, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302531.

vonSchroeder:2020:ECJ

[56] Jonathan von Schroeder and Thorsten Dickhaus. Efficient calculation of the joint distribution of order statistics. Computational Statistics & Data Analysis, 144(??):Article 106899, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947319302543.

Shen:2020:CML

[57] Pao sheng Shen and Huichen Hsu. Conditional maximum likelihood estimation for semiparametric transformation models with doubly truncated data. Computational Statistics & Data Analysis, 144 (??):Article 106862, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302178.

Lai:2020:EQC

[58] Yuanhao Lai and Ian McLeod. Ensemble quantile classifier. Computational Statistics & Data Analysis, 144 (??):Article 106849, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930204X.

Yi:2020:RAI

[59] Fengting Yi, Niansheng Tang, and Jianguo Sun. Regression analysis of intervalcensored failure time data with timedependent covariates. Computational Statistics & Data Analysis, 144(??): Article 106848, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302038.

Arellano-Valle:2020:TPN

[60] Reinaldo B. Arellano-Valle, Adelchi Azzalini, Clécio S. Ferreira, and Karol Santoro. A two-piece normal measurement error model. Computational Statistics & Data Analysis, 144(??): Article 106863, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930218X.

Selosse:2020:MBC

[61] Margot Selosse, Julien Jacques, and Christophe Biernacki. Model-based coclustering for mixed type data. Computational Statistics & Data Analysis, 144 (??):Article 106866, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930221X.

Walder:2020:BAS

[62] Adam Walder and Ephraim M. Hanks. Bayesian analysis of spatial generalized linear mixed models with Laplace moving average random fields. Computational Statistics & Data Analysis, 144 (??):Article 106861, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302166.

Borrajo:2020:BKI

[63] M. I. Borrajo, W. González-Manteiga, and M. D. Martínez-Miranda. Bootstrapping kernel intensity estimation for inhomogeneous point processes with spatial covariates. Computational Statistics & Data Analysis, 144(??):Article 106875, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302300.

Han:2020:SEN

[64] Bo Han and Xiaoguang Wang. Semiparametric estimation for the nonmixture cure model in case-cohort and nested case-control studies. Computational Statistics & Data Analysis, 144 (??):Article 106874, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302294.

Tian:2020:BBR

[65] Yuzhu Tian and Xinyuan Song. Bayesian bridge-randomized penalized quantile regression. Computational Statistics & Data Analysis, 144(??): Article 106876, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302312.

Xue:2020:ELP

[66] Liugen Xue and Jinghua Zhang. Empirical likelihood for partially linear single-index models with missing observations. Computational Statistics & Data Analysis, 144(??):Article 106877, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302324.

Puig:2020:SGF

[67] Pedro Puig and Christian H. Weiß. Some goodness-of-fit tests for the Poisson distribution with applications in biodosimetry. Computational Statistics & Data Analysis, 144(??):Article 106878, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302336.

Paci:2020:SLC

[68] Lucia Paci and Guido Consonni. Structural learning of contemporaneous dependencies in graphical VAR models. Computational Statistics & Data Analysis, 144(??):Article 106880, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930235X.

Deresa:2020:MNR

[69] Negera Wakgari Deresa and Ingrid Van Keilegom. A multivariate normal regression model for survival data subject to different types of dependent censoring. Computational Statistics & Data Analysis, 144(??):Article 106879, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302348.

Liu:2020:SAH

[70] Mengque Liu, Qingzhao Zhang, Kuangnan Fang, and Shuangge Ma. Structured analysis of the high-dimensional FMR model. Computational Statistics & Data Analysis, 144(??):Article 106883, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302385.

Zhang:2020:CCI

[71] Likun Zhang, Enrique del Castillo, Andrew J. Berglund, Martin P. Tingley, and Nirmal Govind. Computing confidence intervals from massive data via penalized quantile smoothing splines. Computational Statistics & Data Analysis, 144(??):Article 106885, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302403.

Feng:2020:HDS

[72] Long Feng, Xiaoxu Zhang, and Binghui Liu. A high-dimensional spatial rank test for two-sample location problems. Computational Statistics & Data Analysis, 144(??):Article 106889, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302440.

Zhang:2020:SEL

[73] Ting Zhang and Lei Wang. Smoothed empirical likelihood inference and variable selection for quantile regression with nonignorable missing response. Computational Statistics & Data Analysis, 144(??):Article 106888, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302439.

Zhang:2020:BID

[74] Tingting Zhang, Yinge Sun, Huazhang Li, Guofen Yan, Seiji Tanabe, Ruizhong Miao, Yaotian Wang, Brian S. Caffo, and Mark S. Quigg. Bayesian inference of a directional brain network model for intracranial EEG data. Computational Statistics & Data Analysis, 144 (??):Article 106847, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302026.

Liu:2020:GFT

[75] Juxin Liu, Yanyuan Ma, and Jill Johnstone. A goodness-of-fit test for zeroinflated Poisson mixed effects models in tree abundance studies. Computational Statistics & Data Analysis, 144 (??):Article 106887, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302427.

Wang:2020:EAH

[76] Shuying Wang, Chunjie Wang, Peijie Wang, and Jianguo Sun. Estimation of the additive hazards model with case K interval-censored failure time data in the presence of informative censoring. Computational Statistics & Data Analysis, 144(??): Article 106891, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302464.

Im:2020:LST

[77] Jongho Im, Kosuke Morikawa, and Hyung-Tae Ha. A least squarestype density estimator using a polynomial function. Computational Statistics & Data Analysis, 144(??):Article 106882, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302373.

Chen:2020:QRB

[78] Lanjue Chen and Yong Zhou. Quantile regression in big data: a divide and conquer based strategy. Computational Statistics & Data Analysis, 144 (??):Article 106892, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302476.

Serra:2020:ELD

[79] Paulo Serra and Michel Mandjes. Estimation of local degree distributions via local weighted averaging and Monte Carlo cross-validation. Computational Statistics & Data Analysis, 144(??): Article 106886, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302415.

Chen:2020:MCF

[80] Feifei Chen, Qing Jiang, Zhenghui Feng, and Lixing Zhu. Model checks for functional linear regression models based on projected empirical processes. Computational Statistics & Data Analysis, 144 (??):Article 106897, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930252X.

Qiu:2020:GVS

[81] Debin Qiu and Jeongyoun Ahn. Grouped variable screening for ultrahigh dimensional data for linear model. Computational Statistics & Data Analysis, 144(??):Article 106894, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731930249X.

Xu:2020:BSB

[82] Ganggang Xu, Huirong Zhu, and J. Jack Lee. Borrowing strength and borrowing index for Bayesian hierarchical models. Computational Statistics & Data Analysis, 144(??):Article 106901, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302567.

Harrar:2020:NPP

[83] Solomon W. Harrar, Merga B. Feyasa, and Eshetu Wencheko. Nonparametric procedures for partially paired data in two groups. Computational Statistics & Data Analysis, 144(??): Article 106903, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302580.

Liao:2020:CMC

[84] Jun Liao and Guohua Zou. Corrected Mallows criterion for model averaging. Computational Statistics & Data Analysis, 144(??):Article 106902, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302579.

Lovato:2020:MFT

[85] Ilenia Lovato, Alessia Pini, Aymeric Stamm, and Simone Vantini. Modelfree two-sample test for networkvalued data. Computational Statistics & Data Analysis, 144(??):Article 106896, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302518.

Huskova:2020:TVS

[86] Marie Husková, Simos G. Meintanis, and Charl Pretorius. Tests for validity of the semiparametric heteroskedastic transformation model. Computational Statistics & Data Analysis, 144 (??):Article 106895, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302506.

Santitissadeekorn:2020:AFC

[87] Naratip Santitissadeekorn, David J. B. Lloyd, Martin B. Short, and Sylvain Delahaies. Approximate filtering of conditional intensity process for Poisson count data: Application to urban crime. Computational Statistics & Data Analysis, 144(??):Article 106850, April 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302051.

You:2020:IER

[88] Honglong You, Junyi Guo, and Jiancheng Jiang. Interval estimation of the ruin probability in the classical compound Poisson risk model. Computational Statistics & Data Analysis, 144 (??):Article 106890, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302452.

Carapia:2020:BCL

[89] Gustavo Quintana Carapia, Ivan Markovsky, Rik Pintelon, Péter Zoltán Csurcsia, and Dieter Verbeke. Bias and covariance of the least squares estimate in a structured errors-in-variables problem. Computational Statistics & Data Analysis, 144(??): Article 106893, April 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302488.

Anonymous:2020:Mb

[90] Anonymous. May 2020. Computational Statistics & Data Analysis, 145 (??):??, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2020:EBe

[91] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 145 (??):Article 106935, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320300268.

Fabrizi:2020:RBS

[92] Enrico Fabrizi, Nicola Salvati, and Carlo Trivisano. Robust Bayesian small area estimation based on quantile regression. Computational Statistics & Data Analysis, 145(??):Article 106900, May 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302555.

Kloodt:2020:STS

[93] Nick Kloodt and Natalie Neumeyer. Specification tests in semiparametric transformation models — a multiplier bootstrap approach. Computational Statistics & Data Analysis, 145(??): Article 106908, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302634.

Nakamura:2020:ABR

[94] Kento Nakamura, Keisuke Yano, and Fumiyasu Komaki. Adjacency-based regularization for partially ranked data with non-ignorable missing. Computational Statistics & Data Analysis, 145 (??):Article 106905, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302609.

Li:2020:ALM

[95] Jingjing Li, Zimu Chen, Zhanfeng Wang, and Yuan chin Ivan Chang. Active learning in multiple-class classification problems via individualized binary models. Computational Statistics & Data Analysis, 145(??):Article 106911, May 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320300025.

Sugasawa:2020:EIA

[96] Shonosuke Sugasawa, Genya Kobayashi, and Yuki Kawakubo. Estimation and inference for area-wise spatial income distributions from grouped data. Computational Statistics & Data Analysis, 145 (??):Article 106904, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302592.

Han:2020:SBE

[97] Kyunghee Han, Young K. Lee, and Byeong U. Park. Smooth backfitting for errors-in-variables varying coefficient regression models. Computational Statistics & Data Analysis, 145 (??):Article 106909, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302646.

Oguz-Alper:2020:MMD

[98] Melike Oğuz-Alper and Yves G. Berger. Modelling multilevel data under complex sampling designs: an empirical likelihood approach. Computational Statistics & Data Analysis, 145(??): Article 106906, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319302610.

Gehman:2020:OSA

[99] Andrew Gehman and William W. S. Wei. Optimal spatial aggregation of space-time models and applications. Computational Statistics & Data Analysis, 145(??):Article 106913, May 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300049.

Zhang:2020:NAV

[100] Xiaoke Zhang, Qixian Zhong, and Jane-Ling Wang. A new approach to varyingcoefficient additive models with longitudinal covariates. Computational Statistics & Data Analysis, 145(??): Article 106912, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300037.

Jiang:2020:LRM

[101] Wei Jiang, Julie Josse, Marc Lavielle, and TraumaBase Group. Logistic regression with missing covariates — parameter estimation, model selection and prediction within a jointmodeling framework. Computational Statistics & Data Analysis, 145(??): Article 106907, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319302622.

Chu:2020:BRP

[102] Jeffrey Chu, Yuanyuan Zhang, Stephen Chan, and Saralees Nadarajah. Bias reduction in the population size estimation of large data sets. Computational Statistics & Data Analysis, 145 (??):Article 106914, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300050.

Lin:2020:WQR

[103] Fangzheng Lin, Yanlin Tang, and Zhongyi Zhu. Weighted quantile regression in varying-coefficient model with longitudinal data. Computational Statistics & Data Analysis, 145(??): Article 106915, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300062.

Bedoui:2020:BEL

[104] Adel Bedoui and Nicole A. Lazar. Bayesian empirical likelihood for ridge and lasso regressions. Computational Statistics & Data Analysis, 145(??): Article 106917, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300086.

Yoder:2020:VNC

[105] Jordan Yoder, Li Chen, Henry Pao, Eric Bridgeford, Keith Levin, Donniell E. Fishkind, Carey Priebe, and Vince Lyzinski. Vertex nomination: the canonical sampling and the extended spectral nomination schemes. Computational Statistics & Data Analysis, 145 (??):Article 106916, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300074.

Jhwueng:2020:MRA

[106] Dwueng-Chwuan Jhwueng. Modeling rate of adaptive trait evolution using Cox–Ingersoll–Ross process: an approximate Bayesian computation approach. Computational Statistics & Data Analysis, 145(??): Article 106924, May 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300153.

Anonymous:2020:Ja

[107] Anonymous. June 2020. Computational Statistics & Data Analysis, 146 (??):??, June 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2020:EBf

[108] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 146 (??):Article 106948, June 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300396.

Zhao:2020:NET

[109] Jun Zhao, Hea-Jung Kim, and Hyoung-Moon Kim. New EM-type algorithms for the Heckman selection model. Computational Statistics & Data Analysis, 146 (??):Article 106930, June 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300219.

Rha:2020:DOS

[110] Hyungmin Rha, Ming-Hung Kao, and Rong Pan. Design optimal sampling plans for functional regression models. Computational Statistics & Data Analysis, 146(??):Article 106925, June 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300165.

Zhu:2020:DDT

[111] Xuehu Zhu, Xu Guo, Tao Wang, and Lixing Zhu. Dimensionality determination: a thresholding double ridge ratio approach. Computational Statistics & Data Analysis, 146(??): Article 106910, June 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300013.

Kirsner:2020:MSS

[112] Daniel Kirsner and Bruno Sansó. Multiscale shotgun stochastic search for large spatial datasets. Computational Statistics & Data Analysis, 146(??): Article 106931, June 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300220.

Eberl:2020:ADP

[113] Andreas Eberl and Bernhard Klar. Asymptotic distributions and performance of empirical skewness measures. Computational Statistics & Data Analysis, 146(??):Article 106939, June 2020. CODEN CSDADW.

ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S016794732030030X.

Anonymous:2020:Jb

[114] Anonymous. July 2020. Computational Statistics & Data Analysis, 147 (??):??, July 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2020:EBg

[115] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 147 (??):Article 106967, July 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732030058X.

Filzmoser:2020:CRR

[116] P. Filzmoser, S. Höppner, I. Ortner, S. Serneels, and T. Verdonck. Cellwise robust M regression. Computational Statistics & Data Analysis, 147 (??):Article 106944, July 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300359.

Liu:2020:SVA

[117] Wenchen Liu, Yincai Tang, and Xianyi Wu. Separating variables to accelerate non-convex regularized optimization. Computational Statistics & Data Analysis, 147(??):Article 106943, July 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300347.

Qin:2020:GMQ

[118] Shanshan Qin and Yuehua Wu. General matching quantiles M-estimation. Computational Statistics & Data Analysis, 147(??):Article 106941, July 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0167947320300323.

DeBlasi:2020:IID

[119] Pierpaolo De Blasi, Asael Fabian Martínez, Ramsés H. Mena, and Igor Prünster. On the inferential implications of decreasing weight structures in mixture models. Computational Statistics & Data Analysis, 147(??): Article 106940, July 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320300311.

Lu:2020:JMF

[120] Shuiyun Lu, Xiaolin Chen, Sheng Xu, and Chunling Liu. Joint model-free feature screening for ultra-high dimensional semi-competing risks data. Computational Statistics & Data Analysis, 147 (??):Article 106942, July 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320300335.

Fordellone:2020:FGS

[121] Mario Fordellone and Maurizio Vichi. Finding groups in structural equation modeling through the partial least squares algorithm. Computational Statistics & Data Analysis, 147(??): Article 106957, July 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300487.

Anonymous:2020:Ab

[122] Anonymous. August 2020. Computational Statistics & Data Analysis, 148 (??):??, August 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2020:EBh

[123] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 148(??):Article 106982, August 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947320300736.

Monterrubio-Gomez:2020:PIS

[124] Karla Monterrubio-Gómez, Lassi Roininen, Sara Wade, Theodoros Damoulas, and Mark Girolami. Posterior inference for sparse hierarchical nonstationary models. Computational Statistics & Data Analysis, 148(??): Article 106954, August 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300451.

Jeon:2020:PPA

[125] Jong-June Jeon, Yongdai Kim, Sungho Won, and Hosik Choi. Primal path algorithm for compositional data analysis. Computational Statistics & Data Analysis, 148(??):Article 106958, August 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300499.

Gao:2020:AIC

[126] Zhikun Gao, Yanlin Tang, Huixia Judy Wang, Guangying K. Wu, and Jeff Lin. Automatic identification of curve shapes with applications to ultrasonic vocalization. Computational Statistics & Data Analysis, 148(??):Article 106956, August 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300475.

Najarzadeh:2020:STZ

[127] Dariush Najarzadeh. A simple test for zero multiple correlation coefficient in high-dimensional normal data using random projection. Computational Statistics & Data Analysis, 148(??): Article 106955, August 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300463.

Yang:2020:EFM

[128] Yu-Chen Yang, Tsung-I Lin, Luis M. Castro, and Wan-Lun Wang. Extending finite mixtures of t linear mixed-effects models with concomitant covariates. Computational Statistics & Data Analysis, 148(??):Article 106961, August 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300529.

Anonymous:2020:S

[129] Anonymous. September 2020. Computational Statistics & Data Analysis, 149 (??):??, September 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2020:EBi

[130] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 149(??):Article 107000, September 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947320300918.

Lee:2020:OUM

[131] JooChul Lee, HaiYing Wang, and Elizabeth D. Schifano. Online updating method to correct for measurement error in big data streams. Computational Statistics & Data Analysis, 149(??):Article 106976, September 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947320300670.

Hui:2020:LLI

[132] Francis K. C. Hui, Samuel Müller, and A. H. Welsh. The LASSO on latent indices for regression modeling with ordinal categorical predictors. Computational Statistics & Data Analysis, 149(??):Article 106951, September 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300426.

Rivest:2020:STD

[133] Louis-Paul Rivest and Sergio Ewane Ebouele. Sampling a two dimensional matrix. Computational Statistics & Data Analysis, 149(??):Article 106971, September 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320300621.

Ma:2020:BNT

[134] Zichen Ma and Timothy E. Hanson. Bayesian nonparametric test for independence between random vectors. Computational Statistics & Data Analysis, 149(??):Article 106959, September 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320300505.

Dai:2020:FOD

[135] Wenlin Dai, Tomáš Mrkvička, Ying Sun, and Marc G. Genton. Functional outlier detection and taxonomy by sequential transformations. Computational Statistics & Data Analysis, 149(??):Article 106960, September 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320300517.

Hofmeyr:2020:DFM

[136] David P. Hofmeyr. Degrees of freedom and model selection for k-means clustering. Computational Statistics & Data Analysis, 149(??):Article 106974, September 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300657.

Zhang:2020:FSU

[137] Jing Zhang, Qihua Wang, and Jian Kang. Feature screening under missing indicator imputation with non-ignorable missing response. Computational Statistics & Data Analysis, 149(??):Article 106975, September 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300669.

Chen:2020:RDF

[138] Yaqing Chen, Matthew Dawson, and Hans-Georg Müller. Rank dynamics for functional data. Computational Statistics & Data Analysis, 149(??):Article 106963, September 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300542.

Anonymous:2020:O

[139] Anonymous. October 2020. Computational Statistics & Data Analysis, 150 (??):??, October 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2020:EBj

[140] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 150(??):Article 107023, October 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947320301146.

Vo-Thanh:2020:AQS

[141] Nha Vo-Thanh and Hans-Peter Piepho. Augmented quasi-sudoku designs in field trials. Computational Statistics & Data Analysis, 150(??):Article 106988, October 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320300797.

Jaeger:2020:SSE

[142] Adam Jaeger and Nicole A. Lazar. Split sample empirical likelihood. Computational Statistics & Data Analysis, 150(??):Article 106994, October 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947320300852.

Zhao:2020:CNC

[143] Shi Zhao, Giorgos Bakoyannis, Spencer Lourens, and Wanzhu Tu. Comparison of nonlinear curves and surfaces. Computational Statistics & Data Analysis, 150(??):Article 106987, October 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320300785.

Cheng:2020:TPT

[144] Guanghui Cheng, Baisen Liu, Guoliang Tian, and Shurong Zheng. Testing proportionality of two high-dimensional covariance matrices. Computational Statistics & Data Analysis, 150(??): Article 106962, October 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300530.

Emura:2020:EMW

[145] Takeshi Emura and Jiun-Huang Hsu.

Estimation of the Mann–Whitney effect in the two-sample problem under dependent censoring. Computational Statistics & Data Analysis, 150(??): Article 106990, October 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300815.

Lo:2020:NCD

[146] Simon M. S. Lo, Enno Mammen, and Ralf A. Wilke. A nested copula duration model for competing risks with multiple spells. Computational Statistics & Data Analysis, 150(??): Article 106986, October 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300773.

Li:2020:GCC

[147] Gen Li. Generalized co-clustering analysis via regularized alternating least squares. Computational Statistics & Data Analysis, 150(??):Article 106989, October 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300803.

Lee:2020:BWB

[148] Taewook Lee and Changryong Baek.

Block wild bootstrap-based CUSUM tests robust to high persistence and misspecification. Computational Statistics & Data Analysis, 150(??):Article 106996, October 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300876.

Xie:2020:GKB

[149] Chuanlong Xie and Lixing Zhu. Generalized kernel-based inverse regression methods for sufficient dimension reduction. Computational Statistics & Data Analysis, 150(??):Article 106995, October 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300864.

Cui:2020:PFL

[150] Xia Cui, Hongmei Lin, and Heng Lian. Partially functional linear regression in reproducing kernel Hilbert spaces. Computational Statistics & Data Analysis, 150(??):Article 106978, October 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300694.

Chung:2020:RBQ

[151] Ray S. W. Chung, Mike K. P. So, Amanda M. Y. Chu, and Thomas W. C. Chan. Regularization of Bayesian quasilikelihoods constructed from complex estimating functions. Computational Statistics & Data Analysis, 150(??): Article 106977, October 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300682.

Park:2020:SMT

[152] Jun Young Park, Joerg Polzehl, Snigdhansu Chatterjee, André Brechmann, and Mark Fiecas. Semiparametric modeling of time-varying activation and connectivity in task-based fMRI data. Computational Statistics & Data Analysis, 150(??):Article 107006, October 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300979.

Yang:2020:FLR

[153] Seong J. Yang, Hyejin Shin, Sang Han Lee, Seokho Lee, and Alzheimerś Disease Neuroimaging Initiative. Functional linear regression model with randomly censored data: Predicting conversion time to Alzheimer ’s disease. Computational Statistics & Data Analysis, 150(??):Article 107009, October 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301006.

Albano:2020:ITN

[154] G. Albano and V. Giorno. Inferring time non-homogeneous Ornstein Uhlenbeck type stochastic process. Computational Statistics & Data Analysis, 150(??): Article 107008, October 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300992.

Balabdaoui:2020:CMD

[155] Fadoua Balabdaoui and Yulia Kulagina. Completely monotone distributions: Mixing, approximation and estimation of number of species. Computational Statistics & Data Analysis, 150(??): Article 107014, October 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301055.

Bhattacharya:2020:SBE

[156] Rabi Bhattacharya and Rachel Oliver. Superiority of Bayes estimators over the MLE in high dimensional multinomial models and its implication for nonparametric Bayes theory. Computational Statistics & Data Analysis, 150(??): Article 107011, October 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732030102X.

Anonymous:2020:N

[157] Anonymous. November 2020. Computational Statistics & Data Analysis, 151 (??):??, November 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2020:EBk

[158] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 151(??):Article 107035, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947320301262.

Nguyen:2020:VIH

[159] Hoang Nguyen, M. Concepción Ausín, and Pedro Galeano. Variational inference for high dimensional structured factor copulas. Computational Statistics & Data Analysis, 151(??):Article 107012, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301031.

Philipson:2020:FMC

[160] Pete Philipson, Graeme L. Hickey, Michael J. Crowther, and Ruwanthi Kolamunnage-Dona. Faster Monte Carlo estimation of joint models for time-to-event and multivariate longitudinal data. Computational Statistics & Data Analysis, 151(??):Article 107010, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301018.

Edwards:2020:MPS

[161] Matthew Edwards, Stefano Castruccio, and Dorit Hammerling. Marginally parameterized spatio-temporal models and stepwise maximum likelihood estimation. Computational Statistics & Data Analysis, 151(??):Article 107018, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301092.

He:2020:HDT

[162] Yong He, Mingjuan Zhang, Xinsheng Zhang, and Wang Zhou. Highdimensional two-sample mean vectors test and support recovery with factor adjustment. Computational Statistics & Data Analysis, 151(??):Article 107004, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300955.

Lin:2020:HTP

[163] Liang-Ching Lin, Ray-Bing Chen, Mong-Na Lo Huang, and Meihui Guo. Huber-type principal expectile component analysis. Computational Statistics & Data Analysis, 151(??):Article 106992, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300839.

Ivanovic:2020:CST

[164] Blagoje Ivanović, Bojana Milošević, and Marko Obradović. Comparison of symmetry tests against some skewsymmetric alternatives in i.i.d. and noni.i.d. setting. Computational Statistics & Data Analysis, 151(??):Article 106991, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300827.

Hao:2020:CRM

[165] Meiling Hao, Xingqiu Zhao, and Wei Xu. Competing risk modeling and testing for X-chromosome genetic association. Computational Statistics & Data Analysis, 151(??):Article 107007, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300980.

Wang:2020:JAS

[166] Xiaoqing Wang, Xiangnan Feng, and Xinyuan Song. Joint analysis of semicontinuous data with latent variables. Computational Statistics & Data Analysis, 151(??):Article 107005, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300967.

Han:2020:ASE

[167] Zhong-Cheng Han, Jin-Guan Lin, and Yan-Yong Zhao. Adaptive semiparametric estimation for single index models with jumps. Computational Statistics & Data Analysis, 151(??):Article 107013, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301043.

Zhang:2020:RBA

[168] Haixiang Zhang, Jian Huang, and Liuquan Sun. A rank-based approach to estimating monotone individualized two treatment regimes. Computational Statistics & Data Analysis, 151(??):Article 107015, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947320301067.

Ahn:2020:RMU

[169] Kyungmin Ahn, J. Derek Tucker, Wei Wu, and Anuj Srivastava. Regression models using shapes of functions as predictors. Computational Statistics & Data Analysis, 151(??):Article 107017, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301080.

Zhang:2020:NBI

[170] Shibin Zhang. Nonparametric Bayesian inference for the spectral density based on irregularly spaced data. Computational Statistics & Data Analysis, 151(??):Article 107019, November 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947320301109.

Anonymous:2020:D

[171] Anonymous. December 2020. Computational Statistics & Data Analysis, 152 (??):??, December 2020. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2020:EB

[172] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 152(??):Article 107076, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947320301675.

Liu:2020:DTL

[173] Yehong Liu and Guosheng Yin. The Delaunay triangulation learner and its ensembles. Computational Statistics & Data Analysis, 152(??):Article 107030, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301213.

Babkin:2020:LSE

[174] Sergii Babkin, Jonathan R. Stewart, Xiaochen Long, and Michael Schweinberger. Large-scale estimation of random graph models with local dependence. Computational Statistics & Data Analysis, 152(??):Article 107029, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301201.

Zhao:2020:JEL

[175] Yichuan Zhao, Yueju Su, and Hanfang Yang. Jackknife empirical likelihood inference for the Pietra ratio. Computational Statistics & Data Analysis, 152(??):Article 107049, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947320301407.

Nie:2020:SFP

[176] Yunlong Nie and Jiguo Cao. Sparse functional principal component analysis in a new regression framework. Computational Statistics & Data Analysis, 152(??):Article 107016, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947320301079.

Le:2020:ALD

[177] Khuyen T. Le, Caroline Chaux, Frédéric J. P. Richard, and Eric Guedj. An adapted linear discriminant analysis with variable selection for the classification in highdimension, and an application to medical data. Computational Statistics & Data Analysis, 152(??):Article 107031, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301225.

Lv:2020:RKH

[178] Shaogao Lv, Zengyan Fan, Heng Lian, Taiji Suzuki, and Kenji Fukumizu. A reproducing kernel Hilbert space approach to high dimensional partially varying coefficient model. Computational Statistics & Data Analysis, 152(??):Article 107039, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301304.

Dong:2020:MFV

[179] Yuexiao Dong, Zhou Yu, and Liping Zhu. Model-free variable selection for conditional mean in regression. Computational Statistics & Data Analysis, 152(??):Article 107042, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732030133X.

Baak:2020:NCC

[180] M. Baak, R. Koopman, H. Snoek, and S. Klous. A new correlation coefficient between categorical, ordinal and interval variables with Pearson characteristics. Computational Statistics & Data Analysis, 152(??):Article 107043, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301341.

Redivo:2020:BCS

[181] Edoardo Redivo, Hien D. Nguyen, and Mayetri Gupta. Bayesian clustering of skewed and multimodal data using geometric skewed normal distributions. Computational Statistics & Data Analysis, 152(??):Article 107040, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301316.

Fang:2020:MAA

[182] Fang Fang and Zhou Yu. Model averaging assisted sufficient dimension reduction. Computational Statistics & Data Analysis, 152(??):Article 106993, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320300840.

Boente:2020:RES

[183] Graciela Boente, Matías Salibian-Barrera, and Pablo Vena. Robust estimation for semi-functional linear regression models. Computational Statistics & Data Analysis, 152(??):Article 107041, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301328.

Tonellato:2020:BNC

[184] Stefano F. Tonellato. Bayesian nonparametric clustering as a community detection problem. Computational Statistics & Data Analysis, 152(??):Article 107044, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301353.

Li:2020:SRN

[185] Xin Li, Dongya Wu, Chong Li, Jinhua Wang, and Jen-Chih Yao. Sparse recovery via nonconvex regularized M-estimators over ℓ_{q}-balls. Computational Statistics & Data Analysis, 152(??):Article 107047, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947320301389.

Liu:2020:EEM

[186] Bowen Liu and Sujit K. Ghosh. On empirical estimation of mode based on weakly dependent samples. Computational Statistics & Data Analysis, 152(??):Article 107046, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947320301377.

Tomarchio:2020:TNM

[187] Salvatore D. Tomarchio, Antonio Punzo, and Luca Bagnato. Two new matrix-variate distributions with application in model-based clustering. Computational Statistics & Data Analysis, 152(??):Article 107050, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947320301419.

Lee:2020:VCM

[188] Jihui Lee, Gen Li, and James D. Wilson. Varying-coefficient models for dynamic networks. Computational Statistics & Data Analysis, 152(??):Article 107052, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301432.

Ludkin:2020:IGS

[189] Matthew Ludkin. Inference for a generalised stochastic block model with unknown number of blocks and non-conjugate edge models. Computational Statistics & Data Analysis, 152(??):Article 107051, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947320301420.

Feng:2020:MDE

[190] Sanying Feng, Wenqi He, and Feng Li. Model detection and estimation for varying coefficient panel data models with fixed effects. Computational Statistics & Data Analysis, 152(??):Article 107054, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301456.

Chee:2020:SEL

[191] Chew-Seng Chee and Byungtae Seo. Semiparametric estimation for linear regression with symmetric errors. Computational Statistics & Data Analysis, 152(??):Article 107053, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947320301444.

Stoner:2020:AHM

[192] Oliver Stoner and Theo Economou. An advanced hidden Markov model for hourly rainfall time series. Computational Statistics & Data Analysis, 152(??):Article 107045, December 2020. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947320301365.

Anonymous:2021:Ja

[193] Anonymous. January 2021. Computational Statistics & Data Analysis, 153 (??):??, January 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2021:EBa

[194] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 153(??):Article 107101, January 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947320301924.

Bak:2021:PLD

[195] Kwan-Young Bak, Jae-Hwan Jhong, JungJun Lee, Jae-Kyung Shin, and Ja-Yong Koo. Penalized logspline density estimation using total variation penalty. Computational Statistics & Data Analysis, 153(??):Article 107060, January 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301511.

Gromping:2021:ABR

[196] Ulrike Grömping. An algorithm for blocking regular fractional factorial 2-level designs with clear two-factor interactions. Computational Statistics & Data Analysis, 153(??):Article 107059, January 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732030150X.

Zilber:2021:VLA

[197] Daniel Zilber and Matthias Katzfuss.

Vecchia–Laplace approximations of generalized Gaussian processes for big non-Gaussian spatial data. Computational Statistics & Data Analysis, 153(??): Article 107081, January 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301729.

Zhang:2021:DSS

[198] Haixiang Zhang and HaiYing Wang.

Distributed subdata selection for big data via sampling-based approach. Computational Statistics & Data Analysis, 153(??):Article 107072, January 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301638.

Zeng:2021:HSS

[199] Yaohui Zeng, Tianbao Yang, and Patrick Breheny. Hybrid safe–strong rules for efficient optimization in lassotype problems. Computational Statistics & Data Analysis, 153(??):Article 107063, January 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301547.

Ito:2021:ICR

[200] Tsubasa Ito and Shonosuke Sugasawa. Improved confidence regions in meta-analysis of diagnostic test accuracy. Computational Statistics & Data Analysis, 153(??):Article 107068, January 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301596.

Bagkavos:2021:FDL

[201] Dimitrios Bagkavos and Dimitrios Ioannides. Fixed design local polynomial smoothing and bandwidth selection for right censored data. Computational Statistics & Data Analysis, 153(??): Article 107064, January 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301559.

Liu:2021:WRE

[202] Tianqing Liu, Xiaohui Yuan, and Jianguo Sun. Weighted rank estimation for nonparametric transformation models with nonignorable missing data. Computational Statistics & Data Analysis, 153(??):Article 107061, January 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301523.

Junker:2021:ESI

[203] Robert R. Junker, Florian Griessenberger, and Wolfgang Trutschnig. Estimating scale-invariant directed dependence of bivariate distributions. Computational Statistics & Data Analysis, 153(??):Article 107058, January 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947320301493.

Machado:2021:PML

[204] Robson J. M. Machado, Ardo van den Hout, and Giampiero Marra. Penalised maximum likelihood estimation in multi-state models for intervalcensored data. Computational Statistics & Data Analysis, 153(??):Article 107057, January 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301481.

Baek:2021:TST

[205] Changryong Baek, Katheleen M. Gates, Benjamin Leinwand, and Vladas Pipiras. Two sample tests for highdimensional autocovariances. Computational Statistics & Data Analysis, 153(??):Article 107067, January 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947320301584.

Ju:2021:RBR

[206] Xiaomeng Ju and Matías Salibián-Barrera. Robust boosting for regression problems. Computational Statistics & Data Analysis, 153(??):Article 107065, January 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301560.

Fan:2021:CSM

[207] Xinyan Fan, Qingzhao Zhang, Shuangge Ma, and Kuangnan Fang. Conditional score matching for high-dimensional partial graphical models. Computational Statistics & Data Analysis, 153(??): Article 107066, January 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301572.

Chau:2021:ANP

[208] Thi Tuyet Trang Chau, Pierre Ailliot, and Valérie Monbet. An algorithm for non-parametric estimation in state–space models. Computational Statistics & Data Analysis, 153(??): Article 107062, January 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301535.

Feng:2021:RBT

[209] Long Feng, Ping Zhao, Yanling Ding, and Binghui Liu. Rank-based tests of cross-sectional dependence in panel data models. Computational Statistics & Data Analysis, 153(??):Article 107070, January 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301614.

Hebert:2021:ADP

[210] Florian Hébert, David Causeur, and Mathieu Emily. An adaptive decorrelation procedure for signal detection. Computational Statistics & Data Analysis, 153(??):Article 107082, January 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301730.

Ke:2021:EME

[211] Rui Ke, Wanbo Lu, and Jing Jia. Evaluating multiplicative error models: a residual-based approach. Computational Statistics & Data Analysis, 153(??): Article 107086, January 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301778.

Anonymous:2021:F

[212] Anonymous. February 2021. Computational Statistics & Data Analysis, 154 (??):??, February 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2021:EBb

[213] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 154(??):Article 107119, February 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947320302103.

Gressani:2021:LAF

[214] Oswaldo Gressani and Philippe Lambert. Laplace approximations for fast Bayesian inference in generalized additive models based on Psplines. Computational Statistics & Data Analysis, 154(??):Article 107088, February 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301791.

Watanabe:2021:GFT

[215] Chihiro Watanabe and Taiji Suzuki. Goodness-of-fit test for latent block models. Computational Statistics & Data Analysis, 154(??):Article 107090, February 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732030181X.

Chen:2021:SPE

[216] Tianbo Chen, Ying Sun, and Ta-Hsin Li. A semi-parametric estimation method for the quantile spectrum with an application to earthquake classification using convolutional neural network. Computational Statistics & Data Analysis, 154(??):Article 107069, February 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301602.

Fiebig:2021:DDC

[217] Ewelina Marta Fiebig. On datadriven choice of λ in nonparametric Gaussian regression via Propagation– Separation approach. Computational Statistics & Data Analysis, 154(??): Article 107087, February 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732030178X.

Chaoubi:2021:HCA

[218] Ihsan Chaoubi, Hélène Cossette, Etienne Marceau, and Christian Y. Robert. Hierarchical copulas with Archimedean blocks and asymmetric between-block pairs. Computational Statistics & Data Analysis, 154(??): Article 107071, February 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301626.

Park:2021:LHD

[219] Gunwoong Park and Yesool Kim. Learning high-dimensional Gaussian linear structural equation models with heterogeneous error variances. Computational Statistics & Data Analysis, 154(??): Article 107084, February 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301754.

vanGeloven:2021:EID

[220] N. van Geloven, Y. He, A. H. Zwinderman, and H. Putter. Estimation of incident dynamic AUC in practice. Computational Statistics & Data Analysis, 154(??):Article 107095, February 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301869.

Wang:2021:GRP

[221] Zhanfeng Wang, Maengseok Noh, Youngjo Lee, and Jian Qing Shi. A general robust t-process regression model. Computational Statistics & Data Analysis, 154(??):Article 107093, February 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301845.

Davis:2021:PNS

[222] Casey B. Davis, Christopher M. Hans, and Thomas J. Santner. Prediction of non-stationary response functions using a Bayesian composite Gaussian process. Computational Statistics & Data Analysis, 154(??):Article 107083, February 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301742.

Burgard:2021:ERT

[223] Jan Pablo Burgard, Joscha Krause, and Simon Schmaus. Estimation of regional transition probabilities for spatial dynamic microsimulations from survey data lacking in regional detail. Computational Statistics & Data Analysis, 154(??):Article 107048, February 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301390.

Anonymous:2021:Ma

[224] Anonymous. March 2021. Computational Statistics & Data Analysis, 155 (??):??, March 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2021:EBc

[225] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 155(??):Article 107135, March 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0167947320302267.

Wu:2021:MAD

[226] Runxiong Wu and Xin Chen. MM algorithms for distance covariance based sufficient dimension reduction and sufficient variable selection. Computational Statistics & Data Analysis, 155(??): Article 107089, March 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301808.

Gerber:2021:PCV

[227] Florian Gerber and Douglas W. Nychka. Parallel cross-validation: a scalable fitting method for Gaussian process models. Computational Statistics & Data Analysis, 155(??):Article 107113, March 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302048.

Zhou:2021:LPN

[228] Lin Zhou and Yayong Tang. Linearly preconditioned nonlinear conjugate gradient acceleration of the PX-EM algorithm. Computational Statistics & Data Analysis, 155(??):Article 107056, March 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732030147X.

Song:2021:RVS

[229] Yunquan Song, Xijun Liang, Yanji Zhu, and Lu Lin. Robust variable selection with exponential squared loss for the spatial autoregressive model. Computational Statistics & Data Analysis, 155(??):Article 107094, March 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301857.

Hees:2021:SII

[230] Katharina Hees, Smarak Nayak, and Peter Straka. Statistical inference for inter-arrival times of extreme events in bursty time series. Computational Statistics & Data Analysis, 155(??): Article 107096, March 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301870.

Huang:2021:JGE

[231] Youjun Huang and Jianxin Pan. Joint generalized estimating equations for longitudinal binary data. Computational Statistics & Data Analysis, 155(??): Article 107110, March 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302012.

Gorecki:2021:OPT

[232] Jan Górecki, Marius Hofert, and Ostap Okhrin. Outer power transformations of hierarchical Archimedean copulas: Construction, sampling and estimation. Computational Statistics & Data Analysis, 155(??):Article 107109, March 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302000.

Mestre:2021:FTS

[233] Guillermo Mestre, José Portela, Gregory Rice, Antonio Muñoz San Roque, and Estrella Alonso. Functional time series model identification and diagnosis by means of auto- and partial autocorrelation analysis. Computational Statistics & Data Analysis, 155(??): Article 107108, March 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301997.

Zhu:2021:SVB

[234] Kailun Zhu, Dorota Kurowicka, and Gabriela F. Nane. Simplified R-vine based forward regression. Computational Statistics & Data Analysis, 155 (??):Article 107091, March 2021. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S0167947320301821.

Wang:2021:BMR

[235] Qihua Wang, Miaomiao Su, and Ruoyu Wang. A beyond multiple robust approach for missing response problem. Computational Statistics & Data Analysis, 155(??):Article 107111, March 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320302024.

Desgagne:2021:ERE

[236] Alain Desgagné. Efficient and robust estimation of regression and scale parameters, with outlier detection. Computational Statistics & Data Analysis, 155(??):Article 107114, March 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732030205X.

Ciuperca:2021:VSH

[237] Gabriela Ciuperca. Variable selection in high-dimensional linear model with possibly asymmetric errors. Computational Statistics & Data Analysis, 155(??): Article 107112, March 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302036.

Gu:2021:SSC

[238] Lijie Gu, Suojin Wang, and Lijian Yang. Smooth simultaneous confidence band for the error distribution function in nonparametric regression. Computational Statistics & Data Analysis, 155(??):Article 107106, March 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301973.

Barrientos:2021:BGF

[239] Andrés F. Barrientos and Antonio Canale. A Bayesian goodness-of-fit test for regression. Computational Statistics & Data Analysis, 155(??): Article 107104, March 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732030195X.

Marra:2021:LBS

[240] Giampiero Marra, Alessio Farcomeni, and Rosalba Radice. Link-based survival additive models under mixed censoring to assess risks of hospitalacquired infections. Computational Statistics & Data Analysis, 155(??): Article 107092, March 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320301833.

Pun:2021:SCD

[241] Chi Seng Pun and Matthew Zakharia Hadimaja. A self-calibrated direct approach to precision matrix estimation and linear discriminant analysis in high dimensions. Computational Statistics & Data Analysis, 155(??): Article 107105, March 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301961.

Anonymous:2021:Aa

[242] Anonymous. April 2021. Computational Statistics & Data Analysis, 156 (??):??, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2021:EBd

[243] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 156 (??):Article 107172, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000062.

Yang:2021:ABC

[244] Yi Yang, Yuxuan Guo, and Xiangyu Chang. Angle-based cost-sensitive multicategory classification. Computational Statistics & Data Analysis, 156(??): Article 107107, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301985.

Lee:2021:DRN

[245] Hangsuck Lee, Hongjun Ha, and Taewon Lee. Decrement rates and a numerical method under competing risks. Computational Statistics & Data Analysis, 156 (??):Article 107125, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302164.

Goto:2021:CTV

[246] Satoshi Goto, Mariko Takagishi, and Hiroshi Yadohisa. Clustering for timevarying relational count data. Computational Statistics & Data Analysis, 156 (??):Article 107123, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302140.

Xu:2021:VSG

[247] Yang Xu, Shishun Zhao, Tao Hu, and Jianguo Sun. Variable selection for generalized odds rate mixture cure models with interval-censored failure time data. Computational Statistics & Data Analysis, 156(??): Article 107115, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302061.

Carey:2021:FSP

[248] M. Carey and J. O. Ramsay. Fast stable parameter estimation for linear dynamical systems. Computational Statistics & Data Analysis, 156(??): Article 107124, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302152.

Liu:2021:DEN

[249] Yang Liu and David Ruppert. Density estimation on a network. Computational Statistics & Data Analysis, 156 (??):Article 107128, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732030219X.

Wang:2021:NEB

[250] Yihe Wang and Sihai Dave Zhao. A nonparametric empirical Bayes approach to large-scale multivariate regression. Computational Statistics & Data Analysis, 156(??):Article 107130, April 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320302218.

Eckardt:2021:GMP

[251] Matthias Eckardt, Jonatan A. González, and Jorge Mateu. Graphical modelling and partial characteristics for multitype and multivariate-marked spatiotemporal point processes. Computational Statistics & Data Analysis, 156 (??):Article 107139, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320302309.

Li:2021:DRB

[252] Junlan Li and Tao Wang. Dimension reduction in binary response regression: a joint modeling approach. Computational Statistics & Data Analysis, 156 (??):Article 107131, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732030222X.

Zhang:2021:IGP

[253] Hong-Fan Zhang. Iterative GMM for partially linear single-index models with partly endogenous regressors. Computational Statistics & Data Analysis, 156 (??):Article 107145, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732030236X.

Lee:2021:AML

[254] Keunbaik Lee, Chang-Hoon Lee, Min-Sun Kwak, and Eun Jin Jang. Analysis of multivariate longitudinal data using ARMA Cholesky and hypersphere decompositions. Computational Statistics & Data Analysis, 156(??): Article 107144, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302358.

Wang:2021:DRS

[255] Lei Wang, Puying Zhao, and Jun Shao. Dimension-reduced semiparametric estimation of distribution functions and quantiles with nonignorable nonresponse. Computational Statistics & Data Analysis, 156(??):Article 107142, April 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302334.

Wang:2021:CNL

[256] Bingling Wang and Qing Zhou. Causal network learning with non-invertible functional relationships. Computational Statistics & Data Analysis, 156(??): Article 107141, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320302322.

Freund:2021:IGD

[257] Fabian Freund and Arno Siri-Jégousse. The impact of genetic diversity statistics on model selection between coalescents. Computational Statistics & Data Analysis, 156(??):Article 107055, April 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301468.

Byrd:2021:BRG

[258] Michael Byrd, Linh H. Nghiem, and Monnie McGee. Bayesian regularization of Gaussian graphical models with measurement error. Computational Statistics & Data Analysis, 156(??): Article 107085, April 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320301766.

Anonymous:2021:Mb

[259] Anonymous. May 2021. Computational Statistics & Data Analysis, 157 (??):??, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2021:EBe

[260] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 157 (??):Article 107196, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732100030X.

Brown:2021:NMM

[261] Paul T. Brown, Chaitanya Joshi, Stephen Joe, and Håvard Rue. A novel method of marginalisation using low discrepancy sequences for integrated nested Laplace approximations. Computational Statistics & Data Analysis, 157(??): Article 107147, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302383.

Wiqvist:2021:EIS

[262] Samuel Wiqvist, Andrew Golightly, Ashleigh T. McLean, and Umberto Picchini. Efficient inference for stochastic differential equation mixed-effects models using correlated particle pseudomarginal algorithms. Computational Statistics & Data Analysis, 157(??): Article 107151, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302425.

Zhou:2021:CED

[263] Ping Zhou, Zhen Yu, Jingyi Ma, Maozai Tian, and Ye Fan. Communicationefficient distributed estimator for generalized linear models with a diverging number of covariates. Computational Statistics & Data Analysis, 157 (??):Article 107154, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302450.

Hashemi:2021:FFA

[264] Farzane Hashemi, Mehrdad Naderi, Ahad Jamalizadeh, and Andriette Bekker. A flexible factor analysis based on the class of mean-mixture of normal distributions. Computational Statistics & Data Analysis, 157(??): Article 107162, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732030253X.

Bansal:2021:FBE

[265] Prateek Bansal, Rico Krueger, and Daniel J. Graham. Fast Bayesian estimation of spatial count data models. Computational Statistics & Data Analysis, 157(??):Article 107152, May 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302437.

Yuan:2021:CDE

[266] Quan Yuan and Binghui Liu. Community detection via an efficient nonconvex optimization approach based on modularity. Computational Statistics & Data Analysis, 157(??):Article 107163, May 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302541.

Dyckerhoff:2021:ACP

[267] Rainer Dyckerhoff, Pavlo Mozharovskyi, and Stanislav Nagy. Approximate computation of projection depths. Computational Statistics & Data Analysis, 157 (??):Article 107166, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302577.

Hintz:2021:NVM

[268] Erik Hintz, Marius Hofert, and Christiane Lemieux. Normal variance mixtures: Distribution, density and parameter estimation. Computational Statistics & Data Analysis, 157(??): Article 107175, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000098.

Liebscher:2021:KRC

[269] Eckhard Liebscher. Kendall regression coefficient. Computational Statistics & Data Analysis, 157(??):Article 107140, May 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302310.

Fanjul-Hevia:2021:NPT

[270] Arís Fanjul-Hevia, Wenceslao González-Manteiga, and Juan Carlos Pardo-Fernández. A non-parametric test for comparing conditional ROC curves. Computational Statistics & Data Analysis, 157(??):Article 107146, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0167947320302371.

Manju:2021:RCE

[271] Md Abu Manju, Math J. J. M. Candel, and Gerard J. P. van Breukelen. Robustness of cost-effectiveness analyses of cluster randomized trials assuming bivariate normality against skewed cost data. Computational Statistics & Data Analysis, 157(??): Article 107143, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302346.

Gijbels:2021:SQR

[272] Irène Gijbels, Rezaul Karim, and Anneleen Verhasselt. Semiparametric quantile regression using family of quantile-based asymmetric densities. Computational Statistics & Data Analysis, 157(??):Article 107129, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947320302206.

Peng:2021:FIS

[273] Heng Peng, Chuanlong Xie, and Jingxin Zhao. Fast inference for semi-varying coefficient models via local averaging. Computational Statistics & Data Analysis, 157(??):Article 107126, May 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302176.

Huang:2021:CDC

[274] Hengzhen Huang and Xueping Chen.

Compromise design for combination experiment of two drugs. Computational Statistics & Data Analysis, 157(??): Article 107150, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302413.

Mishra:2021:GCS

[275] Aditya Mishra, Dipak K. Dey, Yong Chen, and Kun Chen. Generalized co-sparse factor regression. Computational Statistics & Data Analysis, 157 (??):Article 107127, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302188.

Xiao:2021:MBU

[276] Qian Xiao and Hongquan Xu. A mapping-based universal kriging model for order-of-addition experiments in drug combination studies. Computational Statistics & Data Analysis, 157 (??):Article 107155, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302462.

Wu:2021:EHD

[277] Jianhong Wu. Estimation of high dimensional factor model with multiple threshold-type regime shifts. Computational Statistics & Data Analysis, 157 (??):Article 107153, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320302449.

Qiu:2021:TSTa

[278] Zhiping Qiu, Jianwei Chen, and Jin-Ting Zhang. Two-sample tests for multivariate functional data with applications. Computational Statistics & Data Analysis, 157(??):Article 107160, May 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320302516.

Shi:2021:SEP

[279] Jianhong Shi, Yujing Zhang, Ping Yu, and Weixing Song. SIMEX estimation in parametric modal regression with measurement error. Computational Statistics & Data Analysis, 157 (??):Article 107158, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320302498.

Fitzpatrick:2021:SVS

[280] Dylan Fitzpatrick, Yun Ni, and Daniel B. Neill. Support vector subset scan for spatial pattern detection. Computational Statistics & Data Analysis, 157(??):Article 107149, May 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320302401.

Zhang:2021:GBN

[281] Hongmei Zhang, Xianzheng Huang, Shengtong Han, Faisal I. Rezwan, Wilfried Karmaus, Hasan Arshad, and John W. Holloway. Gaussian Bayesian network comparisons with graph ordering unknown. Computational Statistics & Data Analysis, 157(??):Article 107156, May 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302474.

Du:2021:RAC

[282] Mingyue Du, Huiqiong Li, and Jianguo Sun. Regression analysis of censored data with nonignorable missing covariates and application to Alzheimer Disease. Computational Statistics & Data Analysis, 157(??): Article 107157, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302486.

Sundararajan:2021:PCA

[283] Raanju R. Sundararajan. Principal component analysis using frequency components of multivariate time series. Computational Statistics & Data Analysis, 157(??):Article 107164, May 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302553.

Sun:2021:RAA

[284] Dayu Sun, Hui Zhao, and Jianguo Sun. Regression analysis of asynchronous longitudinal data with informative observation processes. Computational Statistics & Data Analysis, 157 (??):Article 107161, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320302528.

Lu:2021:PBF

[285] Jun Lu, Lu Lin, and WenWu Wang. Partition-based feature screening for categorical data via RKHS embeddings. Computational Statistics & Data Analysis, 157(??):Article 107176, May 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000104.

UlHassan:2021:EAO

[286] Mahmood Ul Hassan and Frank Miller. An exchange algorithm for optimal calibration of items in computerized achievement tests. Computational Statistics & Data Analysis, 157(??): Article 107177, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000116.

Fermanian:2021:ELS

[287] Adeline Fermanian. Embedding and learning with signatures. Computational Statistics & Data Analysis, 157 (??):Article 107148, May 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320302395.

Cao:2021:SKP

[288] Jian Cao, Marc G. Genton, David E. Keyes, and George M. Turkiyyah. Sum of Kronecker products representation and its Cholesky factorization for spatial covariance matrices from large grids. Computational Statistics & Data Analysis, 157(??):Article 107165, May 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947320302565.

Anonymous:2021:Jb

[289] Anonymous. June 2021. Computational Statistics & Data Analysis, 158 (??):??, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2021:EBf

[290] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 158 (??):Article 107214, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000487.

Mirfarah:2021:MLE

[291] Elham Mirfarah, Mehrdad Naderi, and Ding-Geng Chen. Mixture of linear experts model for censored data: a novel approach with scale-mixture of normal distributions. Computational Statistics & Data Analysis, 158(??): Article 107182, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000165.

Gangloff:2021:UIS

[292] Hugo Gangloff, Jean-Baptiste Courbot, Emmanuel Monfrini, and Christophe Collet. Unsupervised image segmentation with Gaussian Pairwise Markov Fields. Computational Statistics & Data Analysis, 158(??):Article 107178, June 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000128.

Cappozzo:2021:RVS

[293] Andrea Cappozzo, Francesca Greselin, and Thomas Brendan Murphy. Robust variable selection for model-based learning in presence of adulteration. Computational Statistics & Data Analysis, 158 (??):Article 107186, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000207.

Luati:2021:EDH

[294] Alessandra Luati and Marco Novelli. Explicit-duration Hidden Markov Models for quantum state estimation. Computational Statistics & Data Analysis, 158(??):Article 107183, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947321000177.

Bouchouia:2021:HDR

[295] Mohammed Bouchouia and François Portier. High dimensional regression for regenerative time-series: an application to road traffic modeling. Computational Statistics & Data Analysis, 158 (??):Article 107191, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000256.

Chabert-Liddell:2021:SBM

[296] Saint-Clair Chabert-Liddell, Pierre Barbillon, Sophie Donnet, and Emmanuel Lazega. A stochastic block model approach for the analysis of multilevel networks: an application to the sociology of organizations. Computational Statistics & Data Analysis, 158(??): Article 107179, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732100013X.

Zuo:2021:CPR

[297] Yijun Zuo. Computation of projection regression depth and its induced median. Computational Statistics & Data Analysis, 158(??):Article 107184, June 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000189.

Lee:2021:VSF

[298] Kuo-Jung Lee, Martin Feldkircher, and Yi-Chi Chen. Variable selection in finite mixture of regression models with an unknown number of components. Computational Statistics & Data Analysis, 158 (??):Article 107180, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000141.

Wiens:2021:RDD

[299] Douglas P. Wiens. Robust designs for dose–response studies: Model and labelling robustness. Computational Statistics & Data Analysis, 158(??):

Article 107189, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000232.

Kasianova:2021:RAD

[300] Ksenia Kasianova, Mark Kelbert, and Pavel Mozgunov. Response adaptive designs for Phase II trials with binary endpoint based on context-dependent information measures. Computational Statistics & Data Analysis, 158(??): Article 107187, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000219.

Silva:2021:CIS

[301] Ivair R. Silva, Luiz Duczmal, and Martin Kulldorff. Confidence intervals for spatial scan statistic. Computational Statistics & Data Analysis, 158(??): Article 107185, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000190.

Li:2021:CFL

[302] Ting Li, Xinyuan Song, Yingying Zhang, Hongtu Zhu, and Zhongyi Zhu. Clusterwise functional linear regression models. Computational Statistics & Data Analysis, 158(??):Article 107192, June 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000268.

Batool:2021:CAS

[303] Fatima Batool and Christian Hennig.

Clustering with the Average Silhouette Width. Computational Statistics & Data Analysis, 158(??):Article 107190, June 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000244.

He:2021:OTR

[304] Yizeng He, Soyoung Kim, Mi-Ok Kim, Wael Saber, and Kwang Woo Ahn. Optimal treatment regimes for competing risk data using doubly robust outcome weighted learning with bilevel variable selection. Computational Statistics & Data Analysis, 158(??): Article 107167, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000013.

Lim:2021:ERS

[305] Alejandro Lim, Chin-Tsang Chiang, and Jen-Chieh Teng. Estimating robot strengths with application to selection of alliance members in FIRST robotics competitions. Computational Statistics & Data Analysis, 158(??): Article 107181, June 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000153.

Ueki:2021:TCM

[306] Masao Ueki and Alzheimerś Disease Neuroimaging Initiative. Testing conditional mean through regression model sequence using Yanai’s generalized coefficient of determination. Computational Statistics & Data Analysis, 158(??):Article 107168, June 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000025.

Anonymous:2021:Jc

[307] Anonymous. July 2021. Computational Statistics & Data Analysis, 159 (??):??, July 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2021:EBg

[308] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 159 (??):Article 107236, July 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000700.

Allassonniere:2021:NCS

[309] Stéphanie Allassonnière and Juliette Chevallier. A new class of stochastic EM algorithms. Escaping local maxima and handling intractable sampling. Computational Statistics & Data Analysis, 159 (??):Article 107159, July 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947320302504.

Zhong:2021:CMV

[310] Wei Zhong, Jiping Wang, and Xiaolin Chen. Censored mean variance sure independence screening for ultrahigh dimensional survival data. Computational Statistics & Data Analysis, 159 (??):Article 107206, July 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000402.

Zhang:2021:SCE

[311] Yuyang Zhang, Patrick Schnell, Chi Song, Bin Huang, and Bo Lu. Subgroup causal effect identification and estimation via matching tree. Computational Statistics & Data Analysis, 159 (??):Article 107188, July 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000220.

Huang:2021:TFR

[312] Shih-Ting Huang, Fang Xie, and Johannes Lederer. Tuning-free ridge estimators for high-dimensional generalized linear models. Computational Statistics & Data Analysis, 159(??): Article 107205, July 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000396.

Fuchs:2021:DFR

[313] Sebastian Fuchs, F. Marta L. Di Lascio, and Fabrizio Durante. Dissimilarity functions for rank-invariant hierarchical clustering of continuous variables. Computational Statistics & Data Analysis, 159(??):Article 107201, July 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000359.

Zhang:2021:GMG

[314] Tonglin Zhang and Ge Lin. Generalized k-means in GLMs with applications to the outbreak of COVID-19 in the United States. Computational Statistics & Data Analysis, 159(??): Article 107217, July 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000517.

Zhang:2021:PSC

[315] Qingzhao Zhang, Shuangge Ma, and Yuan Huang. Promote sign consistency in the joint estimation of precision matrices. Computational Statistics & Data Analysis, 159(??):Article 107210, July 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732100044X.

Maruotti:2021:HSM

[316] Antonello Maruotti, Lea Petrella, and Luca Sposito. Hidden semi-Markov-switching quantile regression for time series. Computational Statistics & Data Analysis, 159(??):Article 107208, July 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000426.

Park:2021:HTV

[317] Seyoung Park and Eun Ryung Lee. Hypothesis testing of varying coefficients for regional quantiles. Computational Statistics & Data Analysis, 159(??): Article 107204, July 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000384.

Kirkby:2021:NDE

[318] J. Lars Kirkby, Álvaro Leitao, and Duy Nguyen. Nonparametric density estimation and bandwidth selection with B-spline bases: a novel Galerkin method. Computational Statistics & Data Analysis, 159(??):Article 107202, July 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000360.

Huang:2021:TDG

[319] Xianzheng Huang and Hongmei Zhang.

Tests for differential Gaussian Bayesian networks based on quadratic inference functions. Computational Statistics & Data Analysis, 159(??):Article 107209, July 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000438.

Kang:2021:ESE

[320] Xiaoning Kang and Mingqiu Wang.

Ensemble sparse estimation of covariance structure for exploring genetic disease data. Computational Statistics & Data Analysis, 159(??):Article 107220, July 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000542.

Li:2021:DDR

[321] Rui Li, Brian J. Reich, and Howard D. Bondell. Deep distribution regression. Computational Statistics & Data Analysis, 159(??):Article 107203, July 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000372.

Anonymous:2021:Ab

[322] Anonymous. August 2021. Computational Statistics & Data Analysis, 160 (??):??, August 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2021:EBh

[323] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 160(??):??, August 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000918.

Guadarrama:2021:TSE

[324] María Guadarrama, Domingo Morales, and Isabel Molina. Time stable empirical best predictors under a unit-level model. Computational Statistics & Data Analysis, 160(??): ??, August 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000608.

Kim:2021:PSN

[325] Nam-Hwui Kim and Ryan P. Browne. In the pursuit of sparseness: a new rank-preserving penalty for a finite mixture of factor analyzers. Computational Statistics & Data Analysis, 160(??):??, August 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000785.

Reynolds:2021:LAG

[326] David Reynolds and Luis Carvalho.

Latent association graph inference for binary transaction data. Computational Statistics & Data Analysis, 160 (??):??, August 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000633.

Wang:2021:RDM

[327] Kangning Wang and Shaomin Li.

Robust distributed modal regression for massive data. Computational Statistics & Data Analysis, 160(??): ??, August 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000591.

Jin:2021:RTT

[328] Lei Jin. Robust tests for time series comparison based on Laplace periodograms. Computational Statistics & Data Analysis, 160(??):??, August 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000578.

Zheng:2021:FDV

[329] Nan Zheng and Noel Cadigan. Frequentist delta-variance approximations with mixed-effects models and TMB. Computational Statistics & Data Analysis, 160(??):??, August 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732100061X.

Qiu:2021:TSTb

[330] Tao Qiu, Wangli Xu, and Liping Zhu. Two-sample test in high dimensions through random selection. Computational Statistics & Data Analysis, 160(??):??, August 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000529.

Guo:2021:CQR

[331] Chaohui Guo, Jing Lv, and Jibo Wu. Composite quantile regression for ultra-high dimensional semiparametric model averaging. Computational Statistics & Data Analysis, 160 (??):??, August 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000657.

Wei:2021:BCK

[332] Yuting Wei, Qihua Wang, Xiaogang Duan, and Jing Qin. Biascorrected Kullback–Leibler distance criterion based model selection with covariables missing at random. Computational Statistics & Data Analysis, 160(??):??, August 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732100058X.

Dong:2021:PIL

[333] Ruipeng Dong, Daoji Li, and Zemin Zheng. Parallel integrative learning for large-scale multi-response regression with incomplete outcomes. Computational Statistics & Data Analysis, 160(??):??, August 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000773.

Liang:2021:MFD

[334] Weijuan Liang, Shuangge Ma, and Cunjie Lin. Marginal false discovery rate for a penalized transformation survival model. Computational Statistics & Data Analysis, 160 (??):??, August 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000669.

Zhang:2021:RMT

[335] Jing Zhang, Qin Wang, and D’Arcy Mays. Robust MAVE through nonconvex penalized regression. Computational Statistics & Data Analysis, 160(??):??, August 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000815.

Lai:2021:KBM

[336] Tingyu Lai, Zhongzhan Zhang, and Yafei Wang. A kernel-based measure for conditional mean dependence. Computational Statistics & Data Analysis, 160(??):??, August 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000803.

Galvani:2021:FNB

[337] Marta Galvani, Agostino Torti, Alessandra Menafoglio, and Simone Vantini. FunCC: a new bi-clustering algorithm for functional data with misalignment. Computational Statistics & Data Analysis, 160(??):??, August 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000530.

Anonymous:2021:S

[338] Anonymous. September 2021. Computational Statistics & Data Analysis, 161 (??):??, September 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2021:EBi

[339] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 161(??):??, September 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001055.

Jones:2021:BLA

[340] Matthew Jones, Michael Goldstein, David Randell, and Philip Jonathan. Bayes linear analysis for ordinary differential equations. Computational Statistics & Data Analysis, 161(??):

??, September 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000621.

Rodriguez:2021:CPF

[341] Carlos E. Rodríguez and Stephen G.

Walker. Copula particle filters. Computational Statistics & Data Analysis, 161(??):??, September 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000645.

Ghorbani:2021:TFO

[342] Mohammad Ghorbani, Nafiseh Vafaei, Jiří Dvořák, and Mari Myllymäki. Testing the first-order separability hypothesis for spatio-temporal point patterns. Computational Statistics & Data Analysis, 161(??):??, September 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000797.

Wang:2021:EIM

[343] Qin Wang and Yuan Xue. An ensemble of inverse moment estimators for sufficient dimension reduction. Computational Statistics & Data Analysis, 161(??):??, September 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732100075X.

Liu:2021:GAH

[344] Xiaoyu Liu and Liming Xiang. Generalized accelerated hazards mixture cure models with interval-censored data. Computational Statistics & Data Analysis, 161(??):??, September 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947321000827.

Kakizawa:2021:CBS

[345] Yoshihide Kakizawa. A class of Birnbaum–Saunders type kernel density estimators for nonnegative data. Computational Statistics & Data Analysis, 161(??):??, September 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000839.

Fan:2021:TEH

[346] Caiyun Fan, Wenbin Lu, and Yong Zhou. Testing error heterogeneity in censored linear regression. Computational Statistics & Data Analysis, 161 (??):??, September 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000414.

Lambert:2021:FBI

[347] Philippe Lambert. Fast Bayesian inference using Laplace approximations in nonparametric double additive location-scale models with rightand interval-censored data. Computational Statistics & Data Analysis, 161(??):??, September 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000840.

Wang:2021:CHS

[348] Craig Wang, Reinhard Furrer, and SNC Study Group. Combining heterogeneous spatial datasets with processbased spatial fusion models: a unifying framework. Computational Statistics & Data Analysis, 161(??): ??, September 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000748.

Ahfock:2021:HLN

[349] Daniel Ahfock and Geoffrey J. McLachlan. Harmless label noise and informative soft-labels in supervised classification. Computational Statistics & Data Analysis, 161(??):??, September 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000876.

Donelli:2021:BSV

[350] Nicola Donelli, Stefano Peluso, and Antonietta Mira. A Bayesian semiparametric vector Multiplicative Error Model. Computational Statistics & Data Analysis, 161(??):??, September 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000761.

Shi:2021:CED

[351] Jianwei Shi, Guoyou Qin, Huichen Zhu, and Zhongyi Zhu. Communicationefficient distributed M-estimation with missing data. Computational Statistics & Data Analysis, 161(??):??, September 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000852.

Lee:2021:BML

[352] Jung Wun Lee, Hwan Chung, and Saebom Jeon. Bayesian multivariate latent class profile analysis: Exploring the developmental progression of youth depression and substance use. Computational Statistics & Data Analysis, 161(??):??, September 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000955.

Wang:2021:RCE

[353] Kangning Wang, Shaomin Li, and Benle Zhang. Robust communicationefficient distributed composite quantile regression and variable selection for massive data. Computational Statistics & Data Analysis, 161(??): ??, September 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000967.

Anonymous:2021:O

[354] Anonymous. October 2021. Computational Statistics & Data Analysis, 162 (??):??, October 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2021:EBj

[355] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 162(??):??, October 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001250.

Polansky:2021:MBP

[356] Alan M. Polansky and Paramahansa Pramanik. A motif building process for simulating random networks. Computational Statistics & Data Analysis, 162(??):??, October 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000979.

Ferreira:2021:FSC

[357] Marco A. R. Ferreira, Erica M.

Porter, and Christopher T. Franck. Fast and scalable computations for Gaussian hierarchical models with intrinsic conditional autoregressive spatial random effects. Computational Statistics & Data Analysis, 162(??): ??, October 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321000980.

Kang:2021:FJA

[358] Yicheng Kang, Yueyong Shi, Yuling Jiao, Wendong Li, and Dongdong Xiang. Fitting jump additive models. Computational Statistics & Data Analysis, 162(??):??, October 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001006.

Langrene:2021:FME

[359] Nicolas Langrené and Xavier Warin. Fast multivariate empirical cumulative distribution function with connection to kernel density estimation. Computational Statistics & Data Analysis, 162(??):??, October 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001018.

Im:2021:BSA

[360] Yunju Im and Aixin Tan. Bayesian subgroup analysis in regression using mixture models. Computational Statistics & Data Analysis, 162(??): ??, October 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000864.

Wang:2021:DOS

[361] Feifei Wang, Yingqiu Zhu, Danyang Huang, Haobo Qi, and Hansheng Wang. Distributed one-step upgraded estimation for non-uniformly and nonrandomly distributed data. Computational Statistics & Data Analysis, 162(??):??, October 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321000992.

Acosta:2021:AES

[362] Jonathan Acosta, Alfredo Alegría, Felipe Osorio, and Ronny Vallejos. Assessing the effective sample size for large spatial datasets: a block likelihood approach. Computational Statistics & Data Analysis, 162(??): ??, October 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732100116X.

Anonymous:2021:N

[363] Anonymous. November 2021. Computational Statistics & Data Analysis, 163 (??):??, November 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2021:EBk

[364] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 163(??):??, November 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001468.

Dumbgen:2021:ASA

[365] Lutz Dümbgen, Alexandre Mösching, and Christof Strähl. Active set algorithms for estimating shape-constrained density ratios. Computational Statistics & Data Analysis, 163(??):??, November 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001341.

Wang:2021:FFE

[366] Pei Wang, Xiangrong Yin, Qingcong Yuan, and Richard Kryscio. Feature filter for estimating central mean subspace and its sparse solution. Computational Statistics & Data Analysis, 163(??):??, November 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001195.

Ghosh:2021:TSH

[367] Santu Ghosh, Deepak Nag Ayyala, and Rafael Hellebuyck. Two-sample high dimensional mean test based on prepivots. Computational Statistics & Data Analysis, 163(??):??, November 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001183.

Tucker:2021:MBR

[368] J. Derek Tucker, Lyndsay Shand, and Kenny Chowdhary. Multimodal Bayesian registration of noisy functions using Hamiltonian Monte Carlo. Computational Statistics & Data Analysis, 163(??):??, November 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001328.

Rodwell:2021:CCB

[369] D. T. Rodwell, C. J. van der Merwe, and S. Gardner-Lubbe. Categorical CVA biplots. Computational Statistics & Data Analysis, 163(??): ??, November 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732100133X.

Zhang:2021:CBF

[370] Xiaoke Zhang, Wu Xue, and Qiyue Wang. Covariate balancing functional propensity score for functional treatments in cross-sectional observational studies. Computational Statistics & Data Analysis, 163(??): ??, November 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001377.

Jimenez:2021:ADE

[371] Johnatan Cardona Jiménez and Carlos A. de B. Pereira. Assessing dynamic effects on a Bayesian matrix-variate dynamic linear model: an application to task-based fMRI data analysis. Computational Statistics & Data Analysis, 163(??):??, November 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001316.

Anonymous:2021:D

[372] Anonymous. December 2021. Computational Statistics & Data Analysis, 164 (??):??, December 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2021:EBl

[373] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 164(??):??, December 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732100164X.

Mao:2021:CTS

[374] Shanjun Mao, Xiaodan Fan, and Jie Hu. Correlation for tree-shaped datasets and its Bayesian estimation. Computational Statistics & Data Analysis, 164(??):??, December 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001419.

Pircalabelu:2021:GIS

[375] Eugen Pircalabelu and Andreas Artemiou. Graph informed sliced inverse regression. Computational Statistics & Data Analysis, 164(??):??, December 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947321001365.

Castelletti:2021:ECS

[376] Federico Castelletti and Stefano Peluso. Equivalence class selection of categorical graphical models. Computational Statistics & Data Analysis, 164 (??):??, December 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001389.

Singh:2021:EEE

[377] Satya Prakash Singh and Ori Davidov. On efficient exact experimental designs for ordered treatments. Computational Statistics & Data Analysis, 164(??):??, December 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001390.

Choi:2021:SLS

[378] Taehwa Choi, Arlene K. H. Kim, and Sangbum Choi. Semiparametric least-squares regression with doubly-censored data. Computational Statistics & Data Analysis, 164(??): ??, December 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001407.

Gaucher:2021:ODN

[379] Solenne Gaucher, Olga Klopp, and Geneviève Robin. Outlier detection in networks with missing links. Computational Statistics & Data Analysis, 164(??):??, December 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001420.

Xu:2021:PAB

[380] Kai Xu and Yeqing Zhou. Projectionaveraging-based cumulative covariance and its use in goodness-of-fit testing for single-index models. Computational Statistics & Data Analysis, 164(??):??, December 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001353.

Narci:2021:IPO

[381] Romain Narci, Maud Delattre, Catherine Larédo, and Elisabeta Vergu. Inference for partially observed epidemic dynamics guided by Kalman filtering techniques. Computational Statistics & Data Analysis, 164(??): ??, December 2021. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001535.

Zhang:2021:MPT

[382] Huaiyu Zhang and Haiyan Wang. A more powerful test of equality of highdimensional two-sample means. Computational Statistics & Data Analysis, 164(??):??, December 2021. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001523.

Anonymous:2022:Ja

[383] Anonymous. January 2022. Computational Statistics & Data Analysis, 165 (??):??, January 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2022:EBa

[384] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 165(??):??, January 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001742.

Zhao:2022:ODO

[385] Yuna Zhao, Dennis K. J. Lin, and Min-Qian Liu. Optimal designs for order-of-addition experiments. Computational Statistics & Data Analysis, 165(??):??, January 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001547.

Park:2022:BMS

[386] Jaewoo Park, Ick Hoon Jin, and Michael Schweinberger. Bayesian model selection for high-dimensional Ising models, with applications to educational data. Computational Statistics & Data Analysis, 165(??):??, January 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001596.

Kelter:2022:PAT

[387] Riko Kelter. Power analysis and type I and type II error rates of Bayesian nonparametric two-sample tests for location-shifts based on the Bayes factor under Cauchy priors. Computational Statistics & Data Analysis, 165(??):??, January 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001602.

Mishra:2022:RRC

[388] Aditya Mishra and Christian L.

Müller. Robust regression with compositional covariates. Computational Statistics & Data Analysis, 165(??): ??, January 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001493.

Kim:2022:NCT

[389] Ahhyoun Kim and Hyunjoong Kim. A new classification tree method with interaction detection capability. Computational Statistics & Data Analysis, 165(??):??, January 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001584.

Li:2022:BMP

[390] Muyi Li and Yanfen Zhang. Bootstrapping multivariate portmanteau tests for vector autoregressive models with weak assumptions on errors. Computational Statistics & Data Analysis, 165(??):??, January 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001559.

Jia:2022:DLQ

[391] Yichen Jia and Jong-Hyeon Jeong. Deep learning for quantile regression under right censoring: DeepQuantreg. Computational Statistics & Data Analysis, 165(??):??, January 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001572.

Anonymous:2022:F

[392] Anonymous. February 2022. Computational Statistics & Data Analysis, 166 (??):??, February 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2022:EBb

[393] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 166(??):??, February 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001900.

Ghosal:2022:BIG

[394] Rahul Ghosal and Sujit K. Ghosh. Bayesian inference for generalized linear model with linear inequality constraints. Computational Statistics & Data Analysis, 166(??):??, February 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001699.

Wang:2022:LLP

[395] Yijun Wang, Weiwei Wang, and Xiaobing Zhao. Local logarithm partial likelihood estimation of panel count data model with an unknown link function. Computational Statistics & Data Analysis, 166(??):??, February 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001808.

Kim:2022:PGM

[396] Kyongwon Kim. On principal graphical models with application to gene network. Computational Statistics & Data Analysis, 166(??): ??, February 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732100178X.

Shin:2022:JEM

[397] Yei Eun Shin, Lan Zhou, and Yu Ding.

Joint estimation of monotone curves via functional principal component analysis. Computational Statistics & Data Analysis, 166(??):??, February 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001778.

Xing:2022:MBS

[398] Xin Xing, Rui Xie, and Wenxuan Zhong. Model-based sparse coding beyond Gaussian independent model. Computational Statistics & Data Analysis, 166(??):??, February 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001705.

Anonymous:2022:Ma

[399] Anonymous. March 2022. Computational Statistics & Data Analysis, 167 (??):??, March 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2022:EBc

[400] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002280.

Rugamer:2022:SIA

[401] David Rügamer, Philipp F. M. Baumann, and Sonja Greven. Selective inference for additive and linear mixed models. Computational Statistics & Data Analysis, 167(??): ??, March 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001845.

Kruse:2022:MAL

[402] René-Marcel Kruse, Alexander Silbersdorff, and Benjamin Säfken. Model averaging for linear mixed models via augmented Lagrangian. Computational Statistics & Data Analysis, 167 (??):??, March 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001857.

Nattino:2022:PAO

[403] Giovanni Nattino, Chi Song, and Bo Lu. Polymatching algorithm in observational studies with multiple treatment groups. Computational Statistics & Data Analysis, 167 (??):??, March 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001985.

Ollier:2022:FSN

[404] Edouard Ollier. Fast selection of nonlinear mixed effect models using penalized likelihood. Computational Statistics & Data Analysis, 167 (??):??, March 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002073.

Franke:2022:AQC

[405] Jürgen Franke, Mario Hefter, André Herzwurm, Klaus Ritter, and Stefanie Schwaar. Adaptive quantile computation for Brownian bridge in change-point analysis. Computational Statistics & Data Analysis, 167 (??):??, March 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002097.

Smida:2022:WMW

[406] Zaineb Smida, Lionel Cucala, Ali Gannoun, and Ghislain Durif. A Wilcoxon–Mann–Whitney spatial scan statistic for functional data. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002127.

Heuchenne:2022:IMS

[407] Cédric Heuchenne and Alexandre Jacquemain. Inference for monotone single-index conditional means: a Lorenz regression approach. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732100181X.

Zhou:2022:BBR

[408] Haiming Zhou and Xianzheng Huang. Bayesian beta regression for bounded responses with unknown supports. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001791.

Chen:2022:HBM

[409] Jiaxun Chen, Athanasios C. Micheas, and Scott H. Holan. Hierarchical Bayesian modeling of spatio-temporal area-interaction processes. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001833.

Xiao:2022:DRB

[410] Zhen Xiao and Qi Zhang. Dimension reduction for block-missing data based on sparse sliced inverse regression. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001821.

Wang:2022:OEE

[411] Jiangyan Wang, Lijie Gu, and Lijian Yang. Oracle-efficient estimation for functional data error distribution with simultaneous confidence band. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001973.

Conti:2022:EUP

[412] Pier Luigi Conti, Fulvia Mecatti, and Federica Nicolussi. Efficient unequal probability resampling from finite populations. Computational Statistics & Data Analysis, 167(??): ??, March 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002000.

Boland:2022:SLT

[413] Joanna Boland, Donatello Telesca, Catherine Sugar, Shafali Jeste, Cameron Goldbeck, and Damla Senturk. A study of longitudinal trends in timefrequency transformations of EEG data during a learning experiment. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002012.

Yang:2022:MAH

[414] Qi Yang, Haijin He, Bin Lu, and Xinyuan Song. Mixture additive hazards cure model with latent variables: Application to corporate default data. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001997.

Pavithra:2022:PEC

[415] Celeste R. Pavithra and T. G. Deepak. Parameter estimation and computation of the Fisher information matrix for functions of phase type random variables. Computational Statistics & Data Analysis, 167 (??):??, March 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001961.

Brunet-Saumard:2022:KBR

[416] Camille Brunet-Saumard, Edouard Genetay, and Adrien Saumard. KbMOM: a robust Lloyd-type clustering algorithm based on bootstrap median-of-means. Computational Statistics & Data Analysis, 167 (??):??, March 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002048.

Liu:2022:PQC

[417] Jicai Liu, Yuefeng Si, Yong Niu, and Riquan Zhang. Projection quantile correlation and its use in high-dimensional grouped variable screening. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002036.

Wu:2022:UQH

[418] Suofei Wu, Jan Hannig, and Thomas C. M. Lee. Uncertainty quantification for honest regression trees. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002115.

Zhang:2022:GGF

[419] Hong Zhang and Zheyang Wu. The general goodness-of-fit tests for correlated data. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947321002139.

Samaddar:2022:GCF

[420] Arunava Samaddar, Brooke S. Jackson, Christopher J. Helms, Nicole A. Lazar, Jennifer E. McDowell, and Cheolwoo Park. A group comparison in fMRI data using a semiparametric model under shape invariance. Computational Statistics & Data Analysis, 167(??):??, March 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732100195X.

Anonymous:2022:Aa

[421] Anonymous. April 2022. Computational Statistics & Data Analysis, 168 (??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2022:EBd

[422] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732200007X.

Wiemann:2022:CSS

[423] Paul F. V. Wiemann, Nadja Klein, and Thomas Kneib. Correcting for sample selection bias in Bayesian distributional regression models. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002164.

Rios:2022:AOO

[424] Nicholas Rios, Peter Winker, and Dennis K. J. Lin. TA algorithms for Doptimal OofA Mixture designs. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002450.

Battauz:2022:LBB

[425] Michela Battauz and Paolo Vidoni. A likelihood-based boosting algorithm for factor analysis models with binary data. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002462.

Coube-Sisqueille:2022:IPM

[426] Sébastien Coube-Sisqueille and Benoît Liquet. Improving performances of MCMC for Nearest Neighbor Gaussian Process models with full data augmentation. Computational Statistics & Data Analysis, 168(??): ??, April 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002024.

Weng:2022:FTS

[427] Jiaying Weng. Fourier transform sparse inverse regression estimators for sufficient variable selection. Computational Statistics & Data Analysis, 168 (??):??, April 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002140.

Luo:2022:CGE

[428] Renwen Luo and Jianxin Pan. Conditional generalized estimating equations of mean-variance-correlation for clustered data. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947321002206.

Zhang:2022:NNR

[429] Jin-Ting Zhang and Tianming Zhu. A new normal reference test for linear hypothesis testing in high-dimensional heteroscedastic one-way MANOVA. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732100219X.

Weiss:2022:NPA

[430] Christian H. Weiß, Manuel Ruiz Marín, Karsten Keller, and Mariano Matilla-García. Non-parametric analysis of serial dependence in time series using ordinal patterns. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002152.

Golovkine:2022:CMF

[431] Steven Golovkine, Nicolas Klutchnikoff, and Valentin Patilea. Clustering multivariate functional data using unsupervised binary trees. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002103.

Liu:2022:CMC

[432] Jie Liu and Huilin Ge. Collaboration mechanisms and community detection of statisticians based on ERGMs and k NN-walktrap. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002061.

Su:2022:CPS

[433] Miaomiao Su and Qihua Wang. A convex programming solution based debiased estimator for quantile with missing response and high-dimensional covariables. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S016794732100205X.

Deng:2022:SSB

[434] Jianqiu Deng, Xiaojie Yang, and Qihua Wang. Surrogate space based dimension reduction for nonignorable nonresponse. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002085.

Ahfock:2022:SFM

[435] Daniel Ahfock, Saumyadipta Pyne, and Geoffrey J. McLachlan. Statistical file-matching of non-Gaussian data: a game theoretic approach. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002218.

Wang:2022:SHD

[436] Yue Wang, Yan Zhou, Rui Li, and Heng Lian. Sparse high-dimensional semi-nonparametric quantile regression in a reproducing kernel Hilbert space. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732100222X.

Wei:2022:DRE

[437] Kecheng Wei, Guoyou Qin, Jiajia Zhang, and Xuemei Sui. Doubly robust estimation in causal inference with missing outcomes: With an application to the Aerobics Center Longitudinal Study. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002334.

Kapla:2022:FSD

[438] Daniel Kapla, Lukas Fertl, and Efstathia Bura. Fusing sufficient dimension reduction with neural networks. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002243.

Ning:2022:CIT

[439] Jing Ning, Daewoo Pak, Hong Zhu, and Jing Qin. Conditional independence test of failure and truncation times: Essential tool for method selection. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732100236X.

Villa:2022:OBF

[440] Cristiano Villa and Stephen G. Walker.

An objective Bayes factor with improper priors. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947321002383.

Agarwal:2022:FQC

[441] Gaurav Agarwal, Wei Tu, Ying Sun, and Linglong Kong. Flexible quantile contour estimation for multivariate functional data: Beyond convexity. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002346.

Ameijeiras-Alonso:2022:FTP

[442] Jose Ameijeiras-Alonso, Irène Gijbels, and Anneleen Verhasselt. On a family of two-piece circular distributions. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002371.

Szarek:2022:STA

[443] Dawid Szarek, Katarzyna Maraj-Zygma̧t, Grzegorz Sikora, Diego Krapf, and Agnieszka Wyłomańska. Statistical test for anomalous diffusion based on empirical anomaly measure for Gaussian processes. Computational Statistics & Data Analysis, 168 (??):??, April 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002358.

Kirkby:2022:MLE

[444] J. L. Kirkby, Dang H. Nguyen, Duy Nguyen, and Nhu N. Nguyen. Maximum likelihood estimation of diffusions by continuous time Markov chain. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002425.

Zhu:2022:CIP

[445] Ke Zhu and Hanzhong Liu. Confidence intervals for parameters in high-dimensional sparse vector autoregression. Computational Statistics & Data Analysis, 168(??):??, April 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947321002176.

Anonymous:2022:Mb

[446] Anonymous. May 2022. Computational Statistics & Data Analysis, 169 (??):??, May 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2022:EBe

[447] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 169(??):??, May 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000238.

Lai:2022:VBI

[448] Wei-Ting Lai, Ray-Bing Chen, Ying Chen, and Thorsten Koch. Variational Bayesian inference for network autoregression models. Computational Statistics & Data Analysis, 169(??):??, May 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002401.

Thompson:2022:RSS

[449] Ryan Thompson. Robust subset selection. Computational Statistics & Data Analysis, 169(??):??, May 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947321002498.

Hamel:2022:CQS

[450] Andreas H. Hamel and Daniel Kostner. Computation of quantile sets for bivariate ordered data. Computational Statistics & Data Analysis, 169(??):??, May 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000020.

DeGooijer:2022:KBH

[451] Jan G. De Gooijer, Gustav Eje Henter, and Ao Yuan. Kernel-based hidden Markov conditional densities. Computational Statistics & Data Analysis, 169(??):??, May 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000111.

Bigot:2022:LRM

[452] Jérémie Bigot and Charles Deledalle. Low-rank matrix denoising for count data using unbiased Kullback–Leibler risk estimation. Computational Statistics & Data Analysis, 169(??):??, May 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947322000032.

Gamiz:2022:MLS

[453] María Luz Gámiz, Enno Mammen, María Dolores Martínez-Miranda, and Jens Perch Nielsen. Missing link survival analysis with applications to available pandemic data. Computational Statistics & Data Analysis, 169 (??):??, May 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002395.

Zhao:2022:ESV

[454] Yan-Yong Zhao, Jin-Guan Lin, Jian-Qiang Zhao, and Zhang-Xiao Miao. Estimation of semi-varying coefficient models for longitudinal data with irregular error structure. Computational Statistics & Data Analysis, 169 (??):??, May 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002231.

Nibbering:2022:MPL

[455] Didier Nibbering and Trevor J. Hastie.

Multiclass-penalized logistic regression. Computational Statistics & Data Analysis, 169(??):??, May 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002486.

Yang:2022:MSS

[456] Kai Yang, Xinyang Yu, Qingqing Zhang, and Xiaogang Dong. On MCMC sampling in self-exciting integer-valued threshold time series models. Computational Statistics & Data Analysis, 169(??):??, May 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002449.

Kim:2022:STS

[457] Seongho Kim and Weng Kee Wong. Spatial two-stage designs for phase II clinical trials. Computational Statistics & Data Analysis, 169 (??):??, May 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002541.

Luo:2022:DAH

[458] Jiyu Luo, Qiang Sun, and Wen-Xin Zhou. Distributed adaptive Huber regression. Computational Statistics & Data Analysis, 169 (??):??, May 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732100253X.

Liu:2022:SIH

[459] Yang Liu, Wei Sun, Li Hsu, and Qianchuan He. Statistical inference for high-dimensional pathway analysis with multiple responses. Computational Statistics & Data Analysis, 169(??):??, May 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002528.

Zhu:2022:TIC

[460] Qiansheng Zhu and Joseph B. Lang. Test-inversion confidence intervals for estimands in contingency tables subject to equality constraints. Computational Statistics & Data Analysis, 169(??):??, May 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002474.

Sosa:2022:LSM

[461] Juan Sosa and Brenda Betancourt.

A latent space model for multilayer network data. Computational Statistics & Data Analysis, 169 (??):??, May 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000123.

Grazian:2022:ABC

[462] Clara Grazian, Luciana Dalla Valle, and Brunero Liseo. Approximate Bayesian conditional copulas. Computational Statistics & Data Analysis, 169(??):??, May 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002516.

Liu:2022:TEF

[463] Xi Liu, Afshin A. Divani, and Alexander Petersen. Truncated estimation in functional generalized linear regression models. Computational Statistics & Data Analysis, 169 (??):??, May 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000019.

Anonymous:2022:Jcb

[464] Anonymous. June 2022. Computational Statistics & Data Analysis, 170 (??):??, June 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2022:EBf

[465] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 170(??):??, June 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000457.

Cheng:2022:RHD

[466] Chao Cheng, Xingdong Feng, Jian Huang, Yuling Jiao, and Shuang Zhang. ℓ_{0}-regularized high-dimensional accelerated failure time model. Computational Statistics & Data Analysis, 170(??):??, June 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732200010X.

Murray:2022:FAE

[467] James Murray and Pete Philipson. A fast approximate EM algorithm for joint models of survival and multivariate longitudinal data. Computational Statistics & Data Analysis, 170 (??):??, June 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000184.

Hediger:2022:URF

[468] Simon Hediger, Loris Michel, and Jeffrey Näf. On the use of random forest for two-sample testing. Computational Statistics & Data Analysis, 170(??):??, June 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000159.

Mao:2022:NFS

[469] Xiaojun Mao, Liuhua Peng, and Zhonglei Wang. Nonparametric feature selection by random forests and deep neural networks. Computational Statistics & Data Analysis, 170 (??):??, June 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000160.

Corsini:2022:DOM

[470] Noemi Corsini and Cinzia Viroli.

Dealing with overdispersion in multivariate count data. Computational Statistics & Data Analysis, 170 (??):??, June 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000275.

Rhee:2022:RMM

[471] Anbin Rhee, Min-Sun Kwak, and Keunbaik Lee. Robust modeling of multivariate longitudinal data using modified Cholesky and hypersphere decompositions. Computational Statistics & Data Analysis, 170 (??):??, June 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000196.

Shi:2022:CSM

[472] Xuesheng Shi, Colin Gallagher, Robert Lund, and Rebecca Killick. A comparison of single and multiple changepoint techniques for time series data. Computational Statistics & Data Analysis, 170(??):??, June 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000135.

Santos-Fernandez:2022:BST

[473] Edgar Santos-Fernandez, Jay M. Ver Hoef, Erin E. Peterson, James Mc-Gree, Daniel J. Isaak, and Kerrie Mengersen. Bayesian spatio-temporal models for stream networks. Computational Statistics & Data Analysis, 170(??):??, June 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000263.

Anonymous:2022:J

[474] Anonymous. July 2022. Computational Statistics & Data Analysis, 171 (??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2022:EBg

[475] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000688.

Lai:2022:CSO

[476] Jianfa Lai, Lin-Chen Weng, Xiaoling Peng, and Kai-Tai Fang. Construction of symmetric orthogonal designs with deep Q-network and orthogonal complementary design. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000287.

Rosa:2022:CMV

[477] Samuel Rosa and Radoslav Harman.

Computing minimum-volume enclosing ellipsoids for large datasets. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000329.

You:2022:PEM

[478] Kisung You and Changhee Suh. Parameter estimation and model-based clustering with spherical normal distribution on the unit hypersphere. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000378.

Marino:2022:HML

[479] Maria Francesca Marino and Silvia Pandolfi. Hybrid maximum likelihood inference for stochastic block models. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000299.

Palm:2022:DRA

[480] Bruna G. Palm, Fábio M. Bayer, and Renato J. Cintra. 2-D Rayleigh autoregressive moving average model for SAR image modeling. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000330.

Diani:2022:MCW

[481] Cecilia Diani, Giuliano Galimberti, and Gabriele Soffritti. Multivariate clusterweighted models based on seemingly unrelated linear regression. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000317.

Zhao:2022:GLN

[482] Yuxuan Zhao, David S. Matteson, Stewart H. Mostofsky, Mary Beth Nebel, and Benjamin B. Risk. Group linear non-Gaussian component analysis with applications to neuroimaging. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000342.

Sottile:2022:RER

[483] Gianluca Sottile and Paolo Frumento.

Robust estimation and regression with parametric quantile functions. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000512.

Schnurr:2022:GOP

[484] Alexander Schnurr and Svenja Fischer. Generalized ordinal patterns allowing for ties and their applications in hydrology. Computational Statistics & Data Analysis, 171 (??):??, July 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000524.

Zhao:2022:NIT

[485] Yunpeng Zhao. Network inference from temporally dependent grouped observations. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947322000500.

Liu:2022:GCC

[486] Sisheng Liu and Xiaoli Kong. A generalized correlated C_{p} criterion for derivative estimation with dependent errors. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000536.

Zhou:2022:ACE

[487] Jing Zhou, Wei Lan, and Hansheng Wang. Asymptotic covariance estimation by Gaussian random perturbation. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000391.

Hornung:2022:IFI

[488] Roman Hornung and Anne-Laure Boulesteix. Interaction forests: Identifying and exploiting interpretable quantitative and qualitative interaction effects. Computational Statistics & Data Analysis, 171(??):??, July 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000408.

Anonymous:2022:Ab

[489] Anonymous. August 2022. Computational Statistics & Data Analysis, 172 (??):??, August 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2022:EBh

[490] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 172(??):??, August 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000925.

Zhu:2022:RVS

[491] Kailun Zhu and Dorota Kurowicka. Regular vines with strongly chordal pattern of (conditional) independence. Computational Statistics & Data Analysis, 172(??):??, August 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732200041X.

Betancourt:2022:PRL

[492] Brenda Betancourt, Juan Sosa, and Abel Rodríguez. A prior for record linkage based on allelic partitions. Computational Statistics & Data Analysis, 172(??):??, August 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000548.

Pokojovy:2022:RDA

[493] Michael Pokojovy and J. Marcus Jobe. A robust deterministic affineequivariant algorithm for multivariate location and scatter. Computational Statistics & Data Analysis, 172(??):??, August 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732200055X.

Ayub:2022:EEP

[494] Kanwal Ayub, Weixing Song, and Jianhong Shi. Extrapolation estimation in parametric regression models with measurement error. Computational Statistics & Data Analysis, 172(??):??, August 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000585.

Hudecova:2022:MRB

[495] Šárka Hudecová and Miroslav Šiman. Multivariate ranks based on randomized lift-interdirections. Computational Statistics & Data Analysis, 172 (??):??, August 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000603.

Pohle:2022:FES

[496] Jennifer Pohle, Timo Adam, and Larissa T. Beumer. Flexible estimation of the state dwell-time distribution in hidden semi-Markov models. Computational Statistics & Data Analysis, 172(??):??, August 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000597.

Osei:2022:BLM

[497] Prince P. Osei and Ori Davidov. Bayesian linear models for cardinal paired comparison data. Computational Statistics & Data Analysis, 172 (??):??, August 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000615.

Du:2022:VSC

[498] Mingyue Du, Xingqiu Zhao, and Jianguo Sun. Variable selection for case-cohort studies with informatively interval-censored outcomes. Computational Statistics & Data Analysis, 172(??):??, August 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000640.

Florez:2022:EAA

[499] Alvaro J. Flórez, Geert Molenberghs, Wim Van der Elst, and Ariel Alonso Abad. An efficient algorithm to assess multivariate surrogate endpoints in a causal inference framework. Computational Statistics & Data Analysis, 172(??):??, August 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000743.

Qu:2022:VSV

[500] Lianqiang Qu, Xiaoyu Wang, and Liuquan Sun. Variable screening for varying coefficient models with ultrahighdimensional survival data. Computational Statistics & Data Analysis, 172(??):??, August 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000780.

Anonymous:2022:S

[501] Anonymous. September 2022. Computational Statistics & Data Analysis, 173 (??):??, September 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2022:EBi

[502] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 173(??):??, September 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001141.

Pan:2022:FSF

[503] Yingli Pan. Feature screening and FDR control with knockoff features for ultrahigh-dimensional rightcensored data. Computational Statistics & Data Analysis, 173(??):??, September 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000846.

Blier-Wong:2022:SRF

[504] Christopher Blier-Wong, Hélène Cossette, and Etienne Marceau. Stochastic representation of FGM copulas using multivariate Bernoulli random variables. Computational Statistics & Data Analysis, 173(??):??, September 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732200086X.

Fan:2022:SSS

[505] Yiwei Fan and Junlong Zhao. Safe sample screening rules for multicategory angle-based support vector machines. Computational Statistics & Data Analysis, 173(??):??, September 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000883.

Liu:2022:VCH

[506] Hefei Liu, Xinyuan Song, and Baoxue Zhang. Varying-coefficient hidden Markov models with zero-effect regions. Computational Statistics & Data Analysis, 173(??):??, September 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947322000627.

Nanshan:2022:DMN

[507] Muye Nanshan, Nan Zhang, Xiaolei Xun, and Jiguo Cao. Dynamical modeling for non-Gaussian data with high-dimensional sparse ordinary differential equations. Computational Statistics & Data Analysis, 173(??): ??, September 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000639.

Merlo:2022:MQR

[508] Luca Merlo, Lea Petrella, Nicola Salvati, and Nikos Tzavidis. Marginal M-quantile regression for multivariate dependent data. Computational Statistics & Data Analysis, 173(??): ??, September 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000809.

Yuan:2022:MMA

[509] Chaoxia Yuan, Fang Fang, and Lyu Ni. Mallows model averaging with effective model size in fragmentary data prediction. Computational Statistics & Data Analysis, 173(??): ??, September 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000779.

Ouyang:2022:RBH

[510] Yanyan Ouyang, Jiamin Liu, Tiejun Tong, and Wangli Xu. A rankbased high-dimensional test for equality of mean vectors. Computational Statistics & Data Analysis, 173(??): ??, September 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000755.

Cadirci:2022:EBT

[511] Mehmet Siddik Cadirci, Dafydd Evans, Nikolai Leonenko, and Vitalii Makogin. Entropy-based test for generalised Gaussian distributions. Computational Statistics & Data Analysis, 173(??):??, September 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000822.

Wang:2022:LGT

[512] Jason Wang and Robert E. Weiss. Local and global topics in text modeling of web pages nested in web sites. Computational Statistics & Data Analysis, 173(??):??, September 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000986.

Sardy:2022:TTB

[513] Sylvain Sardy, Jairo Diaz-Rodriguez, and Caroline Giacobino. Thresholding tests based on affine LASSO to achieve non-asymptotic nominal level and high power under sparse and dense alternatives in high dimension. Computational Statistics & Data Analysis, 173(??):??, September 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000871.

Arcagni:2022:CRA

[514] Alberto Arcagni, Alessandro Avellone, and Marco Fattore. Complexity reduction and approximation of multidomain systems of partially ordered data. Computational Statistics & Data Analysis, 173(??):??, September 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001001.

Hamura:2022:LRV

[515] Yasuyuki Hamura, Kaoru Irie, and Shonosuke Sugasawa. Log-regularly varying scale mixture of normals for robust regression. Computational Statistics & Data Analysis, 173(??): ??, September 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000974.

Paige:2022:BMM

[516] John Paige, Geir-Arne Fuglstad, Andrea Riebler, and Jon Wakefield. Bayesian multiresolution modeling of georeferenced data: an extension of ‘LatticeKrig’. Computational Statistics & Data Analysis, 173(??):??, September 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000834.

Usefi:2022:CMS

[517] Hamid Usefi. Clustering, multicollinearity, and singular vectors. Computational Statistics & Data Analysis, 173 (??):??, September 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001037.

Su:2022:TSO

[518] Miaomiao Su, Ruoyu Wang, and Qihua Wang. A two-stage optimal subsampling estimation for missing data problems with large-scale data. Computational Statistics & Data Analysis, 173(??):??, September 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000858.

Wei:2022:SVC

[519] Honglei Wei, Hongfan Zhang, Hui Jiang, and Lei Huang. On the semi-varying coefficient dynamic panel data model with autocorrelated errors. Computational Statistics & Data Analysis, 173 (??):??, September 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732200038X.

Krivobokova:2022:JNP

[520] Tatyana Krivobokova, Paulo Serra, Francisco Rosales, and Karolina Klockmann. Joint non-parametric estimation of mean and auto-covariances for Gaussian processes. Computational Statistics & Data Analysis, 173 (??):??, September 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000998.

Anonymous:2022:O

[521] Anonymous. October 2022. Computational Statistics & Data Analysis, 174 (??):??, October 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2022:EBj

[522] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001220.

Wang:2022:GAE

[523] Shikun Wang, Zhao Li, Lan Lan, Jieyi Zhao, W. Jim Zheng, and Liang Li. GPU accelerated estimation of a shared random effect joint model for dynamic prediction. Computational Statistics & Data Analysis, 174 (??):??, October 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001086.

Li:2022:DDL

[524] Peili Li, Yuling Jiao, Xiliang Lu, and Lican Kang. A data-driven line search rule for support recovery in highdimensional data analysis. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001049.

Cole:2022:LSL

[525] D. Austin Cole, Robert B. Gramacy, and Mike Ludkovski. Large-scale local surrogate modeling of stochastic simulation experiments. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001177.

Chau:2022:TVS

[526] Joris Chau and Rainer von Sachs. Timevarying spectral matrix estimation via intrinsic wavelet regression for surfaces of Hermitian positive definite matrices. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000573.

Uemoto:2022:SVR

[527] Takumi Uemoto and Kanta Naito. Support vector regression with penalized likelihood. Computational Statistics & Data Analysis, 174(??): ??, October 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001025.

Aushev:2022:LFI

[528] Alexander Aushev, Henri Pesonen, Markus Heinonen, Jukka Corander, and Samuel Kaski. Likelihood-free inference with deep Gaussian processes. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001098.

Obst:2022:ILR

[529] David Obst, Badih Ghattas, Sandra Claudel, Jairo Cugliari, Yannig Goude, and Georges Oppenheim. Improved linear regression prediction by transfer learning. Computational Statistics & Data Analysis, 174(??): ??, October 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000792.

Sugasawa:2022:RFM

[530] Shonosuke Sugasawa and Genya Kobayashi.

Robust fitting of mixture models using weighted complete estimating equations. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001062.

Yuan:2022:IIS

[531] Qingcong Yuan, Xianyan Chen, Chenlu Ke, and Xiangrong Yin. Independence index sufficient variable screening for categorical responses. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001104.

Arnone:2022:RPA

[532] Eleonora Arnone, Federico Ferraccioli, Clara Pigolotti, and Laura M. Sangalli. A roughness penalty approach to estimate densities over twodimensional manifolds. Computational Statistics & Data Analysis, 174(??): ??, October 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001074.

Li:2022:EEP

[533] Mengyan Li, Yanyuan Ma, and Jiwei Zhao. Efficient estimation in a partially specified nonignorable propensity score model. Computational Statistics & Data Analysis, 174(??): ??, October 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321001560.

Huang:2022:ART

[534] Whitney K. Huang, Yu-Min Chung, Yu-Bo Wang, Jeff E. Mandel, and Hau-Tieng Wu. Airflow recovery from thoracic and abdominal movements using synchrosqueezing transform and locally stationary Gaussian process regression. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002188.

Williamson:2022:GBD

[535] S. Faye Williamson, Peter Jacko, and Thomas Jaki. Generalisations of a Bayesian decision-theoretic randomisation procedure and the impact of delayed responses. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947321002413.

Granados-Garcia:2022:BWA

[536] Guilllermo Granados-Garcia, Mark Fiecas, Shahbaba Babak, Norbert J. Fortin, and Hernando Ombao. Brain waves analysis via a non-parametric Bayesian mixture of autoregressive kernels. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947321002437.

Codazzi:2022:GGM

[537] Laura Codazzi, Alessandro Colombi, Matteo Gianella, Raffaele Argiento, Lucia Paci, and Alessia Pini. Gaussian graphical modeling for spectrometric data analysis. Computational Statistics & Data Analysis, 174(??): ??, October 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321002504.

To:2022:EVU

[538] Duc-Khanh To, Gianfranco Adimari, and Monica Chiogna. Estimation of the volume under a ROC surface in presence of covariates. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000147.

Banerjee:2022:HSM

[539] Sayantan Banerjee. Horseshoe shrinkage methods for Bayesian fusion estimation. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000305.

Embleton:2022:WTR

[540] Jonathan Embleton, Marina I. Knight, and Hernando Ombao. Wavelet testing for a replicate-effect within an ordered multiple-trial experiment. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000366.

Cho:2022:NPD

[541] Youngjoo Cho, Xiang Zhan, and Debashis Ghosh. Nonlinear predictive directions in clinical trials. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000561.

Amovin-Assagba:2022:ODM

[542] Martial Amovin-Assagba, Irène Gannaz, and Julien Jacques. Outlier detection in multivariate functional data through a contaminated mixture model. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000767.

Cai:2022:HDC

[543] Xizhen Cai, Yeying Zhu, Yuan Huang, and Debashis Ghosh. High-dimensional causal mediation analysis based on partial linear structural equation models. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322000810.

Ghosh:2022:HTP

[544] Santu Ghosh, Wenge Guo, and Samiran Ghosh. A hierarchical testing procedure for three arm non-inferiority trials. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001013.

Degras:2022:MSS

[545] David Degras, Chee-Ming Ting, and Hernando Ombao. Markov-switching state-space models with applications to neuroimaging. Computational Statistics & Data Analysis, 174(??): ??, October 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001050.

Ghaderinezhad:2022:WIM

[546] Fatemeh Ghaderinezhad, Christophe Ley, and Ben Serrien. The Wasserstein impact measure (WIM): a practical tool for quantifying prior impact in Bayesian statistics. Computational Statistics & Data Analysis, 174(??):??, October 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947321001869.

Seri:2022:CAD

[547] Raffaello Seri. Computing the asymptotic distribution of second-order Uand V -statistics. Computational Statistics & Data Analysis, 174(??): ??, October 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322000172.

Anonymous:2022:N

[548] Anonymous. November 2022. Computational Statistics & Data Analysis, 175 (??):??, November 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2022:EBk

[549] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 175(??):??, November 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001530.

Gao:2022:ANR

[550] Min Gao, Wenzhi Yang, Shipeng Wu, and Wei Yu. Asymptotic normality of residual density estimator in stationary and explosive autoregressive models. Computational Statistics & Data Analysis, 175(??): ??, November 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001293.

Palzer:2022:SSJ

[551] Elise F. Palzer, Christine H. Wendt, Russell P. Bowler, Craig P. Hersh, Sandra E. Safo, and Eric F. Lock. sJIVE: Supervised joint and individual variation explained. Computational Statistics & Data Analysis, 175 (??):??, November 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732200127X.

Pan:2022:PTE

[552] Shenyi Pan and Harry Joe. Predicting times to event based on vine copula models. Computational Statistics & Data Analysis, 175(??): ??, November 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001268.

Cho:2022:BCI

[553] Haeran Cho and Claudia Kirch. Bootstrap confidence intervals for multiple change points based on moving sum procedures. Computational Statistics & Data Analysis, 175(??): ??, November 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001323.

Petti:2022:CLB

[554] Danilo Petti, Alessia Eletti, Giampiero Marra, and Rosalba Radice. Copula link-based additive models for bivariate time-to-event outcomes with general censoring scheme. Computational Statistics & Data Analysis, 175(??):??, November 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732200130X.

Xu:2022:UFD

[555] Zishen Xu, Chenran Wang, and Wei Wu. A unified framework on defining depth for point process using function smoothing. Computational Statistics & Data Analysis, 175(??): ??, November 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001256.

Chen:2022:GMC

[556] Feifei Chen, M. Dolores Jiménez-Gamero, Simos Meintanis, and Lixing Zhu. A general Monte Carlo method for multivariate goodness-of-fit testing applied to elliptical families. Computational Statistics & Data Analysis, 175(??):??, November 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001281.

Zhang:2022:ELI

[557] Yuexia Zhang, Guoyou Qin, Zhongyi Zhu, and Jiajia Zhang. Empirical likelihood inference for longitudinal data with covariate measurement errors: an application to the LEAN study. Computational Statistics & Data Analysis, 175(??):??, November 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001335.

Feng:2022:SIH

[558] Sanying Feng, Kaidi Kong, Yinfei Kong, Gaorong Li, and Zhaoliang Wang. Statistical inference of heterogeneous treatment effect based on single-index model. Computational Statistics & Data Analysis, 175(??): ??, November 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001347.

Modarres:2022:HDD

[559] Reza Modarres. A high dimensional dissimilarity measure. Computational Statistics & Data Analysis, 175(??): ??, November 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001402.

Anonymous:2022:D

[560] Anonymous. December 2022. Computational Statistics & Data Analysis, 176 (??):??, December 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2022:EBl

[561] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 176(??):??, December 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001712.

Cruz:2022:IIS

[562] Ivette Raices Cruz, Johan Lindström, Matthias C. M. Troffaes, and Ullrika Sahlin. Iterative importance sampling with Markov chain Monte Carlo sampling in robust Bayesian analysis. Computational Statistics & Data Analysis, 176(??):??, December 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001384.

Park:2022:CDS

[563] Yujin Park, Kyongwon Kim, and Jae Keun Yoo. On cross-distance selection algorithm for hybrid sufficient dimension reduction. Computational Statistics & Data Analysis, 176(??):??, December 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001426.

Chu:2022:VCS

[564] Amanda M. Y. Chu, Chun Yin Ip, Benson S. Y. Lam, and Mike K. P. So. Vine copula statistical disclosure control for mixed-type data. Computational Statistics & Data Analysis, 176(??):??, December 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001414.

Mavrogonatou:2022:CAS

[565] Lida Mavrogonatou, Yuxuan Sun, David S. Robertson, and Sofía S. Villar. A comparison of allocation strategies for optimising clinical trial designs under variance heterogeneity. Computational Statistics & Data Analysis, 176(??):??, December 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001396.

Centofanti:2022:SLE

[566] Fabio Centofanti, Matteo Fontana, Antonio Lepore, and Simone Vantini. Smooth LASSO estimator for the function-on-function linear regression model. Computational Statistics & Data Analysis, 176(??): ??, December 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001360.

Wang:2022:HDR

[567] Yibo Wang and Rohana J. Karunamuni. High-dimensional robust regression with L_{q}-loss functions. Computational Statistics & Data Analysis, 176(??):??, December 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001475.

Bissiri:2022:NBM

[568] Pier Giovanni Bissiri, Galatia Cleanthous, Xavier Emery, Bernardo Nipoti, and Emilio Porcu. Nonparametric Bayesian modelling of longitudinally integrated covariance functions on spheres. Computational Statistics & Data Analysis, 176(??): ??, December 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001359.

Burghardt:2022:ADH

[569] Elliot Burghardt, Daniel Sewell, and Joseph Cavanaugh. Agglomerative and divisive hierarchical Bayesian clustering. Computational Statistics & Data Analysis, 176(??):??, December 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001463.

He:2022:EQA

[570] Fengyang He, Huixia Judy Wang, and Yuejin Zhou. Extremal quantile autoregression for heavy-tailed time series. Computational Statistics & Data Analysis, 176(??):??, December 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001438.

Goffinet:2022:FNP

[571] Etienne Goffinet, Mustapha Lebbah, Hanane Azzag, Giraldi Loïc, and Anthony Coutant. Functional nonparametric latent block model: a multivariate time series clustering approach for autonomous driving validation. Computational Statistics & Data Analysis, 176(??):??, December 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001451.

Dallakyan:2022:TSG

[572] Aramayis Dallakyan, Rakheon Kim, and Mohsen Pourahmadi. Time series graphical lasso and sparse VAR estimation. Computational Statistics & Data Analysis, 176(??):??, December 2022. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001372.

Miron:2022:RPL

[573] Julien Miron, Benjamin Poilane, and Eva Cantoni. Robust polytomous logistic regression. Computational Statistics & Data Analysis, 176(??): ??, December 2022. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732200144X.

Anonymous:2023:Ja

[574] Anonymous. January 2023. Computational Statistics & Data Analysis, 177 (??):??, January 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2023:EBa

[575] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 177(??):??, January 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001852.

Bianconcini:2023:DWQ

[576] Silvia Bianconcini and Silvia Cagnone. The dimension-wise quadrature estimation of dynamic latent variable models for count data. Computational Statistics & Data Analysis, 177(??): ??, January 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001657.

Hector:2023:PSA

[577] Emily C. Hector, Lan Luo, and Peter X.-K. Song. Parallel-and-stream accelerator for computationally fast supervised learning. Computational Statistics & Data Analysis, 177(??): ??, January 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001670.

Uddin:2023:SBV

[578] Md Nazir Uddin and Jeremy T. Gaskins. Shared Bayesian variable shrinkage in multinomial logistic regression. Computational Statistics & Data Analysis, 177(??):??, January 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001487.

Lee:2023:MIM

[579] Sangyeol Lee, Dongwon Kim, and Byungsoo Kim. Modeling and inference for multivariate time series of counts based on the INGARCH scheme. Computational Statistics & Data Analysis, 177(??):??, January 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001591.

Austin:2023:ONP

[580] Edward Austin, Gaetano Romano, Idris A. Eckley, and Paul Fearnhead. Online non-parametric changepoint detection with application to monitoring operational performance of network devices. Computational Statistics & Data Analysis, 177(??): ??, January 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001311.

McElroy:2023:IDO

[581] Tucker S. McElroy and Agnieszka Jach.

Identification of the differencing operator of a non-stationary time series via testing for zeroes in the spectral density. Computational Statistics & Data Analysis, 177(??):??, January 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001608.

Abramowicz:2023:NBC

[582] Konrad Abramowicz, Sara Sjöstedt de Luna, and Johan Strandberg. Nonparametric bagging clustering methods to identify latent structures from a sequence of dependent categorical data. Computational Statistics & Data Analysis, 177(??):??, January 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001633.

Zhong:2023:SSC

[583] Yan Zhong, Huiyan Sang, Scott J. Cook, and Paul M. Kellstedt. Sparse spatially clustered coefficient model via adaptive regularization. Computational Statistics & Data Analysis, 177(??):??, January 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732200161X.

Tang:2023:EPF

[584] Qingguo Tang, Wei Tu, and Linglong Kong. Estimation for partial functional partially linear additive model. Computational Statistics & Data Analysis, 177(??):??, January 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001645.

Gonzalez:2023:AEC

[585] Juan Claramunt González, Arnout van Delden, and Ton de Waal. Assessment of the effect of constraints in a new multivariate mixed method for statistical matching. Computational Statistics & Data Analysis, 177 (??):??, January 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001499.

Robin:2023:MBE

[586] Stéphane Robin and Luca Scrucca.

Mixture-based estimation of entropy. Computational Statistics & Data Analysis, 177(??):??, January 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001621.

Caubet:2023:BJM

[587] Miguel Caubet, Mariia Samoilenko, Simon Drouin, Daniel Sinnett, Maja Krajinovic, Caroline Laverdière, Valérie Marcil, and Geneviève Lefebvre. Bayesian joint modeling for causal mediation analysis with a binary outcome and a binary mediator: Exploring the role of obesity in the association between cranial radiation therapy for childhood acute lymphoblastic leukemia treatment and the long-term risk of insulin resistance. Computational Statistics & Data Analysis, 177(??): ??, January 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001669.

Anonymous:2023:F

[588] Anonymous. February 2023. Computational Statistics & Data Analysis, 178 (??):??, February 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2023:EBb

[589] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 178(??):??, February 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002183.

Merrell:2023:MDP

[590] David Merrell, Thevaa Chandereng, and Yeonhee Park. A Markov decision process for response-adaptive randomization in clinical trials. Computational Statistics & Data Analysis, 178(??):??, February 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001797.

Qin:2023:VAM

[591] Yichen Qin, Linna Wang, Yang Li, and Rong Li. Visualization and assessment of model selection uncertainty. Computational Statistics & Data Analysis, 178(??):??, February 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001785.

Hirose:2023:HCM

[592] Kei Hirose, Kanta Miura, and Atori Koie. Hierarchical clustered multiclass discriminant analysis via crossvalidation. Computational Statistics & Data Analysis, 178(??):??, February 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947322001931.

Hu:2023:FBI

[593] Zhixiong Hu and Raquel Prado. Fast Bayesian inference on spectral analysis of multivariate stationary time series. Computational Statistics & Data Analysis, 178(??):??, February 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001761.

Acharki:2023:RPI

[594] Naoufal Acharki, Antoine Bertoncello, and Josselin Garnier. Robust prediction interval estimation for Gaussian processes by cross-validation method. Computational Statistics & Data Analysis, 178(??):??, February 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001773.

Huang:2023:GSA

[595] Danyang Huang, Wei Hu, Bingyi Jing, and Bo Zhang. Grouped spatial autoregressive model. Computational Statistics & Data Analysis, 178(??):??, February 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001815.

Hormann:2023:PFR

[596] Siegfried Hörmann and Fatima Jammoul. Prediction in functional regression with discretely observed and noisy covariates. Computational Statistics & Data Analysis, 178(??): ??, February 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001803.

Ghosal:2023:SCE

[597] Rahul Ghosal, Sujit Ghosh, Jacek Urbanek, Jennifer A. Schrack, and Vadim Zipunnikov. Shape-constrained estimation in functional regression with Bernstein polynomials. Computational Statistics & Data Analysis, 178(??): ??, February 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001943.

Boente:2023:RSA

[598] Graciela Boente and Alejandra Mercedes Martínez. A robust spline approach in partially linear additive models. Computational Statistics & Data Analysis, 178(??):??, February 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947322001918.

Ke:2023:STQ

[599] Baofang Ke, Weihua Zhao, and Lei Wang. Smoothed tensor quantile regression estimation for longitudinal data. Computational Statistics & Data Analysis, 178(??):??, February 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732200189X.

Yu:2023:ODS

[600] Jun Yu, Xiran Meng, and Yaping Wang. Optimal designs for semiparametric dose-response models under random contamination. Computational Statistics & Data Analysis, 178(??):??, February 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001955.

Yang:2023:MSL

[601] Yuehan Yang, Siwei Xia, and Hu Yang.

Multivariate sparse Laplacian shrinkage for joint estimation of two graphical structures. Computational Statistics & Data Analysis, 178(??): ??, February 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002006.

Chakraborty:2023:NMB

[602] Nilanjan Chakraborty and Lyudmila Sakhanenko. Novel multiplier bootstrap tests for high-dimensional data with applications to MANOVA. Computational Statistics & Data Analysis, 178(??):??, February 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001992.

Andrade:2023:CQR

[603] Ana C. C. Andrade, Gustavo H. A.

Pereira, and Rinaldo Artes. The circular quantile residual. Computational Statistics & Data Analysis, 178(??):??, February 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732200192X.

Zhang:2023:ALS

[604] Yangchun Zhang, Yirui Zhou, and Xiaowei Liu. Applications on linear spectral statistics of high-dimensional sample covariance matrix with divergent spectrum. Computational Statistics & Data Analysis, 178(??): ??, February 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001979.

Anonymous:2023:Ma

[605] Anonymous. March 2023. Computational Statistics & Data Analysis, 179 (??):??, March 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2023:EBc

[606] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002523.

Turnbull:2023:SET

[607] Kathryn Turnbull, Christopher Nemeth, Matthew Nunes, and Tyler McCormick. Sequential estimation of temporally evolving latent space network models. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002079.

Jeong:2023:OLD

[608] Kuhwan Jeong, Minwoo Chae, and Yongdai Kim. Online learning for the Dirichlet process mixture model via weakly conjugate approximation. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002067.

Anyosa:2023:ASD

[609] Susan Anyosa, Jo Eidsvik, and Oscar Pizarro. Adaptive spatial designs minimizing the integrated Bernoulli variance in spatial logistic regression models — with an application to benthic habitat mapping. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002237.

Kim:2023:RTS

[610] Seungkyu Kim, Seongoh Park, Johan Lim, and Sang Han Lee. Robust tests for scatter separability beyond Gaussianity. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002134.

Chen:2023:UFM

[611] Sixia Chen and David Haziza. A unified framework of multiply robust estimation approaches for handling incomplete data. Computational Statistics & Data Analysis, 179 (??):??, March 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002262.

Hatjispyros:2023:MMD

[612] Spyridon J. Hatjispyros, Christos Merkatas, and Stephen G. Walker. Mixture models with decreasing weights. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002316.

Fernandez-de-Marcos:2023:DDS

[613] Alberto Fernández de Marcos and Eduardo García-Portugués. Data-driven stabilizations of goodness-of-fit tests. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732200233X.

Li:2023:LSS

[614] Yanxin Li and Stephen G. Walker. A latent slice sampling algorithm. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002328.

Pereira:2023:BNH

[615] Luz Adriana Pereira, Luis Gutiérrez, Daniel Taylor-Rodríguez, and Ramsés H. Mena. Bayesian nonparametric hypothesis testing for longitudinal data analysis. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947322002092.

Wang:2023:GSL

[616] Xiaodi Wang and Hengzhen Huang.

Group symmetric Latin hypercube designs for symmetrical global sensitivity analysis. Computational Statistics & Data Analysis, 179(??): ??, March 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002055.

Bhatnagar:2023:SAM

[617] Sahir R. Bhatnagar, Tianyuan Lu, Amanda Lovato, David L. Olds, Michael S. Kobor, Michael J. Meaney, Kieran O’Donnell, Archer Y. Yang, and Celia M. T. Greenwood. A sparse additive model for high-dimensional interactions with an exposure variable. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002043.

Zhao:2023:MMA

[618] Yi Zhao and Xi Luo. Multilevel mediation analysis with structured unmeasured mediator-outcome confounding. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002031.

Welz:2023:CRE

[619] Thilo Welz, Wolfgang Viechtbauer, and Markus Pauly. Cluster-robust estimators for multivariate mixedeffects meta-regression. Computational Statistics & Data Analysis, 179 (??):??, March 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002110.

Zhao:2023:CMC

[620] Xin Zhao, Jingru Zhang, and Wei Lin. Clustering multivariate count data via Dirichlet-multinomial network fusion. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002146.

Yang:2023:BDP

[621] Yihe Yang, Hongsheng Dai, and Jianxin Pan. Block-diagonal precision matrix regularization for ultrahigh dimensional data. Computational Statistics & Data Analysis, 179 (??):??, March 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002109.

An:2023:VSB

[622] Hyowon An, Kai Zhang, Hannu Oja, and J. S. Marron. Variable screening based on Gaussian centered L-moments.

Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002122.

McGonigle:2023:RME

[623] Euan T. McGonigle and Haeran Cho. Robust multiscale estimation of time-average variance for time series segmentation. Computational Statistics & Data Analysis, 179(??): ??, March 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002286.

Li:2023:MRM

[624] Xun-Jian Li, Yuan Sun, Guo-Liang Tian, Jiajuan Liang, and Jianhua Shi. Mean regression model for the zero-truncated Poisson distribution and its generalization. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002304.

Nikoloulopoulos:2023:EFI

[625] Aristidis K. Nikoloulopoulos. Efficient and feasible inference for high-dimensional normal copula regression models. Computational Statistics & Data Analysis, 179(??): ??, March 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002341.

Hao:2023:NIS

[626] Meiling Hao, Yuanyuan Lin, Guohao Shen, and Wen Su. Nonparametric inference on smoothed quantile regression process. Computational Statistics & Data Analysis, 179 (??):??, March 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002250.

Kim:2023:SVR

[627] Daeju Kim, Shuichi Kawano, and Yoshiyuki Ninomiya. Smoothly varying regularization. Computational Statistics & Data Analysis, 179(??): ??, March 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002249.

Cipriani:2023:TBG

[628] Alessandra Cipriani, Christian Hirsch, and Martina Vittorietti. Topologybased goodness-of-fit tests for sliced spatial data. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947322002353.

Ma:2023:RFP

[629] Xuan Ma, Jianhua Zhao, Yue Wang, Changchun Shang, and Fen Jiang. Robust factored principal component analysis for matrix-valued outlier accommodation and detection. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002377.

Weishampel:2023:CSM

[630] Anthony Weishampel, Ana-Maria Staicu, and William Rand. Classification of social media users with generalized functional data analysis. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002274.

Vo:2023:EFS

[631] Thanh Huan Vo, Guillaume Chauvet, André Happe, Emmanuel Oger, Stéphane Paquelet, and Valérie Garès. Extending the Fellegi–Sunter record linkage model for mixed-type data with application to the French national health data system. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002365.

Yang:2023:BCM

[632] Haoyu Yang, Yichen Qin, Fan Wang, Yang Li, and Feifang Hu. Balancing covariates in multi-arm trials via adaptive randomization. Computational Statistics & Data Analysis, 179 (??):??, March 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002225.

Hannaford:2023:SBH

[633] Naomi E. Hannaford, Sarah E. Heaps, Tom M. W. Nye, Thomas P. Curtis, Ben Allen, Andrew Golightly, and Darren J. Wilkinson. A sparse Bayesian hierarchical vector autoregressive model for microbial dynamics in a wastewater treatment plant. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002390.

Schmid:2023:TBE

[634] Lena Schmid, Alexander Gerharz, Andreas Groll, and Markus Pauly. Treebased ensembles for multi-output regression: Comparing multivariate approaches with separate univariate ones. Computational Statistics & Data Analysis, 179(??):??, March 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002080.

Anonymous:2023:Aa

[635] Anonymous. April 2023. Computational Statistics & Data Analysis, 180 (??):??, April 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2023:EBd

[636] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 180(??):??, April 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323000075.

Gangloff:2023:DPP

[637] Hugo Gangloff, Katherine Morales, and Yohan Petetin. Deep parameterizations of pairwise and triplet Markov models for unsupervised classification of sequential data. Computational Statistics & Data Analysis, 180(??):??, April 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002432.

Lamprinakou:2023:BBI

[638] Stamatina Lamprinakou, Mauricio Barahona, Seth Flaxman, Sarah Filippi, Axel Gandy, and Emma J. McCoy. BART-based inference for Poisson processes. Computational Statistics & Data Analysis, 180(??): ??, April 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002389.

Mendizabal:2023:FFA

[639] Valentina Zelaya Mendizábal, Marc Boullé, and Fabrice Rossi. Fast and fully-automated histograms for large-scale data sets. Computational Statistics & Data Analysis, 180 (??):??, April 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002481.

Zhuang:2023:PTM

[640] Haoxin Zhuang, Liqun Diao, and Grace Y. Yi. Polya tree Monte Carlo method. Computational Statistics & Data Analysis, 180(??):??, April 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947322002456.

Eendebak:2023:SET

[641] Pieter T. Eendebak, Eric D. Schoen, Alan R. Vazquez, and Peter Goos. Systematic enumeration of two-level evenodd designs of strength 3. Computational Statistics & Data Analysis, 180(??):??, April 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002584.

Cao:2023:AUC

[642] Yongxiu Cao and Jichang Yu. Adjusting for unmeasured confounding in survival causal effect using validation data. Computational Statistics & Data Analysis, 180(??):??, April 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002407.

Yang:2023:JIA

[643] Xi Yang, Katherine A. Hoadley, Jan Hannig, and J. S. Marron. Jackstraw inference for AJIVE data integration. Computational Statistics & Data Analysis, 180(??):??, April 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002298.

Chen:2023:CDB

[644] Kunzhi Chen, Weining Shen, and Weixuan Zhu. Covariate dependent Beta-GOS process. Computational Statistics & Data Analysis, 180 (??):??, April 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002420.

Naderi:2023:RMR

[645] Mehrdad Naderi, Elham Mirfarah, Wan-Lun Wang, and Tsung-I Lin. Robust mixture regression modeling based on the normal mean-variance mixture distributions. Computational Statistics & Data Analysis, 180 (??):??, April 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002419.

Soale:2023:PER

[646] Abdul-Nasah Soale. Projection expectile regression for sufficient dimension reduction. Computational Statistics & Data Analysis, 180(??): ??, April 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002468.

Wang:2023:SHA

[647] Xin Wang, Zhengyuan Zhu, and Hao Helen Zhang. Spatial heterogeneity automatic detection and estimation. Computational Statistics & Data Analysis, 180(??):??, April 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732200247X.

Wang:2023:CML

[648] Jiangzhou Wang, Tingting Cui, Wensheng Zhu, and Pengfei Wang. Covariatemodulated large-scale multiple testing under dependence. Computational Statistics & Data Analysis, 180 (??):??, April 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002444.

Yamaguchi:2023:MAT

[649] Hikaru Yamaguchi and Hidetoshi Murakami. The multi-aspect tests in the presence of ties. Computational Statistics & Data Analysis, 180 (??):??, April 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002602.

DAngelo:2023:LWM

[650] Nicoletta D’Angelo, Giada Adelfio, and Jorge Mateu. Locally weighted minimum contrast estimation for spatiotemporal log-Gaussian Cox processes. Computational Statistics & Data Analysis, 180(??):??, April 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002596.

daSilva:2023:DRT

[651] Murilo da Silva, T. N. Sriram, and Yuan Ke. Dimension reduction in time series under the presence of conditional heteroscedasticity. Computational Statistics & Data Analysis, 180 (??):??, April 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002626.

Xiong:2023:UMF

[652] Wei Xiong, Yaxian Chen, and Shuangge Ma. Unified model-free interaction screening via CV-entropy filter. Computational Statistics & Data Analysis, 180(??):??, April 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732200264X.

Li:2023:EGI

[653] Xiao Li, Yuqiang Li, and Xianyi Wu. Empirical Gittins index strategies with ε-explorations for multiarmed bandit problems. Computational Statistics & Data Analysis, 180 (??):??, April 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001906.

Liu:2023:STO

[654] Ran Liu and Lixing Zhu. Specification testing for ordinary differential equation models with fixed design and applications to COVID-19 epidemic models. Computational Statistics & Data Analysis, 180(??):??, April 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322001967.

Li:2023:GMD

[655] Lu Li, Chenlu Ke, Xiangrong Yin, and Zhou Yu. Generalized martingale difference divergence: Detecting conditional mean independence with applications in variable screening. Computational Statistics & Data Analysis, 180(??):??, April 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322001980.

Goedhart:2023:EPP

[656] Jeroen M. Goedhart, Thomas Klausch, and Mark A. van de Wiel. Estimation of predictive performance in high-dimensional data settings using learning curves. Computational Statistics & Data Analysis, 180(??): ??, April 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732200202X.

Anonymous:2023:M

[657] Anonymous. May 2023. Computational Statistics & Data Analysis, 181 (??):??, May 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2023:EBe

[658] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 181(??):??, May 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323000312.

Niekerk:2023:NAB

[659] Janet Van Niekerk, Elias Krainski, Denis Rustand, and Håvard Rue. A new avenue for Bayesian inference with INLA.

Computational Statistics & Data Analysis, 181(??):??, May 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000038.

Greco:2023:ITA

[660] Luca Greco, Simona Pacillo, and Piera Maresca. An impartial trimming algorithm for robust circle fitting. Computational Statistics & Data Analysis, 181(??):??, May 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002663.

Zhang:2023:GCJ

[661] Zili Zhang, Christiana Charalambous, and Peter Foster. A Gaussian copula joint model for longitudinal and time-to-event data with random effects. Computational Statistics & Data Analysis, 181(??):??, May 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002651.

Hapfelmeier:2023:EPT

[662] Alexander Hapfelmeier, Roman Hornung, and Bernhard Haller. Efficient permutation testing of variable importance measures by the example of random forests. Computational Statistics & Data Analysis, 181 (??):??, May 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002699.

Lopez-Diaz:2023:OBC

[663] María Concepción López-Díaz, Miguel López-Díaz, and Sergio Martínez-Fernández. On the optimal binary classifier with an application. Computational Statistics & Data Analysis, 181(??):??, May 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002638.

Choi:2023:DCS

[664] Semin Choi, Yesool Kim, and Gunwoong Park. Densely connected sub-Gaussian linear structural equation model learning via ℓ_{1}- and ℓ_{2}- regularized regressions. Computational Statistics & Data Analysis, 181 (??):??, May 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323000026.

Pini:2023:LIF

[665] Alessia Pini, Helle Sørensen, Anders Tolver, and Simone Vantini. Local inference for functional linear mixed models. Computational Statistics & Data Analysis, 181(??):??, May 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002687.

Sui:2023:BCL

[666] Yuelei Sui, Scott H. Holan, and Wen-Hsi Yang. Bayesian circular lattice filters for computationally efficient estimation of multivariate timevarying autoregressive models. Computational Statistics & Data Analysis, 181(??):??, May 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000014.

Srinivasan:2023:IMF

[667] Arun Srinivasan, Lingzhou Xue, and Xiang Zhan. Identification of microbial features in multivariate regression under false discovery rate control. Computational Statistics & Data Analysis, 181(??):??, May 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002018.

Febrero-Bande:2023:FCB

[668] Manuel Febrero-Bande, Wenceslao González-Manteiga, Brenda Prallon, and Yuri F. Saporito. Functional classification of bitcoin addresses. Computational Statistics & Data Analysis, 181(??):??, May 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002675.

Fokianos:2023:ESI

[669] Konstantinos Fokianos, Claudia Kirch, and Hernando Ombao. Editorial for the special issue on Time Series Analysis. Computational Statistics & Data Analysis, 181(??):??, May 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947322002559.

Ma:2023:SIB

[670] Shuangge Ma, Martina Mittlboeck, F. Javier Rubio, and Catherine C. Liu. 2^{nd} special issue on BIOSTATISTICS. Computational Statistics & Data Analysis, 181(??):??, May 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947322002614.

Anonymous:2023:Jb

[671] Anonymous. June 2023. Computational Statistics & Data Analysis, 182 (??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2023:EBf

[672] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 182(??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323000452.

Andersson:2023:FEM

[673] Björn Andersson, Shaobo Jin, and Maoxin Zhang. Fast estimation of multiple group generalized linear latent variable models for categorical observed variables. Computational Statistics & Data Analysis, 182 (??):??, June 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732300021X.

Wang:2023:QTF

[674] Kai Y. K. Wang, Cathy W. S. Chen, and Mike K. P. So. Quantile threefactor model with heteroskedasticity, skewness, and leptokurtosis. Computational Statistics & Data Analysis, 182(??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000130.

Soave:2023:RRT

[675] David Soave and Jerald F. Lawless. Regularized regression for two phase failure time studies. Computational Statistics & Data Analysis, 182(??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000142.

Roy:2023:BSM

[676] Arkaprava Roy and Abhra Sarkar.

Bayesian semiparametric multivariate density deconvolution via stochastic rotation of replicates. Computational Statistics & Data Analysis, 182 (??):??, June 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000178.

Cui:2023:CPT

[677] Junfeng Cui, Guanghui Wang, Changliang Zou, and Zhaojun Wang. Changepoint testing for parallel data sets with FDR control. Computational Statistics & Data Analysis, 182(??): ??, June 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000166.

Song:2023:NEP

[678] Jun Song, Kyongwon Kim, and Jae Keun Yoo. On a nonlinear extension of the principal fitted component model. Computational Statistics & Data Analysis, 182(??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732300018X.

Khan:2023:TSN

[679] Ruhul Ali Khan. Two-sample nonparametric test for proportional reversed hazards. Computational Statistics & Data Analysis, 182(??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947323000191.

Cheng:2023:GEP

[680] Suli Cheng and Jianbao Chen. GMM estimation of partially linear additive spatial autoregressive model. Computational Statistics & Data Analysis, 182(??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323000233.

Zou:2023:CQR

[681] Yuye Zou and Chengxin Wu. Composite quantile regression analysis of survival data with missing cause-of-failure information and its application to breast cancer clinical trial. Computational Statistics & Data Analysis, 182(??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000221.

Borgonovo:2023:ECM

[682] Emanuele Borgonovo, Valentina Ghidini, Roman Hahn, and Elmar Plischke. Explaining classifiers with measures of statistical association. Computational Statistics & Data Analysis, 182(??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000129.

Zeng:2023:ODS

[683] Yicheng Zeng and Lixing Zhu. Order determination for spiked-type models with a divergent number of spikes. Computational Statistics & Data Analysis, 182(??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000154.

Chassan:2023:HTM

[684] Malika Chassan and Didier Concordet.

How to test the missing data mechanism in a hidden Markov model. Computational Statistics & Data Analysis, 182(??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000348.

Xu:2023:DES

[685] Danli Xu and Yong Wang. Density estimation for spherical data using nonparametric mixtures. Computational Statistics & Data Analysis, 182(??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000269.

Motegi:2023:SSM

[686] Ryosuke Motegi and Yoichi Seki. SML-SOM: the shrinking maximum likelihood self-organizing map. Computational Statistics & Data Analysis, 182(??):??, June 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000257.

Zhu:2023:SFF

[687] Hanbing Zhu, Yuanyuan Zhang, Yehua Li, and Heng Lian. Semiparametric function-on-function quantile regression model with dynamic singleindex interactions. Computational Statistics & Data Analysis, 182(??): ??, June 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000385.

Liang:2023:IQT

[688] Jinwen Liang, Wolfgang Karl Härdle, and Maozai Tian. Imputed quantile tensor regression for near-sited spatial-temporal data. Computational Statistics & Data Analysis, 182 (??):??, June 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000245.

Anonymous:2023:Jc

[689] Anonymous. July 2023. Computational Statistics & Data Analysis, 183 (??):??, July 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2023:EBg

[690] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 183(??):??, July 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000622.

Gromping:2023:UIS

[691] Ulrike Grömping. A unifying implementation of stratum (aka strong) orthogonal arrays. Computational Statistics & Data Analysis, 183 (??):??, July 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000506.

Escobar-Bach:2023:NEC

[692] Mikael Escobar-Bach and Ingrid Van Keilegom. Nonparametric estimation of conditional cure models for heavy-tailed distributions and under insufficient follow-up. Computational Statistics & Data Analysis, 183(??):??, July 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000397.

Kepplinger:2023:RVS

[693] David Kepplinger. Robust variable selection and estimation via adaptive elastic net S-estimators for linear regression. Computational Statistics & Data Analysis, 183(??):??, July 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000415.

DeIaco:2023:STG

[694] S. De Iaco. Spatio-temporal generalized complex covariance models based on convolution. Computational Statistics & Data Analysis, 183 (??):??, July 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000208.

Holter:2023:TPS

[695] Julia C. Holter and Jonathan W. Stallrich. Tuning parameter selection for penalized estimation via R^{2}. Computational Statistics & Data Analysis, 183(??):??, July 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000403.

Anonymous:2023:Ab

[696] Anonymous. August 2023. Computational Statistics & Data Analysis, 184 (??):??, August 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2023:EBh

[697] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 184(??):??, August 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000853.

Vexler:2023:HIS

[698] Albert Vexler, Xinyu Gao, and Jiaojiao Zhou. How to implement signedrank wilcox.test() type procedures when a center of symmetry is unknown. Computational Statistics & Data Analysis, 184(??):??, August 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000579.

Wang:2023:BTG

[699] Xuqin Wang and Muyi Li. Bootstrapping the transformed goodness-of-fit test on heavy-tailed GARCH models. Computational Statistics & Data Analysis, 184(??):??, August 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000555.

Zhao:2023:AWS

[700] Yujie Zhao and Xiaoming Huo. Accelerate the warm-up stage in the Lasso computation via a homotopic approach. Computational Statistics & Data Analysis, 184(??):??, August 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000580.

Dong:2023:WLS

[701] Qingkai Dong, Binxia Liu, and Hui Zhao. Weighted least squares model averaging for accelerated failure time models. Computational Statistics & Data Analysis, 184(??): ??, August 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000543.

Ranjbar:2023:EHC

[702] Setareh Ranjbar, Nicola Salvati, and Barbara Pacini. Estimating heterogeneous causal effects in observational studies using small area predictors. Computational Statistics & Data Analysis, 184(??):??, August 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000531.

Kawakubo:2023:SAE

[703] Yuki Kawakubo and Genya Kobayashi. Small area estimation of general finite-population parameters based on grouped data. Computational Statistics & Data Analysis, 184(??): ??, August 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732300052X.

Capanu:2023:SBV

[704] Marinela Capanu, Mihai Giurcanu, Colin B. Begg, and Mithat Gönen. Subsampling based variable selection for generalized linear models. Computational Statistics & Data Analysis, 184(??):??, August 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000518.

Park:2023:FPC

[705] Yeonjoo Park, Hyunsung Kim, and Yaeji Lim. Functional principal component analysis for partially observed elliptical process. Computational Statistics & Data Analysis, 184 (??):??, August 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000567.

Chiou:2023:ESI

[706] Jeng-Min Chiou, Frederic Ferraty, Jeff Goldsmith, Debashis Paul, and Jian Qing Shi. Editorial for the 2nd special issue on high-dimensional and functional data analysis. Computational Statistics & Data Analysis, 184(??):??, August 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000373.

Anonymous:2023:S

[707] Anonymous. September 2023. Computational Statistics & Data Analysis, 185 (??):??, September 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2023:EBi

[708] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 185(??):??, September 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001007.

Lowe:2023:AIS

[709] Tom E. Lowe, Andrew Golightly, and Chris Sherlock. Accelerating inference for stochastic kinetic models. Computational Statistics & Data Analysis, 185(??):??, September 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000713.

Bee:2023:UME

[710] Marco Bee. Unsupervised mixture estimation via approximate maximum likelihood based on the Cramér– von Mises distance. Computational Statistics & Data Analysis, 185(??): ??, September 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000750.

Zhang:2023:IBB

[711] Fan Zhang, Ray-Bing Chen, Ying Hung, and Xinwei Deng. Indicatorbased Bayesian variable selection for Gaussian process models in computer experiments. Computational Statistics & Data Analysis, 185(??): ??, September 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000683.

Moscovich:2023:FCP

[712] Amit Moscovich. Fast calculation of p-values for one-sided Kolmogorov– Smirnov type statistics. Computational Statistics & Data Analysis, 185 (??):??, September 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000804.

Li:2023:SMA

[713] Mengyu Li and Xiaoguang Wang. Semiparametric model averaging method for survival probability predictions of patients. Computational Statistics & Data Analysis, 185(??):??, September 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000701.

Miecznikowski:2023:ECF

[714] Jeffrey C. Miecznikowski and Jiefei Wang. Exceedance control of the false discovery proportion via high precision inversion method of Berk– Jones statistics. Computational Statistics & Data Analysis, 185(??):??, September 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000695.

Lenzi:2023:NNP

[715] Amanda Lenzi, Julie Bessac, Johann Rudi, and Michael L. Stein. Neural networks for parameter estimation in intractable models. Computational Statistics & Data Analysis, 185(??):??, September 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000737.

Li:2023:EMT

[716] Xiaoting Li and Harry Joe. Estimation of multivariate tail quantities. Computational Statistics & Data Analysis, 185(??):??, September 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323000725.

Meng:2023:DCE

[717] Rui Meng, Fan Yang, and Won Hwa Kim. Dynamic covariance estimation via predictive Wishart process with an application on brain connectivity estimation. Computational Statistics & Data Analysis, 185(??): ??, September 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323000749.

Zhao:2023:MAH

[718] Zhiwei Zhao, Chixiang Chen, Bhim Mani Adhikari, L. Elliot Hong, Peter Kochunov, and Shuo Chen. Mediation analysis for high-dimensional mediators and outcomes with an application to multimodal imaging data. Computational Statistics & Data Analysis, 185(??):??, September 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323000762.

Pratiwi:2023:PPP

[719] Bunga C. Pratiwi, Elise Dusseldorp, Julian D. Karch, and Mark de Rooij. Predictive performance of psychological tests: Is it better to use items than subscales? Computational Statistics & Data Analysis, 185(??): ??, September 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000786.

Sohn:2023:FCM

[720] Jinwon Sohn, Seonghyun Jeong, Young Min Cho, and Taeyoung Park. Functional clustering methods for binary longitudinal data with temporal heterogeneity. Computational Statistics & Data Analysis, 185(??): ??, September 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000774.

Shi:2023:MST

[721] Xiangyu Shi, Min Xu, and Jiang Du. Max-sum test based on Spearman’s footrule for high-dimensional independence tests. Computational Statistics & Data Analysis, 185(??): ??, September 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000798.

Sun:2023:ORS

[722] Xiaofei Sun, Hongwei Wang, Chao Cai, Mei Yao, and Kangning Wang. Online renewable smooth quantile regression. Computational Statistics & Data Analysis, 185(??):??, September 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323000920.

Zhao:2023:HTC

[723] Kaige Zhao, Tingting Zou, Shurong Zheng, and Jing Chen. Hypothesis testing on compound symmetric structure of high-dimensional covariance matrix. Computational Statistics & Data Analysis, 185(??): ??, September 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323000907.

Anonymous:2023:O

[724] Anonymous. October 2023. Computational Statistics & Data Analysis, 186 (??):??, October 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2023:EBj

[725] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 186(??):??, October 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001159.

Kim:2023:BPM

[726] Jonathan Kim, Brian J. Sandri, Raghavendra B. Rao, and Eric F. Lock. Bayesian predictive modeling of multi-source multi-way data. Computational Statistics & Data Analysis, 186(??):??, October 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323000944.

Liang:2023:LSQ

[727] Weijuan Liang, Qingzhao Zhang, and Shuangge Ma. Locally sparse quantile estimation for a partially functional interaction model. Computational Statistics & Data Analysis, 186(??):??, October 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000932.

Agboola:2023:NBC

[728] Oluwagbenga David Agboola and Han Yu. Neighborhood-based cross fitting approach to treatment effects with high-dimensional data. Computational Statistics & Data Analysis, 186(??):??, October 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000919.

Anonymous:2023:N

[729] Anonymous. November 2023. Computational Statistics & Data Analysis, 187 (??):??, November 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2023:EBk

[730] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001421.

Wu:2023:DCD

[731] Shihao Wu, Zhe Li, and Xuening Zhu.

A distributed community detection algorithm for large scale networks under stochastic block models. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001056.

Priddle:2023:TSP

[732] Jacob W. Priddle and Christopher Drovandi. Transformations in semiparametric Bayesian synthetic likelihood. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947323001081.

Moins:2023:REV

[733] Théo Moins, Julyan Arbel, Stéphane Girard, and Anne Dutfoy. Reparameterization of extreme value framework for improved Bayesian workflow. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001184.

Murray:2023:FEG

[734] James Murray and Pete Philipson. Fast estimation for generalised multivariate joint models using an approximate EM algorithm. Computational Statistics & Data Analysis, 187 (??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001305.

Ajroldi:2023:CPB

[735] Niccolò Ajroldi, Jacopo Diquigiovanni, Matteo Fontana, and Simone Vantini. Conformal prediction bands for twodimensional functional time series. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001329.

Ke:2023:PSV

[736] Chenlu Ke, Wei Yang, Qingcong Yuan, and Lu Li. Partial sufficient variable screening with categorical controls. Computational Statistics & Data Analysis, 187(??): ??, November 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000956.

Ouyang:2023:BAL

[737] Jiangrong Ouyang and Howard Bondell. Bayesian analysis of longitudinal data via empirical likelihood. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323000968.

Tan:2023:EPP

[738] Xin Tan, Haoran Zhan, and Xu Qin.

Estimation of projection pursuit regression via alternating linearization. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001044.

Perepolkin:2023:TQB

[739] Dmytro Perepolkin, Benjamin Goodrich, and Ullrika Sahlin. The tenets of quantile-based inference in Bayesian models. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947323001068.

Boente:2023:REF

[740] Graciela Boente and Daniela Parada. Robust estimation for functional quadratic regression models. Computational Statistics & Data Analysis, 187(??): ??, November 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001093.

Chen:2023:ADM

[741] Yewen Chen, Xiaohui Chang, Fangzhi Luo, and Hui Huang. Additive dynamic models for correcting numerical model outputs. Computational Statistics & Data Analysis, 187(??): ??, November 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732300110X.

Yang:2023:ORM

[742] Yaohong Yang, Weihua Zhao, and Lei Wang. Online regularized matrix regression with streaming data. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001202.

Schmidt:2023:MDS

[743] Rouven Schmidt and Thomas Kneib.

Multivariate distributional stochastic frontier models. Computational Statistics & Data Analysis, 187(??): ??, November 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732300107X.

Kei:2023:PSM

[744] Yik Lun Kei, Yanzhen Chen, and Oscar Hernan Madrid Padilla. A partially separable model for dynamic valued networks. Computational Statistics & Data Analysis, 187 (??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001226.

Yang:2023:CEE

[745] Yaohong Yang, Lei Wang, Jiamin Liu, Rui Li, and Heng Lian. Communicationefficient estimation of quantile matrix regression for massive datasets. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001238.

Zhou:2023:SDP

[746] Xinyu Zhou, Yijia Ma, and Wei Wu. Statistical depth for point process via the isometric log-ratio transformation. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732300124X.

Park:2023:MLS

[747] Beomjin Park and Changyi Park. Multiclass Laplacian support vector machine with functional analysis of variance decomposition. Computational Statistics & Data Analysis, 187(??): ??, November 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001251.

Martinez-Vargas:2023:PCS

[748] Danae Martinez-Vargas and Alejandro Murua-Sazo. Potts–Cox survival regression. Computational Statistics & Data Analysis, 187(??): ??, November 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001275.

Zhou:2023:BRQ

[749] Fei Zhou, Jie Ren, Shuangge Ma, and Cen Wu. The Bayesian regularized quantile varying coefficient model. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001196.

Jang:2023:PWL

[750] Hyun Jung Jang, Seung Jun Shin, and Andreas Artemiou. Principal weighted least square support vector machine: an online dimension-reduction tool for binary classification. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001299.

Liu:2023:OMV

[751] Wei Liu, Lan Luo, and Ling Zhou.

Online missing value imputation for high-dimensional mixed-type data via generalized factor models. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001330.

Bar:2023:GMC

[752] Haim Bar and Martin T. Wells.

On graphical models and convex geometry. Computational Statistics & Data Analysis, 187(??):??, November 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001111.

Anonymous:2023:D

[753] Anonymous. December 2023. Computational Statistics & Data Analysis, 188 (??):??, December 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2023:EBl

[754] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 188(??):??, December 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001573.

Kreuzer:2023:BMN

[755] Alexander Kreuzer, Luciana Dalla Valle, and Claudia Czado. Bayesian multivariate nonlinear state space copula models. Computational Statistics & Data Analysis, 188(??):??, December 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001317.

Gao:2023:ELM

[756] Zhenguo Gao, Xinye Wang, and Xiaoning Kang. Ensemble LDA via the modified Cholesky decomposition. Computational Statistics & Data Analysis, 188(??):??, December 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001342.

Aghabazaz:2023:URT

[757] Zeynab Aghabazaz and Iraj Kazemi. Under-reported time-varying MINAR(1) process for modeling multivariate count series. Computational Statistics & Data Analysis, 188(??):??, December 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947323001366.

Bao:2023:SPD

[758] Yajie Bao and Haojie Ren. Semiprofiled distributed estimation for highdimensional partially linear model. Computational Statistics & Data Analysis, 188(??):??, December 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001354.

Park:2023:RQR

[759] Seyoung Park, Hyunjin Kim, and Eun Ryung Lee. Regional quantile regression for multiple responses. Computational Statistics & Data Analysis, 188(??):??, December 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001378.

Chen:2023:BMS

[760] Cathy W. S. Chen, Chun-Shu Chen, and Mo-Hua Hsiung. Bayesian modeling of spatial integer-valued time series. Computational Statistics & Data Analysis, 188(??):??, December 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732300138X.

Deb:2023:NST

[761] Soudeep Deb and Sayar Karmakar. A novel spatio-temporal clustering algorithm with applications on COVID-19 data from the United States. Computational Statistics & Data Analysis, 188(??):??, December 2023. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001214.

Bagkavos:2023:GFT

[762] Dimitrios Bagkavos and Prakash N. Patil. Goodness-of-fit testing for normal mixture densities. Computational Statistics & Data Analysis, 188(??): ??, December 2023. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001263.

Anonymous:2024:Ja

[763] Anonymous. January 2024. Computational Statistics & Data Analysis, 189 (??):??, January 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2024:EBa

[764] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 189(??):??, January 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001779.

Moya:2024:FUA

[765] Blake Moya and Stephen G. Walker. Full uncertainty analysis for Bayesian nonparametric mixture models. Computational Statistics & Data Analysis, 189(??):??, January 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001494.

Zhu:2024:EBT

[766] Xiaonan Zhu, Yu Chen, and Jie Hu. Estimation of banded timevarying precision matrix based on SCAD and group lasso. Computational Statistics & Data Analysis, 189(??):??, January 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001603.

Liu:2024:VBI

[767] Jie Liu, Zifeng Ye, Kun Chen, and Panpan Zhang. Variational Bayesian inference for bipartite mixed-membership stochastic block model with applications to collaborative filtering. Computational Statistics & Data Analysis, 189(??):??, January 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001470.

Zhang:2024:DRS

[768] Cuihong Zhang, Jing Ning, Jianwen Cai, James E. Squires, Steven H. Belle, and Ruosha Li. Dynamic risk score modeling for multiple longitudinal risk factors and survival. Computational Statistics & Data Analysis, 189(??):??, January 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001482.

Adimari:2024:LTC

[769] Gianfranco Adimari, Duc-Khanh To, Monica Chiogna, Francesca Scatozza, and Antonio Facchiano. Likelihood-type confidence regions for optimal sensitivity and specificity of a diagnostic test. Computational Statistics & Data Analysis, 189(??):??, January 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001512.

Deng:2024:SSC

[770] Jiayi Deng, Danyang Huang, Yi Ding, Yingqiu Zhu, Bingyi Jing, and Bo Zhang. Subsampling spectral clustering for stochastic block models in large-scale networks. Computational Statistics & Data Analysis, 189(??):??, January 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001469.

Andre:2024:JMB

[771] L. M. André, J. L. Wadsworth, and A. O’Hagan. Joint modelling of the body and tail of bivariate data. Computational Statistics & Data Analysis, 189(??):??, January 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001524.

Zhou:2024:STS

[772] Niwen Zhou, Xu Guo, and Lixing Zhu. Significance test for semiparametric conditional average treatment effects and other structural functions. Computational Statistics & Data Analysis, 189(??):??, January 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001500.

Jin:2024:SEE

[773] Shaobo Jin and Youngjo Lee. Standard error estimates in hierarchical generalized linear models. Computational Statistics & Data Analysis, 189(??):??, January 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001639.

Hu:2024:TPF

[774] Yue Hu, Haiqi Li, and Falong Tan. Testing the parametric form of the conditional variance in regressions based on distance covariance. Computational Statistics & Data Analysis, 189 (??):??, January 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001627.

Pelaez:2024:PDE

[775] Rebeca Peláez, Ingrid Van Keilegom, Ricardo Cao, and Juan M. Vilar. Probability of default estimation in credit risk using mixture cure models. Computational Statistics & Data Analysis, 189(??):??, January 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001640.

Anonymous:2024:F

[776] Anonymous. February 2024. Computational Statistics & Data Analysis, 190 (??):??, February 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2024:EBb

[777] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 190(??):??, February 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001925.

Kao:2024:HEA

[778] Ming-Hung Kao and Ping-Han Huang. Hybrid exact-approximate design approach for sparse functional data. Computational Statistics & Data Analysis, 190(??):??, February 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001615.

Godichon-Baggioni:2024:RRR

[779] Antoine Godichon-Baggioni, Wei Lu, and Bruno Portier. Recursive ridge regression using second-order stochastic algorithms. Computational Statistics & Data Analysis, 190(??):??, February 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947323001652.

Maia:2024:GBN

[780] Mateus Maia, Keefe Murphy, and Andrew C. Parnell. GP-BART: a novel Bayesian additive regression trees approach using Gaussian processes. Computational Statistics & Data Analysis, 190(??):??, February 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732300169X.

Guo:2024:EHO

[781] Xiao Guo, Hai Zhang, and Xiangyu Chang. On the efficacy of higherorder spectral clustering under weighted stochastic block models. Computational Statistics & Data Analysis, 190 (??):??, February 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001834.

Zhang:2024:VSU

[782] Tonglin Zhang. Variables selection using L_{0} penalty. Computational Statistics & Data Analysis, 190(??): ??, February 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001718.

Yu:2024:LAQ

[783] Lili Yu and Yichuan Zhao. Laplace approximated quasi-likelihood method for heteroscedastic survival data. Computational Statistics & Data Analysis, 190(??):??, February 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001706.

Mou:2024:APL

[784] Xichen Mou and Dewei Wang. Additive partially linear model for pooled biomonitoring data. Computational Statistics & Data Analysis, 190(??): ??, February 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001731.

Zhang:2024:RST

[785] Qiuyan Zhang, Chen Wang, Baoxue Zhang, and Hu Yang. An RIHT statistic for testing the equality of several high-dimensional mean vectors under homoskedasticity. Computational Statistics & Data Analysis, 190(??):??, February 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001664.

Yoshida:2024:FSS

[786] Wataru Yoshida and Kei Hirose. Fast same-step forecast in SUTSE model and its theoretical properties. Computational Statistics & Data Analysis, 190(??):??, February 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794732300172X.

Li:2024:CRE

[787] Wei Li, Shanshan Luo, and Wangli Xu. Calibrated regression estimation using empirical likelihood under data fusion. Computational Statistics & Data Analysis, 190(??):??, February 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001822.

Zhao:2024:DCS

[788] Wenbiao Zhao and Lixing Zhu. Detecting change structures of nonparametric regressions. Computational Statistics & Data Analysis, 190(??): ??, February 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001676.

Anonymous:2024:Ma

[789] Anonymous. March 2024. Computational Statistics & Data Analysis, 191 (??):??, March 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2024:EBc

[790] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 191(??):??, March 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002074.

Zhao:2024:DDE

[791] Yan-Yong Zhao, Yuchun Zhang, Yuan Liu, and Noriszura Ismail. Distributed debiased estimation of high-dimensional partially linear models with jumps. Computational Statistics & Data Analysis, 191(??):??, March 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001688.

Heaton:2024:IML

[792] Matthew J. Heaton, Benjamin K. Dahl, Caleb Dayley, Richard L. Warr, and Philip White. Integrating machine learning and Bayesian nonparametrics for flexible modeling of point pattern data. Computational Statistics & Data Analysis, 191(??): ??, March 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732300186X.

Wang:2024:NQS

[793] Chuchu Wang and Xinyuan Song. Nonparametric quantile scalar-on-image regression. Computational Statistics & Data Analysis, 191(??):??, March 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947323001846.

Li:2024:BNE

[794] Yunzhe Li, Juhee Lee, and Athanasios Kottas. Bayesian nonparametric Erlang mixture modeling for survival analysis. Computational Statistics & Data Analysis, 191(??):??, March 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001858.

Weiss:2024:CMM

[795] Christian H. Weiß and Fukang Zhu. Conditional-mean multiplicative operator models for count time series. Computational Statistics & Data Analysis, 191(??):??, March 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001962.

Robertson:2024:OPP

[796] Blair Robertson and Chris Price.

One point per cluster spatially balanced sampling. Computational Statistics & Data Analysis, 191(??): ??, March 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001998.

Onizuka:2024:BBT

[797] Takahiro Onizuka, Fumiya Iwashige, and Shintaro Hashimoto. Bayesian boundary trend filtering. Computational Statistics & Data Analysis, 191 (??):??, March 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002001.

He:2024:NAP

[798] Xin He, Xiaojun Mao, and Zhonglei Wang. Nonparametric augmented probability weighting with sparsity. Computational Statistics & Data Analysis, 191(??):??, March 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002013.

Caamano-Carrillo:2024:NNW

[799] Christian Caamaño-Carrillo, Moreno Bevilacqua, Cristian López, and Víctor Morales-Oñate. Nearest neighbors weighted composite likelihood based on pairs for (non-)Gaussian massive spatial data with an application to Tukeyhh random fields estimation. Computational Statistics & Data Analysis, 191(??):??, March 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001986.

Lbath:2024:CBI

[800] Hanâ Lbath, Alexander Petersen, Wendy Meiring, and Sophie Achard. Clustering-based inter-regional correlation estimation. Computational Statistics & Data Analysis, 191(??): ??, March 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001871.

Anonymous:2024:A

[801] Anonymous. April 2024. Computational Statistics & Data Analysis, 192 (??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2024:EBd

[802] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947324000082.

Liang:2024:VSH

[803] Lixing Liang, Yipeng Zhuang, and Philip L. H. Yu. Variable selection for high-dimensional incomplete data.

Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323001883.

Fan:2024:FTL

[804] Xianqiu Fan, Jun Cheng, Hailing Wang, Bin Zhang, and Zhenzhen Chen. A fast trans-lasso algorithm with penalized weighted score function. Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002104.

Yang:2024:ENP

[805] Yuan Yang, Christopher S. McMahan, Yu-Bo Wang, and Yuyuan Ouyang. Estimation of l_{0} norm penalized models: a statistical treatment. Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732300213X.

Wang:2024:HEA

[806] Cheng Wang, Haozhe Chen, and Binyan Jiang. HiQR: an efficient algorithm for high-dimensional quadratic regression with penalties. Computational Statistics & Data Analysis, 192 (??):??, April 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002153.

Zhang:2024:GVS

[807] Xin Zhang and Junlong Zhao. Group variable selection via group sparse neural network. Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167947323002220.

Li:2024:BSS

[808] Yuanbo Li, Chu Kin Chan, Chun Yip Yau, Wai Leong Ng, and Henry Lam. Burn-in selection in simulating stationary time series. Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323001974.

Yamazoe:2024:SCR

[809] Hiroya Yamazoe and Kanta Naito. Simultaneous confidence region of an embedded one-dimensional curve in multi-dimensional space. Computational Statistics & Data Analysis, 192 (??):??, April 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947323002025.

Wu:2024:MBA

[810] Xiaofei Wu, Hao Ming, Zhimin Zhang, and Zhenyu Cui. Multi-block alternating direction method of multipliers for ultrahigh dimensional quantile fused regression. Computational Statistics & Data Analysis, 192 (??):??, April 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002128.

Kalogridis:2024:RAF

[811] Ioannis Kalogridis. Robust and adaptive functional logistic regression. Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002165.

Liang:2024:HFD

[812] Weijuan Liang, Qingzhao Zhang, and Shuangge Ma. Hierarchical false discovery rate control for high-dimensional survival analysis with interactions. Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002177.

Cai:2024:SIU

[813] Leheng Cai and Qirui Hu. Simultaneous inference and uniform test for eigensystems of functional data. Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002116.

Wu:2024:TST

[814] Qianyong Wu and Jiang Hu. Twosample test of stochastic block models.

Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002141.

Tortora:2024:LBM

[815] Cristina Tortora, Brian C. Franczak, Luca Bagnato, and Antonio Punzo. A Laplace-based model with flexible tail behavior. Computational Statistics & Data Analysis, 192(??): ??, April 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002207.

Goepp:2024:GBS

[816] Vivien Goepp and Jan van de Kassteele. Graph-based spatial segmentation of areal data. Computational Statistics & Data Analysis, 192 (??):??, April 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002190.

Cai:2024:SDB

[817] Tingting Cai, Jianbo Li, Qin Zhou, Songlou Yin, and Riquan Zhang. Subgroup detection based on partially linear additive individualized model with missing data in response. Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002219.

Peng:2024:ERO

[818] Cheng Peng, Drew P. Kouri, and Stan Uryasev. Efficient and robust optimal design for quantile regression based on linear programming. Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002037.

Filipiak:2024:DBS

[819] Katarzyna Filipiak, Daniel Klein, and Monika Mokrzycka. Discrepancy between structured matrices in the power analysis of a separability test. Computational Statistics & Data Analysis, 192(??):??, April 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002189.

Anonymous:2024:M

[820] Anonymous. May 2024. Computational Statistics & Data Analysis, 193 (??):??, May 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2024:EBe

[821] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 193(??):??, May 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000215.

Zhong:2024:OEE

[822] Chen Zhong. Oracle-efficient estimation and trend inference in nonstationary time series with trend and heteroscedastic ARMA error. Computational Statistics & Data Analysis, 193(??):??, May 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732400001X.

Shao:2024:ISI

[823] Lihui Shao, Jiaqi Wu, Weiping Zhang, and Yu Chen. Integrated subgroup identification from multi-source data. Computational Statistics & Data Analysis, 193(??):??, May 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000021.

Zhou:2024:CPD

[824] Houlin Zhou, Hanbing Zhu, and Xuejun Wang. Change point detection via feedforward neural networks with theoretical guarantees. Computational Statistics & Data Analysis, 193(??):??, May 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002244.

Fan:2024:GLS

[825] Xinyan Fan, Kuangnan Fang, Dan Pu, and Ruixuan Qin. Generalized latent space model for one-mode networks with awareness of two-mode networks. Computational Statistics & Data Analysis, 193(??):??, May 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002268.

Xue:2024:ELP

[826] Liugen Xue. Empirical likelihood in a partially linear single-index model with censored response data. Computational Statistics & Data Analysis, 193(??):??, May 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002232.

Hivert:2024:PCD

[827] Benjamin Hivert, Denis Agniel, Rodolphe Thiébaut, and Boris P. Hejblum. Postclustering difference testing: Valid inference and practical considerations with applications to ecological and biological data. Computational Statistics & Data Analysis, 193 (??):??, May 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732300227X.

Anonymous:2024:Jb

[828] Anonymous. June 2024. Computational Statistics & Data Analysis, 194 (??):??, June 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2024:EBf

[829] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 194(??):??, June 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000343.

Fu:2024:BWP

[830] Penghui Fu and Zhiqiang Tan. Blockwise primal-dual algorithms for largescale doubly penalized ANOVA modeling. Computational Statistics & Data Analysis, 194(??):??, June 2024. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S0167947324000161.

Bian:2024:UFA

[831] Yuan Bian, Grace Y. Yi, and Wenqing He. A unified framework of analyzing missing data and variable selection using regularized likelihood. Computational Statistics & Data Analysis, 194(??):??, June 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000033.

Winn-Nunez:2024:SAL

[832] Emily T. Winn-Nuñez, Maryclare Griffin, and Lorin Crawford. A simple approach for local and global variable importance in nonlinear regression models. Computational Statistics & Data Analysis, 194(??):??, June 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947323002256.

Hou:2024:HQR

[833] Zhaohan Hou and Lei Wang. Heterogeneous quantile regression for longitudinal data with subgroup structures. Computational Statistics & Data Analysis, 194(??):??, June 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000124.

Borrajo:2024:GFT

[834] M. I. Borrajo, W. González-Manteiga, and M. D. Martínez-Miranda. Goodnessof-fit test for point processes firstorder intensity. Computational Statistics & Data Analysis, 194(??):??, June 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0167947324000136.

Gagnon:2024:RHT

[835] Philippe Gagnon and Yuxi Wang.

Robust heavy-tailed versions of generalized linear models with applications in actuarial science. Computational Statistics & Data Analysis, 194(??):??, June 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000045.

Li:2024:PER

[836] Shuaiyu Li, Yunpei Wu, and Yuzhong Cheng. Parameter estimation and random number generation for student Lévy processes. Computational Statistics & Data Analysis, 194 (??):??, June 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000173.

Marco:2024:FRE

[837] Nicholas Marco, Damla Şentürk, Shafali Jeste, Charlotte C. DiStefano, Abigail Dickinson, and Donatello Telesca. Flexible regularized estimation in highdimensional mixed membership models. Computational Statistics & Data Analysis, 194(??):??, June 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732400015X.

Ma:2024:SPR

[838] Yijia Ma, Xinyu Zhou, and Wei Wu. A stochastic process representation for time warping functions. Computational Statistics & Data Analysis, 194(??):??, June 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000252.

Mondal:2024:IOR

[839] Anjana Mondal and Somesh Kumar. Inference on order restricted means of inverse Gaussian populations under heteroscedasticity. Computational Statistics & Data Analysis, 194(??):??, June 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000276.

You:2024:SEM

[840] Na You, Hongsheng Dai, Xueqin Wang, and Qingyun Yu. Sequential estimation for mixture of regression models for heterogeneous population. Computational Statistics & Data Analysis, 194(??):??, June 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000264.

Singh:2024:FSS

[841] Rakhi Singh and John Stufken. Factor selection in screening experiments by aggregation over random models. Computational Statistics & Data Analysis, 194(??):??, June 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000240.

Anonymous:2024:Jc

[842] Anonymous. July 2024. Computational Statistics & Data Analysis, 195 (??):??, July 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Anonymous:2024:EBg

[843] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 195(??):??, July 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000495.

Pereira:2024:NAI

[844] Diogo Pereira, Cláudia Nunes, and Rui Rodrigues. A new algorithm for inference in HMM’s with lower span complexity. Computational Statistics & Data Analysis, 195 (??):??, July 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000392.

Dargel:2024:PSR

[845] Lukas Dargel and Christine Thomas-Agnan. Pairwise share ratio interpretations of compositional regression models. Computational Statistics & Data Analysis, 195 (??):??, July 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794732400029X.

Lin:2024:VSK

[846] Xiefang Lin and Fang Fang. Variable selection of Kolmogorov–Smirnov maximization with a penalized surrogate loss. Computational Statistics & Data Analysis, 195(??):??, July 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000288.

Lo:2024:SMC

[847] Simon M. S. Lo, Ralf A. Wilke, and Takeshi Emura. A semiparametric model for the cause-specific hazard under risk proportionality. Computational Statistics & Data Analysis, 195(??):??, July 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000379.

Okazaki:2024:MTL

[848] Akira Okazaki and Shuichi Kawano. Multi-task learning regression via convex clustering. Computational Statistics & Data Analysis, 195 (??):??, July 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000409.

Liu:2024:VBA

[849] Wenting Liu, Huiqiong Li, Niansheng Tang, and Jun Lyu. Variational Bayesian approach for analyzing interval-censored data under the proportional hazards model. Computational Statistics & Data Analysis, 195(??):??, July 2024. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000410.

Seo:2024:SAF

[850] Byungtae Seo and Il Do Ha. Semiparametric accelerated failure time models under unspecified random effect distributions. Computational Statistics & Data Analysis, 195 (??):??, July 2024. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947324000422.